### v6

# Robust Design and Evaluation of

Abstract
Predictive algorithms inform consequential decisions in settings with selective labels: outcomes are observed only for units selected by past decision makers. This creates an identification problem under unobserved confounding — when selected and unselected units differ in unobserved ways that affect outcomes. We propose a framework for robust design and evaluation of predictive algorithms that bounds how much outcomes may differ between selected and unselected units with the same observed characteristics. These bounds formalize common empirical strategies including proxy outcomes and instrumental variables. Our estimators work across bounding strategies and performance measures such as conditional likelihoods, mean square error, and true/false positive rates. Using administrative data from a large Australian financial institution, we show that varying confounding assumptions substantially affects credit risk predictions and fairness evaluations across income groups.
Keywords: predictive algorithms, missing data, confounding, partial identification.
JEL Codes: C14, C52, G20
1 Introduction
Predictive algorithms inform high-stakes decisions in pretrial release, consumer lending, health care, and many other domains. Designing and evaluating such tools is often complicated by the “selective labels” problem: outcomes are observed only for units selected by past decision makers. For example, pretrial risk assessments predict the likelihood defendants would commit pretrial misconduct if they were released before trial; but we only observe whether a defendant committed misconduct if a judge decided to release them (e.g., KLLLM(18); Rambachan(21); AngelovaEtAl(22)). Consumer credit scores predict the likelihood of default if an applicant were granted a loan; but we only observe whether an applicant defaulted if the financial institution approved them and they accepted the loan terms (e.g., BlattnerNelson(21); CostonEtAl(21)-Rashomon). [^1] The selective labels problem arises when analyzing predictive algorithms in many other empirical settings as well, such as medical diagnosis (e.g., MullainathanObermeyer(21)), child protective services (e.g., ChouldechovaEtAl(18)), hiring (e.g., LiRaymondBergman(20)), education (BergmanEtAl(21)), and tax audits (e.g., BlackEtAl(22)-TaxAudits; BattagliniEtAl(22); ElzaynEtAl(23)).
The selective labels problem creates an identification challenge under unobserved confounding–when selected and unselected units differ in ways not captured by observed covariates. Judges may release defendants based on courtroom interactions not in administrative data; loan applicants may accept offers based on competing credit options unavailable to researchers. Ignoring unobserved confounding leads to inaccurate predictions and misleading evaluations of predictive algorithms. [^2] While there exists methods for designing and evaluating predictive algorithms in computer science from selectively observed data (e.g., Coston(20)-Evaluation; CostonEtAl(21)-Rashomon; MishlerEtAl(21); mishler2021fade; guerdan2023counterfactual), they tackle this challenge assuming that the historical selection decision was unconfounded conditional on observed covariates. Existing solutions are limited. Applied researchers rely on informal heuristics—imputing missing outcomes using proxy variables (ChouldechovaEtAl(18); BlattnerNelson(21); MullainathanObermeyer(21)) or applying ad-hoc adjustments to observed outcomes among selected units. Recent work uses instrumental variables but typically focuses on specific estimands (lakkaraju2017selective; ArnoldDobbieHull(20)-discrimination_alg) or predictive algorithms that are linear in covariates (ArnoldEtAl(24)).
We develop a framework for robust design and evaluation of predictive algorithms under unobserved confounding. We bound how much outcomes may differ between selected and unselected units with the same observed characteristics. These bounds formalize common empirical strategies — including proxy outcomes and instrumental variables — and partially identify key predictive estimands such as conditional likelihoods, mean square error, calibration, and true/false positive rates. Our estimation approach works across multiple bounding strategies and performance measures, flexibly incorporating machine learning methods.
We consider data $(X_{i},D_{i},Y_{i})$ for $i=1,\ldots,n$ drawn i.i.d. from a joint distribution $\mathbb{P}(\cdot)$, where $X_{i}$ is a vector of covariates and $D_{i}\in\{0,1\}$ is a selection decision by past decision makers (e.g., judges or loan applicants). The observed outcome satisfies $Y_{i}=D_{i}\times Y_{i}^{*}$ for some true outcome $Y_{i}^{*}\in\{0,1\}$; for example, $Y_{i}^{*}$ is whether a defendant would commit pretrial misconduct if released before trial or whether an applicant would default if granted a loan.
Researchers have two objectives. First, design predictive algorithms $s(X_{i})$ that predict the probability $Y_{i}^{*}=1$ given the covariates $X_{i}$ and construct data-driven decision rules. Second, evaluate existing predictive algorithms through performance measures, such as overall performance given by $\mbox{perf}(s;\beta)=\mathbb{E}[\beta_{0}(X_{i};s)+\beta_{1}(X_{i};s)Y_{i}^{*}]$, and class-specific performance given by $\mbox{perf}_{+}(s;\beta)=\mathbb{E}[\beta_{0}(X_{i};s)\mid Y_{i}^{*}=1]$ and $\mbox{perf}_{-}(s;\beta)=\mathbb{E}[\beta_{0}(X_{i};s)\mid Y_{i}^{*}=0]$, where $\beta_{0}(X_{i};s),\beta_{1}(X_{i};s)$ are researcher-specified functions. Alternative choices recover popular diagnostics such as mean square error, true positive rate, and false positive rate. Comparing these measures across sensitive attributes evaluates fairness properties (e.g., MitchellEtAl(19); barocas-hardt-narayanan).
Rather than assuming unconfounded selection, we bound the confounding function $\delta(x):=\mathbb{P}(Y_{i}^{*}=1\mid D_{i}=0,X_{i}=x)-\mathbb{P}(Y_{i}^{*}=1\mid D_{i}=1,X_{i}=x)$ using researcher-specified bounding functions that may depend on covariates and identified nuisance parameters. The confounding function measures how much unselected units differ from selected units with the same characteristics; bounding it formalizes assumptions about unobserved confounding severity without requiring a full selection model. Under these bounds, the conditional likelihood and predictive performance measures are partially identified.
Our framework encompasses popular strategies for addressing selective labels as special cases. When researchers observe proxy outcomes correlated with $Y_{i}^{*}$ — such as default on other credit products (BlattnerNelson(21)) or long-term health outcomes (MullainathanObermeyer(21); ChanGentzkowYu(21)) — this implies specific bounds. When researchers have instrumental variables like randomly assigned decision makers (lakkaraju2017selective; KLLLM(18)), this implies alternative bounds. Our framework accommodates other plausible identifying assumptions.
We apply this framework to two objectives. First, we develop estimators for the bounds on the conditional likelihood $P(Y_{i}^{*}=1\mid X_{i}=x)$ (e.g., the pretrial misconduct rate or default rate given covariates), enabling design of predictive algorithms robust to unobserved confounding. Building on advances in causal machine learning (NieWager(20); kennedy2020optimal), we propose pseudo-outcome regression estimators with sample splitting: estimate nuisance parameters on one fold, construct influence-function-based pseudo-outcomes, then regress them on covariates using machine learning on a second fold. A key innovation is an oracle inequality extending kennedy2020optimal’s pointwise analysis to integrated mean square error: our feasible estimators approximate an infeasible oracle up to a smoothed, doubly-robust product of first-stage errors. This applies to any regression method satisfying a mild stability condition, including random forests, kernel methods, and series estimators. We leverage this to characterize integrated mean square error convergence rates and establish regret bounds for plug-in decision rules, connecting estimation guarantees to decision-theoretic performance.
Second, we develop estimators for bounds on the performance of a given predictive algorithm $s(\cdot)$ under unobserved confounding. For overall performance measures $\mbox{perf}(s;\beta)$ — which nest mean square error, calibration, and precision — we construct debiased estimators using crossfitting and debiased machine learning (e.g., RobinsEtAl(08); ZhengvanderLaan(11); ChernozhukovEtAl(18)-DML). Under mild regularity conditions, these estimators are $\sqrt{n}$-consistent and jointly asymptotically normal. For class-specific performance $\mbox{perf}_{+}(s;\beta)$ and $\mbox{perf}_{-}(s;\beta)$ — such as true and false positive rates — we show sharp bounds correspond to optimal values of linear fractional programs with nuisances in both the objective and constraints. We estimate using two folds: estimate nuisances and bounding functions on the first fold, then solve the sample optimization program on the second fold. We establish partial double robustness: estimation error decomposes into sampling error, a doubly-robust product for objective nuisances, and constraint estimation errors entering through root mean square error. Proving this exploits the optimization structure of linear fractional programs. [^3] Since linear fractional programs arise in several related sensitivity analysis frameworks in causal inference (e.g., AronowLee(2013); MiratrixWagerZubizaretta(18); KallusMaoZhou(18)-IntervalEstimation; ZhaoSmallBhattacharya(19); KallusZhou(21)), our results may be broadly applicable beyond the selective labels setting. Our estimation strategies accommodate multiple bounding strategies, flexibly incorporating domain-specific knowledge while maintaining statistical guarantees.
As an empirical illustration, we apply our framework to administrative data on personal loan applications from a large Australian bank. First, we construct confounding-robust credit risk scores under varying confounding assumptions. The resulting algorithms produce substantially different risk rankings that could meaningfully affect credit access. Second, inspired by recent work in consumer finance (BlattnerNelson(21); FusterEtAl(22)predictably; DiMaggioEtAl(22)-Fintech), we evaluate a benchmark credit score’s performance across income groups. An analysis ignoring unobserved confounding suggests the score is more accurate (in mean square error) for higher-income applicants. However, we show this disparity disappears at plausible levels of unobserved confounding, illustrating how researchers can assess the robustness of fairness conclusions.
Our framework prioritizes breadth and applicability. Rather than bespoke methods for each combination of bounding strategy and performance measure, we provide a unified approach for a variety of assumptions and performance measures. We deliver practical estimation procedures that enable researchers to robustly design and evaluate predictive algorithms under unobserved confounding.
Related Work:
Our analysis builds on foundational work on partial identification with selective labels (e.g., Manski(89); Manski(90); Manski(94)-SampleSelectionBounds). We generalize Manski(03)’s “approximate mean independence” assumption by incorporating covariates and nuisance parameters into the bounding functions, enabling flexible formalization of empirical heuristics using proxy outcomes and other strategies. See also diaz2013sensitivity; luedtke2015statistics; diaz2018sensitivity for similar bounding assumptions without covariates. We characterize identified sets for novel estimands arising in predictive algorithm evaluation.
Instrumental variable strategies are widely used to address unobserved confounding in this setting. lakkaraju2017selective and KLLLM(18) evaluate specific performance measures by returning a single point in the identified set. ArnoldDobbieHull(20)-discrimination_alg; ArnoldEtAl(24) extrapolate across instrument values using identification-at-infinity arguments. We instead derive covariate-dependent bounds following Manski(94)-SampleSelectionBounds, accommodating IVs as a special case while also encompassing other identifying strategies. [^4] In related work, LevisEtAl(23)-CovariateAssistedIV analyze covariate-assisted Balke-Pearl bounds on the average treatment effect. semenova2023generalized generalizes the Lee(09) bounds on the average treatment effect on the always takers, which requires a binary instrument that monotonically affects selection, to incorporate covariate information. Monotonicity is particularly unlikely to hold when the instrument arises from the random assignment of decision makers. See, for example, FrandsenLefgrenLeslie(19); ArnoldDobbieHull(20)-discrimination_bail; ChanGentzkowYu(21).
Our work relates to sensitivity analysis in causal inference, where researchers bound treatment effects under relaxations of unconfoundedness. Influential frameworks include the marginal sensitivity model (e.g., Tan(06); KallusMaoZhou(18)-IntervalEstimation; ZhaoSmallBhattacharya(19); DornGuoKallus(21); KallusZhou(21); DornGuo(21)), Rosenbaum’s sensitivity model (e.g., Rosenbaum(02); yadlowsky2018bounds), and partial $c$-dependence (e.g., MastenPoirier(18); MastenPoirier(20)). These models bound how unobservables affect selection propensities—assumptions that may be difficult to formulate when decisions involve human judgment. We instead bound outcome differences, which may be more intuitive in applied settings. Nonetheless, these models imply specific bounding functions in our framework, as we discuss in Remark 2. chernozhukov2022long develop omitted variable bounds for linear functionals; this covers our overall performance measures but not class-specific performance.
2 Setup and Identification Framework
We consider data $(X_{i},D_{i},Y_{i})$ for $i=1,\ldots,n$ drawn i.i.d. from a joint distribution $\mathbb{P}(\cdot)$, where $X_{i}\in\mathcal{X}\subseteq\mathbb{R}^{d}$ are covariates and $D_{i}\in\{0,1\}$ is the decision. There is a “selective labels” problem: the observed outcome satisfies $Y_{i}=D_{i}\,Y_{i}^{*}$ for some true outcome $Y_{i}^{*}\in\{0,1\}$ (e.g., KLLLM(18)). We assume $0<\mathbb{P}(Y_{i}^{*}=1)<1$ so that we face a non-trivial prediction problem and that the decision satisfies strict overlap (i.e., there exists $\epsilon>0$ such that $\mathbb{P}(D_{i}=1\mid X_{i}=x)\geq\epsilon$ with probability one).
Example: lending decisions
We observe data on loan applications, where $X_{i}$ contains information reported on the application such as income and prior credit history, $D_{i}$ is whether the application was granted a loan, and $Y_{i}=D_{i}\,Y_{i}^{*}$ is whether the applicant defaulted if granted the loan. We study credit scores that predict whether a new applicant would default $Y_{i}^{*}=1$ (e.g., BlattnerNelson(21); DiMaggioEtAl(22)-Fintech; FusterEtAl(22)predictably). $\blacktriangle$
Example: pretrial release
We observe data on defendants, where $X_{i}$ contains information about the defendant such as their current charge and prior conviction history, $D_{i}$ is whether the defendant was released before trial, and $Y_{i}=D_{i}\,Y_{i}^{*}$ is whether the defendant committed pretrial misconduct if released. We study pretrial risk tools that predict whether a new defendant would commit pretrial misconduct $Y_{i}^{*}=1$ (e.g., KLLLM(18); ArnoldDobbieHull(20)-discrimination_alg; AngelovaEtAl(22)). $\blacktriangle$
Example: medical testing
We observe data on patients, where $X_{i}$ contains patient information such as vital signs at admission and prior medical history, $D_{i}$ is whether a doctor decided to conduct a diagnostic test such as a cardiac stress test or lab test, and $Y_{i}=D_{i}\,Y_{i}^{*}$ is whether the patient was diagnosed with a heart attack or bacterial infection. We study medical risk tools that predict whether a new patient suffered a heart attack or bacterial infection $Y_{i}^{*}=1$ (e.g., MullainathanObermeyer(21); HuangEtAl(22)). $\blacktriangle$
Let $\pi_{d}(x):=\mathbb{P}(D_{i}=d\mid X_{i}=x)$ for $d\in\{0,1\}$ be the propensity score and $\mu_{1}(x):=\mathbb{P}(Y_{i}^{*}=1\mid D_{i}=1,X_{i}=x)$. Averages of a random variable $V_{i}$ are $\mathbb{E}_{n}[V_{i}]:=n^{-1}\sum_{i=1}^{n}V_{i}$. We let $\|\cdot\|$ denote the appropriate $L_{2}$-norm by context, where $\|f\|=\left(\int f(v)^{2}d\mathbb{P}(v)\right)^{1/2}$ for a measurable function $f(\cdot)$ and $\|v\|=\left(\sum_{j=1}^{k}v_{j}^{2}\right)^{1/2}$ for a vector $v\in\mathbb{R}^{k}$.
2.1 Predictive Algorithms and Performance Measures
A predictive algorithm $s(\cdot)\colon\mathcal{X}\rightarrow[0,1]$ predicts the probability $Y_{i}^{*}=1$ given covariates $X_{i}$, therefore estimating the conditional likelihood $\mu^{*}(x):=\mathbb{P}(Y_{i}^{*}=1\mid X_{i}=x)$. Given a predictive algorithm, researchers evaluate its performance in two ways.
First, given a predictive algorithm, researchers construct decision rules $d(\cdot)\colon\mathcal{X}\rightarrow\{0,1\}$ by applying a threshold rule; that is, defining $d(X_{i})=1\{s(X_{i})\geq\tau\}$ for some threshold $\tau\in[0,1]$. Alternative decision rules are evaluated based on their expected payoff
$U(d):=\mathbb{E}[(-u_{1,1}(X_{i})Y_{i}^{*}+u_{1,0}(X_{i})(1-Y_{i}^{*}))d(X_{i})+(-u_{0,0}(X_{i})(1-Y_{i}^{*})+u_{0,1}(X_{i})Y_{i}^{*})(1-d(X_{i})],$ (1)
where each $u_{d,y^{*}}(\cdot)\geq 0$ for $d,y^{*}\in\{0,1\}^{2}$ specifies the known payoff associated with a combination of decision $D_{i}$ and outcome $Y_{i}^{*}$ at features $X_{i}$. Payoffs are normalized so that $\sum_{d,y^{*}\in\{0,1\}^{2}}u_{d,y^{*}}(x)=1$ for all $x\in\mathcal{X}$. In consumer lending, the profitability of approving customers that would not default $u_{1,0}(\cdot)$ may vary based on features like the requested loan size. In pretrial release, the societal benefits of releasing defendants that would not commit misconduct $u_{1,0}(\cdot)$ may vary based on the defendant’s age, charge severity, or prior history of pretrial misconduct (e.g., Baughman(17)).
Second, given a predictive algorithm $s(\cdot)$, researchers evaluate its performance over the full population or specific groups. The target parameters are its overall performance
$\mbox{perf}(s;\beta)=\mathbb{E}[\beta_{0}(X_{i};s)+\beta_{1}(X_{i};s)Y_{i}^{*}]$ (2)
and its positive class and negative class performance
$\mbox{perf}_{+}(s;\beta)=\mathbb{E}[\beta_{0}(X_{i};s)\mid Y_{i}^{*}=1]\mbox{ and }\mbox{perf}_{-}(s;\beta)=\mathbb{E}[\beta_{0}(X_{i};s)\mid Y_{i}^{*}=0],$ (3)
where $\beta_{0}(X_{i};s),\beta_{1}(X_{i};s)$ are researcher-specified functions that may depend on covariates and the predictive algorithm. Alternative choices of $\beta_{0}(X_{i};s),\beta_{1}(X_{i};s)$ recover popular evaluations used in empirical research. We discuss several leading examples.
Example: mean square error and calibration
Mean square error is $\mathbb{E}[\left(s(X_{i})-Y_{i}^{*}\right)^{2}]=\mbox{perf}(s;\beta)$ for $\beta_{0}(X_{i};s)=s^{2}(X_{i})$, $\beta_{1}(X_{i};s)=1-2s(X_{i})$. Calibration at prediction bin $r_{1},r_{2}]\subseteq[0,1]$ is $\mathbb{E}[Y_{i}^{*}\mid r_{1}\leq s(X_{i})\leq r_{2}]=\mbox{perf}(s;\beta)$ for $\beta_{0}(X_{i};s):=0$, $\beta_{1}(X_{i};s):=\frac{1\{r_{1}\leq s(X_{i})\leq r_{2}\}}{\mathbb{P}\left(r_{1}\leq s(X_{i})\leq r_{2}\right)}$ assuming $\mathbb{P}(r_{1}\leq s(X_{i})\leq r_{2})>0$. $\blacktriangle$
Example: generalized true positive and false positive rates
Generalized true positive rate is $\mathbb{E}[s(X_{i})\mid Y_{i}^{*}=1]=\mbox{perf}_{+}(s;\beta)$ for $\beta_{0}(X_{i};s)=s(X_{i})$, and generalized false positive rate is $\mathbb{E}[s(X_{i})\mid Y_{i}^{*}=0]=\mbox{perf}_{-}(s;\beta)$ for $\beta_{0}(X_{i};s)=s(X_{i})$. $\blacktriangle$
Example: ROC curve
The true positive rate or recall at threshold $\tau$ is $\mathbb{P}\left(s(X_{i})\geq\tau\mid Y_{i}^{*}=1\right)=\mbox{perf}_{+}(s;\beta)$ for $\beta_{0}(X_{i};s)=1\{s(X_{i})\geq\tau\}$. The false positive rate at threshold $\tau$ is $\mathbb{P}\left(s(X_{i})\geq\tau\mid Y_{i}^{*}=0\right)=\mbox{perf}_{-}(s;\beta)$. The Receiver Operating Characteristic (ROC) curve $\{(\mathbb{P}(s(X_{i})\geq\tau\mid Y_{i}^{*}=0),\mathbb{P}(s(X_{i})\geq\tau\mid Y_{i}^{*}=1))\colon\tau\in[0,1]\}$ summarizes the predictive algorithm’s ability to separate the positive class $Y_{i}^{*}=1$ from the negative class $Y_{i}^{*}=0$ as $\tau$ varies. $\blacktriangle$
Example: precision-recall curve
The precision at threshold $\tau$ is $\mathbb{P}\left(Y_{i}^{*}=1\mid s(X_{i})\geq\tau\right):=\mbox{perf}(s;\beta)$ for $\beta_{0}(X_{i};s)=0$ and $\beta_{1}(X_{i};s)=\frac{1\{s(X_{i})\geq\tau\}}{\mathbb{P}\left(s(X_{i})\geq\tau\right)}$ assuming $\mathbb{P}(s(X_{i})\geq\tau)>0$. The precision-recall curve $\{\left(\mathbb{P}\left(s(X_{i})\geq\tau\mid Y_{i}^{*}=1\right),\mathbb{P}\left(Y_{i}^{*}=1\mid s(X_{i})\geq\tau\right)\right)\colon\tau\in[0,1]\}$ summarizes the predictive algorithm’s ability to accurately classify $Y_{i}^{*}=1$ as $\tau$ varies. $\blacktriangle$
Returning to the earlier examples, BlattnerNelson(21), FusterEtAl(22)predictably, and DiMaggioEtAl(22)-Fintech compare mean square error, ROC curves, and precision-recall curves of alternative predictive algorithms for default likelihood in consumer finance. KLLLM(18) and MullainathanObermeyer(21) analyze predictive algorithms in pretrial release and medical diagnosis by reporting calibration at various prediction bins and ROC curves. [^5] In analyzing pretrial risk assessments, lakkaraju2017selective and AngelovaEtAl(22) consider the “failure rate” or “counterfactual misconduct rate” $\mathbb{P}(Y_{i}^{*}=1,s(X_{i})\leq\tau)$ at threshold $\tau\in[0,1]$. This is recovered by the overall performance $\mbox{perf}(s;\beta)$ for $\beta_{0}(X_{i};s)=0,\beta_{1}(X_{i};s)=1\{s(X_{i})\leq\tau\}$.
2.2 Identification
Selective labels create an identification problem: since $Y_{i}^{*}$ is only observed when $D_{i}=1$, the conditional likelihood $\mathbb{P}(Y_{i}^{*}=1\mid D_{i}=0,X_{i})$ is not point identified from the data alone. Consequently, neither are the conditional likelihood, expected payoffs for any decision rule, nor predictive performance measures defined in Section 2.1.
A natural approach would be to assume unconfounded selection — that is, $Y_{i}^{*}\perp\!\!\!\perp D_{i}\mid X_{i}$ (e.g., Coston(20)-Evaluation; mishler2021fade). Under unconfoundedness, the conditional likelihood among unselected units equals that among selected units with the same covariates: $P(Y_{i}^{*}=1\mid D_{i}=0,X_{i})=P(Y_{i}^{*}=1\mid D_{i}=1,X_{i})$. However, this assumption is often implausible in practice. Selected and unselected units may differ systematically in ways not captured by observed covariates $X_{i}$. In consumer lending, applicants’ decisions to accept loan terms may depend on competing offers from other lenders or their private assessment of their ability to repay — information unavailable to the researcher. In pretrial release, judges may rely on courtroom interactions or details from case files that are not recorded in administrative data. Consequently, unobserved confounding — unobservables correlated with both the outcome $Y_{i}^{*}$ and the decision $D_{i}$ — is a central challenge.
Rather than imposing unconfoundedness, we partially identify the conditional likelihood, expected payoffs, and predictive performance measures by placing restrictions on how much outcomes can differ between unselected and selected units with the same covariates. These restrictions formalize researcher beliefs about the severity of unobserved confounding in a transparent way and nest popular empirical strategies as special cases. The key to our approach is bounding the confounding function $\delta(x):=\mathbb{P}(Y_{i}^{*}=1\mid D_{i}=0,X_{i}=x)-\mathbb{P}(Y_{i}^{*}=1\mid D_{i}=1,X_{i}=x)$, which summarizes the difference in outcomes between unselected and selected units conditional on covariates. When $\delta(x)=0$ for all $x\in\mathcal{X}$, the selection decision is unconfounded. When $\delta(x)\neq 0$, unobserved confounding is present: units with the same observed covariates have systematically different outcomes depending on whether they were selected.
Assumption 2.1.
For researcher-specified bounding functions $\underline{\delta}(\cdot;\eta),\overline{\delta}(\cdot;\eta)$ with (possibly infinite-dimensional) nuisance parameters $\eta$, the confounding function satisfies
$\underline{\delta}(x;\eta)\leq\delta(x)\leq\overline{\delta}(x;\eta)\mbox{ for all }x\in\mathcal{X}.$ (4)
Let $\Delta$ denote the set of all confounding functions satisfying Equation (4).
Assumption 2.1 restricts how much outcomes can differ between unselected and selected units with the same observed covariates. In consumer lending, this bounds the difference in default rates between unfunded and funded applicants conditional on application characteristics. In pretrial release, this bounds the difference in misconduct rates between detained and released defendants conditional on charge and criminal history.
Assumption 2.1 allows researchers to encode beliefs about unobserved confounding without committing to a specific selection model. It nests two important special cases: unconfoundedness with $\underline{\delta}(x;\eta)=\overline{\delta}(x;\eta)=0$ for all $x\in\mathcal{X}$, and the assumption-free, worst-case bounds with $\underline{\delta}(x;\eta)=-\mathbb{P}(Y_{i}^{*}=1\mid D_{i}=1,X_{i}=x)$, $\overline{\delta}(x;\eta)=1-\mathbb{P}(Y_{i}^{*}=1\mid D_{i}=1,X_{i}=x)$ (e.g., Manski(89); Manski(94)-SampleSelectionBounds). Thus Assumption 2.1 interpolates between strong identifying assumptions and possibly uninformative bounds.
More broadly, Assumption 2.1 follows the structure of widely-used partial identification frameworks in econometrics, most notably Manski(03)’s ”approximate mean independence” assumption. It extends this by conditioning on covariates and allowing bounds to depend on identified nuisance parameters. In this sense, Assumption 2.1 formalizes restrictions in terms that align with domain knowledge: researchers in these settings naturally reason about outcome differences rather than, say, how unobservables affect selection propensities. As shown in Section 2.3, the bounding functions can be chosen through direct calibration based on domain expertise or derived indirectly from additional data features such as proxy outcomes or instrumental variables. Each strategy is instantiated by specifying both the functional form of the bounding functions and the collection of nuisance parameters $\eta$ (possibly infinite-dimensional) to be estimated.
Under Assumption 2.1, the conditional likelihood and evaluation estimands are partially identified. We assume throughout that the chosen bounding functions yield valid conditional probabilities. Specifically, for all $\delta\in\Delta$ and $x\in\mathcal{X}$, $\mu_{1}(x)+\pi_{0}\delta(x;\eta)\in[0,1]$ and $\delta(x;\eta)\in[-1,1]$. This is a mild restriction in practice; as we discuss below, for common bounding functions, these constraints can be enforced through the researcher’s specification of sensitivity parameters or are satisfied automatically, as we will discuss next. If these constraints are violated and the researcher proceeds with our identification framework, the resulting bounds will be conservative (non-sharp) but may still be informative.
Since $\mu^{*}(x)=\mu_{1}(x)+\pi_{0}(x)\delta(x)$, the identified set is the set of all values consistent with the bounds on $\delta(x)$,
$\mathcal{H}(\mu^{*}(x))=\left\{\mu_{1}(x)+\delta(x)\pi_{0}(x)\mbox{ such that }\delta(x)\in\Delta\right\}.$
The identified sets for the performance measures, $\mathcal{H}(\mbox{perf}(s;\beta))$ and $\mathcal{H}(\mbox{perf}_{+}(s;\beta))$, and for the expected payoff of a decision rule $\mathcal{H}(U(d))$ are defined analogously. As shorthand, let $\beta_{0,i}:=\beta_{0}(X_{i};s),\beta_{1,i}=\beta_{1}(X_{i};s)$ and $\underline{\delta}_{i}:=\underline{\delta}(X_{i};\eta)$, $\overline{\delta}_{i}:=\overline{\delta}_{i}(X_{i};\eta)$.
Lemma 2.1.
For all $x\in\mathcal{X}$, $\mathcal{H}(\mu^{*}(x))=\left[\underline{\mu}^{*}(x),\overline{\mu}^{*}(x)\right],$ where $\overline{\mu}^{*}(x)=\mu_{1}(x)+\pi_{0}(x)\overline{\delta}(x;\eta)$, $\underline{\mu}^{*}(x)=\mu_{1}(x)+\pi_{0}(x)\underline{\delta}(x;\eta)$. Furthermore,
[…]
where $\overline{\mbox{perf}}(s;\beta)=\mathbb{E}[\beta_{0,i}+\beta_{1,i}\mu_{1}(X_{i})+\beta_{1,i}\pi_{0}(X_{i})\left(1\{\beta_{1,i}>0\}\overline{\delta}_{i}+1\{\beta_{1,i}\leq 0\}\underline{\delta}_{i}\right)]$, $\underline{\mbox{perf}}(s;\beta)=\mathbb{E}[\beta_{0,i}+\beta_{1,i}\mu_{1}(X_{i})+\beta_{1,i}\pi_{0}(X_{i})\left(1\{\beta_{1,i}\leq 0\}\overline{\delta}_{i}+1\{\beta_{1,i}>0\}\underline{\delta}_{i}\right)]$, and
$\overline{\mbox{perf}}_{+}(s;\beta)=\sup_{\delta\in\Delta}\mathbb{E}[\mu_{1}(X_{i})+\pi_{0}(X_{i})\delta(X_{i})]^{-1}\mathbb{E}[\beta_{0,i}\mu_{1}(X_{i})+\beta_{0,i}\pi_{0}(X_{i})\delta(X_{i})],$
$\underline{\mbox{perf}}_{+}(s;\beta)=\inf_{\delta\in\Delta}\mathbb{E}[\mu_{1}(X_{i})+\pi_{0}(X_{i})\delta(X_{i})]^{-1}\mathbb{E}[\beta_{0,i}\mu_{1}(X_{i})+\beta_{0,i}\pi_{0}(X_{i})\delta(X_{i})].$
Lemma 2.2.
For $d(\cdot)\colon\mathcal{X}\rightarrow\{0,1\}$ and $u_{d,y^{*},i}=u_{d,y^{*}}(X_{i})$, $\mathcal{H}(U(d))=[\underline{U}(d),\overline{U}(d)]$ for $\underline{U}(d),\overline{U}(d)$ defined in the proof.
The remainder of the paper applies this partial identification framework to the two objectives for predictive algorithms commonly seen in applied work. Section 3 addresses algorithm design: we estimate bounds on the conditional likelihood, thereby constructing predictive algorithms robust to unobserved confounding. Section 4 addresses algorithm evaluation: we estimate bounds on performance measures for a given predictive algorithm. Before proceeding to these estimation results, we next discuss how researchers might choose the bounding functions in Assumption 2.1.
Remark 1 (Connection to Algorithmic Fairness).
Differences in predictive performance across sensitive attributes (e.g., race, gender) formalize violations of popular fairness criteria studied in the algorithmic fairness literature (see MitchellEtAl(19); barocas-hardt-narayanan, for reviews). Specifically, differences in positive and negative class performance across groups characterize violations of “equality of opportunity” (hardt2016equality), “equalized odds” (chouldechova2017fair; KleinbergMullainathanRaghavan(17)), and “group-level discrimination” (ArnoldDobbieHull(20)-discrimination_alg; ArnoldDobbieHull(20)-discrimination_bail). Differences in overall performance across groups characterize violations of “bounded group loss” (AgarwalEtAl(19)-FairRegression). In Online Appendix C.1, we partially identify these performance disparities under Assumption 2.1, enabling robust fairness audits.
2.3 Choice of Bounding Functions
How should the bounding functions $\underline{\delta}(\cdot;\eta)$ and $\overline{\delta}(\cdot;\eta)$ be chosen? The choice encodes assumptions about how much unselected units might differ from selected units with the same observed characteristics. In consumer lending: How much more likely are unfunded applicants to default than funded applicants with the same credit profile? In pretrial release: How much more likely are detained defendants to commit misconduct than released defendants with the same criminal history? We next show how alternative bounding functions formalize three common strategies for addressing selective labels in applied research.
2.3.1 Observed Outcome Bounds
A natural starting point is to assume that the unobserved outcome rate among unselected units is similar to the observed outcome rate among selected units with the same characteristics. For instance, in pretrial release, detained defendants with a given criminal history might be 1.5 times as likely to commit misconduct as released defendants with the same history. In consumer lending, unfunded applicants might be twice as likely to default as funded applicants with the same credit profile.
Observed outcome bounds formalize this intuition by specifying, for researcher-chosen constants $\underline{\Gamma},\overline{\Gamma}>0$ and all $x\in\mathcal{X}$,
[…]
Under this choice, the unobserved conditional probability among unselected units satisfies $\underline{\Gamma}\mu_{1}(x)\leq P(Y_{i}^{*}=1\mid D_{i}=0,X_{i}=x)\leq\overline{\Gamma}\mu_{1}(x)$ — that is, it is bounded by constant multiples of the observed conditional probability $\mu_{1}(x)$. Setting $\underline{\Gamma}=\overline{\Gamma}=1$ recovers unconfoundedness. For observed outcome bounds, the probability constraints are satisfied for all choices of $\underline{\Gamma},\overline{\Gamma}$ in a neighborhood of one under strict overlap on the decision and observed outcome probabilities being bounded away from zero and one.
This approach formalizes strategies commonly used in applied work. KLLLM(18) evaluate a constructed pretrial risk tool under varying assumptions on how much misconduct rates among detained defendants differ from released defendants (their Table 5). In consumer finance, researchers impute default rates for rejected applicants by applying multiplicative adjustments to observed default rates among accepted applicants (e.g., HandHenley(93); ZengZhao(14)). Observed outcome bounds flexibly encode such adjustments while allowing the imputed values to vary richly with covariates.
Remark 2 (Connection to Sensitivity Analysis Models in Causal Inference).
Several influential sensitivity analysis frameworks — including the marginal sensitivity model (Tan(06); ZhaoSmallBhattacharya(19); DornGuo(21)), Rosenbaum’s $\Gamma$-sensitivity model (Rosenbaum(02); Yadlowsky(21)evaluating), and partial $c$-dependence (MastenPoirier(18); MastenPoirier(20)) — place bounds on how unobservables affect selection propensities. While conceptually distinct from bounding outcome differences, these models imply observed outcome bounds with specific choices of $\underline{\Gamma},\overline{\Gamma}$ (see Online Appendix E). The connection between outcome-based bounds and selection-based sensitivity models arises through Bayes’ rule; though the observed outcome bounds we derive may not exhaust all implications of the original sensitivity model.
2.3.2 Proxy Outcome Bounds
Sometimes researchers observe a proxy outcome $\widetilde{Y}_{i}\in\{0,1\}$ for both selected and unselected units that is statistically related to the true outcome $Y_{i}^{*}$. This proxy can help discipline assumptions about unobserved confounding. For example, in consumer lending, whether an applicant defaulted on other credit products (e.g., credit cards) serves as a proxy for whether they would default on the loan of interest (BlattnerNelson(21)). In medical testing, long-term health outcomes or total medical spending serves as a proxy for whether a patient suffered an acute condition (ObermeyerEtAl(19); ChanGentzkowYu(21); MullainathanObermeyer(21)).
The key assumption is that the proxy outcome is equally informative about the true outcome for both selected and unselected units. Formally, suppose the proxy outcome satisfies, for all $x\in\mathcal{X}$,
$P(Y^{*}_{i}=\tilde{Y}_{i}\mid D_{i}=0,X_{i}=x)=P(Y^{*}_{i}=\tilde{Y}_{i}\mid D_{i}=1,X_{i}=x).$ (6)
Conditional on the covariates, the probability that the true and proxy outcome agree is the same regardless of selection status. This assumption, combined with the observed distribution of $\widetilde{Y}_{i}$ among unselected units, pins down specific bounding functions on $\delta(x)$.
Under mild restrictions satisfied by the proxy outcomes used in BlattnerNelson(21) and MullainathanObermeyer(21), the bounding functions simplify to
$\underline{\delta}(x;\eta)=1-\gamma_{1}(x)-\widetilde{\mu}_{0}(x)-\mu_{1}(x),$
$\overline{\delta}(x;\eta)=1-\gamma_{1}(x)+\widetilde{\mu}_{0}(x)-\mu_{1}(x),$
where $\gamma_{1}(x):=\mathbb{P}(Y_{i}^{*}=\tilde{Y}_{i}\mid D_{i}=1,X_{i}=x)$ is the probability the true and proxy outcome agree among selected units and $\widetilde{\mu}_{0}(x):=\mathbb{P}(\tilde{Y}_{i}=1\mid D_{i}=0,X_{i}=x)$ is the observed proxy outcome rate among unselected units. These bounding functions satisfy the probability constraints under the maintained proxy assumptions under strict overlap on the decision and observed outcome probabilities being bounded away from zero and one. Online Appendix D.1 provides the general characterization without these simplifying restrictions. To our knowledge, proxy outcome assumptions of this form have not been explored in the causal inference sensitivity analysis literature.
2.3.3 Instrumental Variable Bounds
Sometimes researchers observe an instrumental variable $Z_{i}\in\mathcal{Z}$ with finite support that generates quasi-random variation in selection decisions but does not directly affect outcomes (i.e., $Y_{i}^{*}$ independent of $Z_{i}$ given $X_{i}$). For example, in pretrial release, judges randomly assigned to cases vary in their propensity to release defendants (KLLLM(18); ArnoldDobbieHull(20)-discrimination_bail; Rambachan(21); AngelovaEtAl(22)). The day-of-week affects whether certain tests are ordered (e.g., MullainathanObermeyer(21)), and state-level variation in banking regulations affects loan approval rates (e.g., BlattnerNelson(21)).
An instrument $Z_{i}$ implies specific bounding functions through a classic result of Manski(94)-SampleSelectionBounds. For any fixed instrument value $z\in\mathcal{Z}$, the confounding function satisfies, for all $x\in\mathcal{X}$,
$\underline{\delta}_{z}(x;\eta)\leq\delta(x)\leq\overline{\delta}_{z}(x;\eta),$
where $\underline{\delta}_{z}(x)=\left(\mathbb{E}[Y_{i}D_{i}\mid X_{i}=x,Z_{i}=z]-\mu_{1}(x)\right)/\pi_{0}(x)$ and $\overline{\delta}_{z}(x):=(\pi_{0}(x,z)+\mathbb{E}[Y_{i}D_{i}\mid X_{i}=x,Z_{i}=z]-\mu_{1}(x))/\pi_{0}(x)$. Our identification and estimation results apply to these bounds for any fixed $z\in\mathcal{Z}$.
Researchers can obtain sharper bounds by aggregating across all instrument values: $\max_{z\in\mathcal{Z}}\ \underline{\delta}_{z}(x;\eta)\leq\delta(x)\leq\min_{z\in\mathcal{Z}}\ \overline{\delta}_{z}(x;\eta)$. These intersection bounds (Chernozhukov-IntersectionBounds) are tighter but involve maximum and minimum operators, creating non-smoothness that complicates estimation. Following LevisEtAl(23)-CovariateAssistedIV, we address this challenge using smooth approximations to the intersection bounds. Online Appendix D.2 discusses both approaches: bounds based on a fixed instrument value and smooth approximations to intersection bounds. [^6] LevisEtAl(23)-CovariateAssistedIV and semenova2023adaptive consider an alternative strategy for dealing with pointwise intersection bounds that invokes a “margin condition” on the separation of the maximal value of the intersection bounds. Under such a margin condition, existing results would apply to bounds on overall performance. Extending this approach to estimators for the conditional likelihood bounds or positive class performance bounds is not immediate. We leave the full analysis of margin conditions for future work.
3 Robust Design of Predictive Algorithms
In this section, we propose estimators for the bounds on the conditional likelihood $\mu^{*}(\cdot)$ under Assumption 2.1. We build on recent work studying the estimation of heterogeneous treatment effects in causal inference (e.g., KunzelEtAl(19); NieWager(20); Kennedy(22)-towardsoptimal), and our estimators use cross-fitting and doubly-robust bias corrections. Since the bounds on the conditional likelihood are an infinite-dimensional parameter, we establish finite-sample bounds on our estimator’s integrated mean square error (MSE) convergence using a novel oracle inequality for pseudo-outcome regression procedures. These convergence guarantees enable us to construct plug-in decision rules with provable regret bounds.
To simplify exposition, we develop our estimator assuming the researcher specifies observed outcome bounds (Section 2.3.1) and illustrate it with two folds. The analysis for multiple folds and cross-fitting is straightforward. We extend to proxy outcome bounds in Online Appendix D.1 and instrumental variable bounds in Online Appendix D.2.
Estimation Procedure for Observed Outcome Bounds:
Under Assumption 2.1, the unobserved conditional probability among unselected units can be decomposed as $P(Y^{*}_{i}=1|D_{i}=0,X_{i}=x)=\mu_{1}(x)+\delta(x)$, where $\mu_{1}(x)$ is the observed rate among selected units and $\delta(x)$ measures confounding. Our estimation strategy constructs pseudo-outcomes based on influence functions that combine observed outcomes among selected units with extrapolation to unselected units. The pseudo-outcomes correct for both observable selection (via inverse propensity weighting) and unobservable confounding (via the sensitivity parameters $\underline{\Gamma},\overline{\Gamma}$).
[…]
3.1 Bound on Integrated Mean Square Error Convergence
We analyze the integrated MSE convergence rate of our proposed estimators by comparing them against an infeasible oracle nonparametric regression. This oracle has access to the true nuisance functions $\eta$ and estimates the conditional likelihood bounds by regressing the true pseudo-outcomes $\phi_{\mu,i}(\eta)+\left(\overline{\Gamma}-1\right)\phi_{\pi\mu,i}(\eta)$ for the upper bound and $\phi_{\mu,i}(\eta)+\left(\underline{\Gamma}-1\right)\phi_{\pi\mu,i}(\eta)$ for the lower bound on $X_{i}$ in the second fold using the same nonparametric regression procedure specified by the researcher. We denote these oracle estimators by $\widehat{\underline{\mu}}_{oracle}(\cdot),\widehat{\overline{\mu}}_{oracle}(\cdot)$.
This oracle comparison is useful because it decomposes the integrated MSE of our proposed estimators into two sources. The first is approximation error from using a finite-sample nonparametric regression procedure to estimate the infinite-dimensional conditional likelihood bounds — this is captured by the oracle’s integrated MSE. The second is estimation error from first-stage nuisance function estimation. Our next result shows that the latter only affects our proposed estimator through a smoothed, doubly-robust remainder term.
Proposition 3.1.
Let $\widehat{\mathbb{E}}_{n}[\cdot\mid X_{i}=x]$ denote the second-stage pseudo-outcome regression estimator. Suppose $\widehat{\mathbb{E}}_{n}[\cdot\mid X_{i}=x]$ satisfies the $L_{2}(\mathbb{P})$-stability condition (Assumption A.1), and $\mathbb{P}(\epsilon\leq\hat{\pi}_{1}(X_{i})\leq 1-\epsilon)=1$ for some $\epsilon>0$. Define $\tilde{R}(x)=\widehat{\mathbb{E}}_{n}[(\pi_{1}(X_{i})-\hat{\pi}_{1}(X_{i}))(\mu_{1}(X_{i})-\hat{\mu}_{1}(X_{i}))\mid X_{i}=x]$, and $R_{oracle}^{2}=\mathbb{E}[\|\widehat{\overline{\mu}}_{oracle}(\cdot)-\overline{\mu}^{*}(\cdot)\|^{2}]$. Then,
$\|\widehat{\overline{\mu}}(\cdot)-\overline{\mu}^{*}(\cdot)\|\leq\|\widehat{\overline{\mu}}_{oracle}(\cdot)-\overline{\mu}^{*}(\cdot)\|+\epsilon^{-1}\sqrt{2}(\overline{\Gamma}-1)\|\tilde{R}(\cdot)\|+o_{\mathbb{P}}(R_{oracle})$
The analogous result holds for $\widehat{\underline{\mu}}(x)$.
Proposition 3.1 delivers two insights. First, our analysis is modular: researchers can pair choices of nuisance function estimators in the first step with any choice of nonparametric regression method (satisfying Assumption A.1) in the second step, and the integrated MSE is characterized by the oracle rate plus a doubly-robust remainder term. Second, Proposition 3.1 shows that the cost of estimating the nuisance functions in the first step only enters through the product of nuisance function errors.
To prove this result, we establish an oracle inequality on the $L_{2}(\mathbb{P})$-error of pseudo-outcome regression procedures (Lemma A.1), extending Kennedy(22)-towardsoptimal’s pointwise analysis. This result may be of independent interest. The $L_{2}(\mathbb{P})$-stability condition (Assumption A.1) on $\widehat{\mathbb{E}}_{n}[\cdot\mid X_{i}=x]$ is quite mild in practice. It is satisfied by a variety of generic linear smoothers such as linear regression, series regression, nearest neighbor matching, random forest models, and several others (Proposition A.1). Once again, in this sense, Proposition 3.1 is agnostic to both the researcher’s choice of $\widehat{\mathbb{E}}_{n}[\cdot\mid X_{i}=x]$} and nuisance function estimators.
[…]
3.2 Regret Bound for Plug-In Decision Rules
Researchers often apply a threshold rule to an estimated predictive algorithm to make decisions. For example, KLLLM(18); Rambachan(21); AngelovaEtAl(22) compare outcomes under alternative decisions rules that threshold a predictive algorithm in pretrial release. We show that plug-in decision rules based on our estimators of the conditional likelihood bounds achieve low regret, ensuring that the statistical uncertainty from estimation does not lead to substantially worse decisions than if the true conditional probabilities were known.
Since the expected payoff $U(d)$ of any decision rule $d(\cdot)\colon\mathcal{X}\rightarrow\{0,1\}$ is partially identified (Lemma 2.2), we evaluate decision rules by comparing their worst-case expected welfare (e.g., Manski(07)). An immediate consequence of Lemma 2.2 is that the optimal max-min decision rule $d^{*}(\cdot)\in\arg\max_{d(\cdot)\colon\mathcal{X}\rightarrow[0,1]}\underline{U}(d)$ is a threshold rule:
$d^{*}(X_{i})=1\{\widetilde{\mu}^{*}(X_{i})\leq u_{1,0,i}+u_{0,0,i}\},$
[…]
Proposition 3.2.
Under the same conditions as Proposition 3.1,
$R(\widehat{d})^{2}\leq 2\|\widehat{\overline{\mu}}_{oracle}(\cdot)-\overline{\mu}^{*}(\cdot)\|+2\|\widehat{\underline{\mu}}_{oracle}(\cdot)-\underline{\mu}^{*}(\cdot)\|+4\epsilon^{-1}\sqrt{2}\|\tilde{R}(x)\|+o_{\mathbb{P}}(R_{oracle}).$
The proof proceeds in two steps. First, we show that the regret of the plug-in decision rule is bounded by the integrated MSE of our estimators for the conditional probability bounds — this follows because the optimal decision rule is a threshold rule, and threshold rules are sensitive to estimation errors measured in $L_{2}(\mathbb{P})$ norm. Second, we apply Proposition 3.1 to decompose this integrated MSE into the oracle rate plus the doubly-robust remainder term. Consequently, worst-case regret converges to zero whenever the oracle’s integrated MSE converges to zero. Plug-in decision rules based on our nonparametric estimators are therefore robust: they achieve nearly optimal welfare despite estimation uncertainty.
4 Robust Evaluation of Predictive Algorithms
In this section, we construct estimators for the bounds on the performance of a given predictive algorithm $s(\cdot)$ under Assumption 2.1. This setting arises empirically in two scenarios: (i) when researchers evaluate an externally provided algorithm (e.g., a proprietary credit score or a commercial risk assessment tool), and (ii) when researchers conduct hold-out evaluation of an algorithm trained on separate data. In both cases, the algorithm $s(\cdot)$ is taken as fixed, and the goal is to assess its performance under varying assumptions about unobserved confounding. Our estimators are consistent for the bounds on the identified set as sample size grows large, and we characterize their convergence rates in terms of errors in the first-step estimation of nuisance parameters.
4.1 Bounds on Overall Performance
We first construct estimators for the bounds on overall performance $\mbox{perf}(s;\beta)$. To simplify exposition, we develop estimators for observed outcome bounds (Section 2.3.1). We extend to proxy outcome bounds and instrumental variable bounds in Online Appendices D.1-D.2 respectively. Under observed outcome bounds, Lemma 2.1 implies that the upper bound can be written as
$\overline{\mbox{perf}}(s;\beta):=\mathbb{E}\left[\beta_{0,i}+\beta_{1,i}\mu_{1}(X_{i})+\beta_{1,i}\left(1\{\beta_{1,i}>0\}\left(\overline{\Gamma}-1\right)+1\{\beta_{1,i}\leq 0\}(\underline{\Gamma}-1)\right)\pi_{0}(X_{i})\mu_{1}(X_{i})\right].$
The lower bound $\underline{\mbox{perf}}(s;\beta)$ has an analogous expression. Both are linear functionals of known functions of the data and identified nuisance parameters $\eta=(\mu_{1},\pi_{0})$. This linearity enables us to construct debiased estimators using standard arguments based on influence functions and cross-fitting (e.g., ChernozhukovEtAl(18)-DML; ChernozhukovEtAl(22)-LocallRobust; Kennedy(22)-IFreview).
Estimation Procedure for Observed Outcome Bounds:
For completeness, we sketch the construction of our estimators based on $K$-fold cross-fitting. We randomly split the data into $K$ disjoint folds. For each fold $k$, we estimate the nuisance functions $\widehat{\eta}_{-k}$ using all observations not in the $k$-th fold and construct
$\overline{\mbox{perf}}_{i}(\hat{\eta}_{-k}):=\beta_{0,i}+\beta_{1,i}\phi_{\mu,i}(\hat{\eta}_{-k})+\beta_{1,i}\left(1\{\beta_{1,i}>0\}\left(\overline{\Gamma}-1\right)+1\{\beta_{1,i}\leq 0\}(\underline{\Gamma}-1)\right)\phi_{\pi\mu,i}(\hat{\eta}_{-k})$
for each observation $i$ in the $k$-th fold, where $\phi_{\mu,i}(\eta)$ and $\phi_{\pi\mu,i}(\eta)$ are defined as before. We then average over all observations $\widehat{\overline{\mbox{perf}}}(s;\beta):=\mathbb{E}_{n}\left[\overline{\mbox{perf}}_{i}(\hat{\eta}_{-K_{i}})\right],$ where $K_{i}$ is the fold containing observation $i$. Our estimator for the lower bound $\widehat{\underline{\mbox{perf}}}(s;\beta)$ is defined analogously.
Our next result characterizes the convergence rate and asymptotic distribution of our estimators for the bounds on overall performance under observed outcome bounds.
Proposition 4.1.
Define $R_{1,n}^{k}:=\|\hat{\mu}_{1,-k}(\cdot)-\mu_{1}(\cdot)\|\|\hat{\pi}_{1,-k}(\cdot)-\pi_{1}(\cdot)\|$ for each fold $k$. Assume (i) there exists $M<\infty$ such that $\|\beta_{1}(\cdot)\|\leq M$; (ii) $\mathbb{P}(\pi_{1}(X_{i})\geq\delta)=1$ for some $\delta>0$, (iii) there exists $\epsilon>0$ such that $\mathbb{P}(\hat{\pi}_{1,-k}(X_{i})\geq\epsilon)=1$ for each fold $k$, and (iv) $\|\hat{\mu}_{1,-k}(\cdot)-\mu_{1}(\cdot)\|=o_{P}(1)$ and $\|\hat{\pi}_{1,-k}(\cdot)-\pi_{1}(\cdot)\|=o_{P}(1)$ for each fold $k$. Then,
$\left|\widehat{\overline{\mbox{perf}}}(s;\beta)-\overline{\mbox{perf}}(s;\beta)\right|=O_{\mathbb{P}}\left(1/\sqrt{n}+\sum_{k=1}^{K}R_{1,n}^{k}\right),$
and the analogous result holds for $\widehat{\underline{\mbox{perf}}}(s;\beta)$. If further $R_{1,n}^{k}=o_{\mathbb{P}}(1/\sqrt{n})$ for each fold $k$, then
$\sqrt{n}\left(\begin{pmatrix}\widehat{\overline{\mbox{perf}}}(s;\beta)\\ \widehat{\underline{\mbox{perf}}}(s;\beta)\end{pmatrix}-\begin{pmatrix}\overline{\mbox{perf}}(s;\beta)\\ \underline{\mbox{perf}}(s;\beta)\end{pmatrix}\right)\xrightarrow{d}N\left(0,\Sigma\right)$
for $\Sigma=Cov\left((\overline{\mbox{perf}}_{i}(\eta),\underline{\mbox{perf}}_{i}(\eta))^{\prime}\right)$.
The estimators converge at rate $O_{\mathbb{P}}(1/\sqrt{n})$ plus a term that captures the nuisance estimation error. Nuisance errors enter multiplicatively — through the product of propensity score and outcome regression errors — rather than additively. The rate condition required for the estimators to be jointly asymptotically normal is satisfied under the standard condition that all nuisance function estimators converge at a rate faster than $O_{\mathbb{P}}(n^{-1/4})$. This is the familiar condition from debiased machine learning, satisfied by a wide range of nonparametric regression and modern machine learning methods.
Online Appendix C.2 constructs a consistent estimator of the asymptotic covariance matrix in Proposition 4.1. This implies that confidence intervals for the identified set $\mathcal{H}(\mbox{perf}(s;\beta))$ can be constructed using standard methods (e.g., ImbensManski(04); Stoye(09)--PICI).
4.2 Bounds on Positive Class Performance
We next estimate bounds on positive class performance $\mbox{perf}_{+}(s;\beta)$. Unlike overall performance, this is a ratio of expectations, and the bounds are given by the optimal values of linear fractional programs as shown in Lemma 2.1. This creates additional challenges in estimation. In this section, we will develop estimators for these bounds and establish their convergence properties.
4.2.1 Estimation via Linear Fractional Programming:
[…]
$\widehat{\overline{\mbox{perf}}}_{+}(s;\beta)=\max_{\tilde{\delta}\in\widehat{\Delta}_{n}}\frac{\mathbb{E}_{n}\left[\beta_{0,i}\phi_{\mu,i}(\hat{\eta})+\beta_{0,i}\tilde{\delta}_{i}\right]}{\mathbb{E}_{n}\left[\phi_{\mu,i}(\hat{\eta})+\tilde{\delta}_{i}\right]},$ (7)
[…]
We provide an error decomposition for the positive class performance bounds. This result characterizes how errors in estimating the nuisance parameters $(\mu_{1}(\cdot),\pi_{1}(\cdot))$ and the bounding functions $(\underline{\delta}(\cdot),\overline{\delta}(\cdot))$ propagate to the final estimator.
Proposition 4.2.
Define $R_{1,n}=\|\hat{\mu}_{1}(\cdot)-\mu_{1}(\cdot)\|\|\hat{\pi}_{1}(\cdot)-\pi_{1}(\cdot)\|$. Assume (i) there exists $M<\infty$ such that $\|\beta_{0}(\cdot)\|\leq M$; (ii) there exists $\delta>0$ such that $\mathbb{P}(\pi_{1}(X_{i})\geq\delta)=1$; (iii) there exists $\epsilon>0$ such that $\mathbb{P}(\hat{\pi}_{1}(X_{i})\geq\epsilon)=1$; (iv) $\|\hat{\mu}_{1}(\cdot)-\mu_{1}(\cdot)\|=o_{P}(1)$ and $\|\hat{\pi}_{1}(\cdot)-\pi_{1}(\cdot)\|=o_{P}(1)$; and (v) $\widehat{\Delta}_{n}$ is non-empty with probability one. Then,
$\left\|\widehat{\overline{\mbox{perf}}}_{+}(s;\beta)-\overline{\mbox{perf}}_{+}(s;\beta)\right\|=O_{\mathbb{P}}\left(1/\sqrt{n}+R_{1,n}+\sqrt{\mathbb{E}_{n}[(\widehat{\underline{\delta}}(X_{i})-\underline{\delta}(X_{i}))^{2}]}+\sqrt{\mathbb{E}_{n}[(\widehat{\overline{\delta}}(X_{i})-\overline{\delta}(X_{i}))^{2}]}\right).$
Proposition 4.2 reveals the structure of how estimation errors affect our estimator. The first term $O_{\mathbb{P}}(1/\sqrt{n})$ reflects sampling error that would arise even if the nuisance functions $(\mu_{1}(\cdot),\pi_{1}(\cdot))$ and the bounding functions $(\underline{\delta}(\cdot),\overline{\delta}(\cdot))$ were known. The second term $O_{\mathbb{P}}(R_{1,n})$ reflects errors from estimating the nuisances $(\mu_{1}(\cdot),\pi_{1}(\cdot))$ that appear in the numerator and denominator of the linear fractional program. These enter through a doubly robust product, and this behavior is inherited from the influence function structure underlying the numerator and denominator. The final term captures errors from estimating the bounding functions $(\underline{\delta}(\cdot),\overline{\delta}(\cdot))$ that define the constraint set $\Delta$. This error enters through the root mean square error. This error can be controlled using standard nonparametric regression techniques, and researchers can trade off bias and variance in familiar ways.
For this reason, Proposition 4.2 establishes a “partial” double robustness: given estimated constraints, our estimator is doubly robust to errors in the objective nuisances $(\mu_{1}(\cdot),\pi_{1}(\cdot))$. Constraint estimation errors enter additively through RMSE. Achieving full double robustness — where constraint estimation errors would also enter with product structure — remains an open problem, which we discuss in Remark 4. Remark 5 describes a simple alternative approach that obtains fully doubly-robust estimators for non-sharp bounds on positive class performance.
Proposition 4.2 and our estimator are practically valuable. Researchers can use state-of-the-art machine learning methods for estimating the objective nuisances $(\mu_{1}(\cdot),\pi_{1}(\cdot))$, while separately controlling the constraint estimation error through their choice of estimator for the bounding functions. The error decomposition applies to any estimator of the bounding functions with favorable $L_{2}(\mathbb{P})$ properties, and we discuss alternative strategies and their tradeoffs in Section 4.2.2. This modularity is a key practical advantage of our framework.
Remark 3 (Connection to existing work).
Our analysis extends existing work on linear fractional programs, including AronowLee(2013), MiratrixWagerZubizaretta(18), KallusZhou(18)-confoundingrobust, and ZhaoSmallBhattacharya(19). These papers analyze linear fractional programs that arise when bounding population means or treatment effects under alternative sensitivity analysis models. Our analysis advances in two key respects: (i) we establish partial double robustness for objective nuisances – a property not achieved in prior work; and (ii) we provide a unified framework that accommodates multiple bounding strategies (observed outcome bounds, proxy outcome bounds, instrumental variable bounds) through a modular approach to constraint estimation.
Remark 4 (Paths toward full double robustness).
The proof of Proposition 4.2 reveals that the optimizer in the linear fractional program can be characterized as having a threshold structure in $\beta_{0,i}$. The twist is that the cutoff determining the threshold is itself not fixed but rather a functional of the estimated constraints. Consequently, estimation error in the constraints affects not only which functions are feasible but also where the optimal threshold lies. While related, recent work on conditional linear programs (benmichael2025partialidentificationconditionallinear) and aggregated intersection bounds (semenova2023adaptive) are not directly applicable. While the linear fractional program can be reduced to a linear program through the Charnes-Cooper transformation, the transformation uses a global normalization that couples all observations together, whereas existing work exploits conditional separability across $X$-cells. We leave full resolution of this question for future work and view our partial double robustness result as an important first step.
Remark 5 (Estimation of non-sharp bounds).
Our estimator targets sharp bounds on positive class performance, but an alternative approach that restores double robustness at the cost of sharpness is possible. The idea is straightforward: separately estimate the numerator and the denominator using overall performance estimators (Section 4.1) and then combine via the delta method. Concretely, define
$\widetilde{\overline{\mbox{perf}}}_{+}(s;\beta)=\frac{\sup_{\delta\in\Delta}\mathbb{E}[\beta_{0,i}\mu_{1}(X_{i})+\beta_{0,i}\pi_{0}(X_{i})\delta(X_{i})]}{\inf_{\delta\in\Delta}\mathbb{E}[\mu_{1}(X_{i})+\pi_{0}(X_{i})\delta(X_{i})]}.$
Clearly, $\widetilde{\overline{\mbox{perf}}}_{+}(s;\beta)\geq\overline{\mbox{perf}}_{+}(s;\beta)$. The numerator and the denominator can be separately estimated using the methods in Section 4.1. This delivers: (i) full double robustness with respect to all nuisances, and (ii) standard inference with straightforward variance estimation. The cost is that we target non-sharp bounds. Whether this trade-off is attractive depends on the application.
Remark 6 (Interpretation of non-empty constraint sets).
Proposition 4.2 assumes the estimated constraint set $\widehat{\Delta}_{n}$ is non-empty. If finite-sample variability leads to $\widehat{\underline{\delta}}(x)>\widehat{\overline{\delta}}(x)$ for some $x\in\mathcal{X}$, researchers can either: (i) clip the bounds to enforce feasibility, which might introduce bias captured by the root mean square terms in Proposition 4.2; or (ii) restrict attention to the feasible covariate region $\tilde{X}_{n}=\{x:\widehat{\underline{\delta}}(x)\leq\widehat{\overline{\delta}}(x)\}$, thus redefining the estimand. The latter approach is analogous to trimming observations with extreme estimated propensity scores in causal inference, though this may complicate analysis (e.g., HotzEtAl(09); Sasaki_Ura_2022).
4.2.2 Estimation of the Bounding Functions
The error decomposition in Proposition 4.2 applies to any estimator of the bounding functions with favorable mean square error properties. This modularity gives researchers flexibility. We discuss two natural strategies and their relative merits, focusing on observed outcome bounds for concreteness.
The simplest approach directly substitutes nuisance estimates into the bounding function formulas. For observed outcome bounds, this means $\widehat{\underline{\delta}}(x)=(\underline{\Gamma}-1)\widehat{\mu}_{1}(x)$ and $\overline{\delta}(x)=(\overline{\Gamma}-1)\widehat{\mu}_{1}(x)$. The mean square error properties of the estimated bounding functions are then inherited from those of $\widehat{\mu}_{1}(x)$. This approach requires no additional estimation beyond the first-stage nuisances. An alternative approach is to construct influence-function-based pseudo-outcomes and regress them on covariates using the full sample (i.e., our pseudo-outcome regression procedure in Section 3). We develop this approach below as it may offer bias-variance advantages in practice. Ultimately, both strategies are valid under our framework.
To understand why pseudo-outcome regression may be advantageous, consider that $\widehat{\mu}_{1}(x)$ is typically constructed using only observations with $D_{i}=1$. Consequently, standard regression procedures target the mean square error conditional on $D_{i}=1$ However, the error bound in Proposition 4.2 depends on the unconditional mean square error. These two objectives could lead to different bias-variance tradeoffs when the selected population ($D_{i}=1$) is small relative to the full population or there is covariate shift between the selected and unselected populations. In such settings, an estimator that directly targets the unconditional MSE may achieve better performance. Pseudo-outcome regression accomplishes this by constructing unbiased pseudo-outcomes using observations from both the selected and unselected populations, then regressing these pseudo-outcomes on the full sample.
Estimation Procedure for Observed Outcome Bounds:
To illustrate how Proposition 4.2 may be applied, we estimate bounding functions using our pseudo-outcome regression procedure. This illustrates how our partial double robustness result (Proposition 4.2) can be combined with analyses of nonparametric regression estimators — in particular, leveraging our oracle inequality for pseudo-outcome regression (Appendix A).
[…]
Corollary 4.1.
Under the same conditions as Proposition 4.2, if further $\widehat{\mathbb{E}}_{n}[\cdot\mid X_{i}=x]$ satisfies the $L_{2}(\mathbb{P})$-stability condition (Assumption A.1), then
$\sqrt{\mathbb{E}_{n}[(\widehat{\delta}(X_{i})-\mu_{1}(X_{i}))^{2}]}=O_{\mathbb{P}}\left(1/\sqrt{n}+\epsilon^{-1}\|\tilde{R}(\cdot)\|+R_{oracle}\right),$
where $\tilde{R}(x)=\widehat{\mathbb{E}}_{n}[(\pi_{1}(X_{i})-\hat{\pi}_{1}(X_{i}))(\mu_{1}(X_{i})-\hat{\mu}_{1}(X_{i}))\mid X_{i}=x]$, and now $R_{oracle}^{2}=\mathbb{E}[\|\widehat{\delta}_{oracle}(\cdot)-\mu_{1}(\cdot)\|^{2}]$ for the oracle pseudo-outcome regression procedure.
5 Empirical Application to Credit Risk Scores
We validate the finite-sample performance of our estimators through Monte Carlo simulations in Appendix F. We now apply our framework to credit risk prediction, where financial institutions use scores $s(\cdot)$ to predict default risk. We observe data on past loan applications but only observe defaults $Y_{i}=D_{i}Y_{i}^{*}$ for funded applications ($D_{i}=1$). It is implausible to assume unconfounded funding decisions. Applicants who accept loan offers may differ systematically from those who decline in ways that affect their default risk. [^7] Lending institutions face two distinct inference problems. First, ”policy rejects” with $\mathbb{P}(D_{i}=1\mid X_{i})=0$ (i.e., overlap violations) due to underwriting rules require extrapolation-based methods that are beyond our scope. Second, even among fundable applicants with $\mathbb{P}(D_{i}=1\mid X_{i})>0$, those accepting versus declining offers may differ in unobserved ways — the selective labels problem we address. Extending our framework to handle policy rejects is an important direction for future research.
We analyze 372,346 personal loan applications submitted to Commonwealth Bank of Australia from July 2017 to July 2019. These loans are repaid in monthly installments and used to purchase vehicles, refinance debt, or cover major expenses (CostonEtAl(21)-Rashomon). Loan amounts ranged up to AU\$50,000 (median AU\$10,000) with a median offered interest rate of 14.9% per annum. We observe rich application-level covariates including reported income, occupation, and credit history at CommBank. We only observe whether an applicant defaulted within 5 months $Y_{i}^{*}\in\{0,1\}$ if the application was funded $D_{i}=1$. Approximately one-third of applications were funded, with a 2.0% default rate among funded loans.
We consider two exercises. First, we construct credit risk scores robust to varying confounding assumptions. The resulting algorithms produce substantially different risk rankings that could affect credit access. Second, following recent work in consumer finance (BlattnerNelson(21); FusterEtAl(22)predictably; DiMaggioEtAl(22)-Fintech), we evaluate how a benchmark score’s accuracy varies across income groups. While an analysis ignoring confounding suggests better performance for higher-income applicants, this disparity disappears under plausible confounding levels.
5.1 Bounding Individual Default Risk
We construct robust credit risk scores using our estimator from Section 3, exploring how varying confounding assumptions affect individual default risk predictions. We estimate the upper bound on the conditional default probability with observed outcome bounds, $\overline{\mu}(\cdot)$. We split applications into two folds: the first estimates $\pi_{1}(\cdot)$, $\mu_{1}(\cdot)$ using random forests; the second regresses estimated pseudo-outcomes on application-level covariates $X_{i}$ using cross-validated Lasso regression. We set $\underline{\Gamma}=1$ and vary $\overline{\Gamma}\in\{1,2,3\}$.
We compare our robust estimator (with $\underline{\Gamma}=1,\overline{\Gamma}=2$) against a benchmark score trained only on funded applications. Figure 1’s left panel shows the joint distribution of their predictions. For each benchmark decile, the heatmap displays the percentage of applications falling into each decile of our robust predictions. Applications exhibit substantial reranking: among those in the benchmark’s 5th decile, 18.7% shift to extreme deciles (1-3 or 8-10) under the robust score. These differences meaningfully affect credit access. For instance, if only the least risky third of applications receive funding, 10.1% would have their decision reversed when comparing benchmark versus robust scores.
Figure 1’s right panel examines how estimated coefficients vary with $\overline{\Gamma}$ (variable definitions in Online Appendix Table A3). Confounding assumptions affect different covariates heterogeneously. Some coefficients—total net income and credit bureau score—remain zero across all $\overline{\Gamma}$ values. Others, like the number of recent credit card applications, receive zero weight in the benchmark but non-zero weight for all $\overline{\Gamma}$¿1. Still others, such as occupation type and maximum recent delinquency, exhibit large magnitude changes as $\overline{\Gamma}$ varies, highlighting how confounding assumptions shape which applicant characteristics the algorithm prioritizes.
5.2 Bounding the Predictive Performance of a Credit Risk Score
We now evaluate an existing credit risk score’s performance. We split applications into training and evaluation data. On the training data, we construct a benchmark score predicting 5-month default risk among funded applications using random forests. On the evaluation data, we analyze mean square error and ROC curves, following recent mortgage lending studies (BlattnerNelson(21); FusterEtAl(22)predictably). Using observed outcome bounds with $\underline{\Gamma}=1$, we assess how the benchmark’s estimated performance varies as we increase $\overline{\Gamma}$, allowing unfunded applicants to have progressively higher default rates than funded applicants with similar characteristics.
Figure 2(a) shows how estimated MSE bounds vary with $\overline{\Gamma}$. When $\overline{\Gamma}=1$ (the unconfounded case), both bounds equal the benchmark’s MSE on funded applications (dashed red line; $0.143$). As $\overline{\Gamma}$ increases, allowing unfunded applicants higher default rates, bounds widen but remain informative. At $\overline{\Gamma}=2$ — permitting unfunded applicants to default at twice the rate of funded applicants with similar characteristics — the estimated bounds are $[0.138,0.182]$. Figure 2(b) examines how estimated ROC curves vary with $\overline{\Gamma}$. We summarize ROC curves using Area Under the Curve (AUC), computed via the trapezoidal rule. Under unconfoundedness ($\overline{\Gamma}=1$), the benchmark’s AUC is $0.637$. At $\overline{\Gamma}=1.5$, AUC bounds are $[0.575,0.705]$. As with MSE, bounds widen with $\overline{\Gamma}$ but remain informative at empirically plausible confounding levels.
We finally investigate the benchmark credit risk score’s predictive disparities across income groups. We define $G_{i}\in\{0,1\}$ to be whether an application is below or above the median personal income of all submitted applications. We examine how the benchmark credit risk score’s mean square error and ROC curve vary across these income groups, thereby studying whether it is a “noisier” signal of 5-month default risk for lower income applications, as in BlattnerNelson(21); FusterEtAl(22)predictably.
Figure 3(a) compares the mean square error of the benchmark credit risk across income groups. Interestingly, for $\overline{\Gamma}=1$, the mean square error of the benchmark credit risk score is significantly larger on applications below the median income (0.169) than those above it (0.136). This difference persists as we vary assumptions on unobserved confounding. The upper bound on the mean square error for applications above the median income is only larger than the lower bound on the mean square error for applications below the median income for values $\overline{\Gamma}\geq 1.69$. Ruling out this mean square error disparity across income groups would require that unfunded applications are at least 1.69 times as likely to default within 5 months as funded applications conditional on covariates. [^8] Rather than examining disparity bounds across a range of $\overline{\Gamma}$ values, researchers could compute the “identification breakdown point” $\overline{\Gamma}^{*}$: the minimal value of $\overline{\Gamma}$ (holding $\underline{\Gamma}=1$) such that $0\in H(\text{disp}(s;\beta))$. A natural estimator is $\widehat{\overline{\Gamma}}^{*}$, the minimal $\overline{\Gamma}$ such that zero is contained in the estimated $(1-\alpha)$ confidence set $\widehat{H}(\text{disp}(s;\beta))$. Since our confidence sets have valid coverage by the results in Section 4.1 and Appendix C.1, $P(\widehat{\overline{\Gamma}}^{*}\geq\overline{\Gamma}^{*})\geq 1-\alpha$. Figure 3(b) also compares the ROC curves of the benchmark credit risk score across income groups. Intriguingly, there exists effectively no disparity across income groups based on the ROC curves. When $\overline{\Gamma}=1$, the implied AUC for applications above the median income is $0.603$, whereas the implied AUC is $0.616$ for applications below the median income. Any disparity in terms of ROC curves is already ruled out at $\overline{\Gamma}=1.25$ as the bounds on the AUC for applications above the median income are $[0.567,0.653]$ and $[0.581,0.654]$ for applications below the median income.
6 Conclusion
We developed a framework for robust design and evaluation of predictive algorithms under unobserved confounding. Our approach nests popular empirical strategies — observed outcome bounds, proxy outcomes, and instrumental variables — by bounding outcome differences between selected and unselected units (Assumption 2.1). We provide estimators that work well across bounding strategies and performance measures: conditional likelihoods for algorithm design, overall metrics like MSE and calibration, and class-specific measures like TPR and FPR. This breadth makes our framework valuable across domains where predictive algorithms face selective labels—pretrial release, consumer lending, medical diagnosis, hiring, and child welfare.
Several open questions present promising research directions. First, achieving full double robustness for positive class performance bounds remains open, though our technical results suggest possible paths forward. Second, uniform inference for breakdown frontiers would let researchers formally test how strong confounding must be to overturn conclusions (MastenPoirier(20)). Third, we evaluate fixed algorithms; extending our results to settings where the same data is used to train and evaluate an algorithm is an important direction. Finally, extending to multi-valued or continuous outcomes would broaden applicability, though defining class-specific performance in these settings requires care.
As algorithms increasingly drive high-stakes decisions in credit, criminal justice, and healthcare, robust evaluation under realistic confounding becomes essential. Our framework and these open questions offer pathways to strengthen algorithmic accountability where the stakes are highest.
Main Figures
(a) Benchmark risk score vs. robust risk score.
(b) Coefficients in risk score as $\overline{\Gamma}$ varies.
Figure 1: Estimated personal loan credit risk scores as assumptions on unobserved confounding vary.
(a) Mean square error
(b) ROC curve
Figure 2: Bounds on mean square error and ROC curve of benchmark risk score as we vary the assumption on unobserved confounding.
(a) Mean square error
(b) ROC curve
Figure 3: Bounds on mean square error and ROC curve of benchmark credit risk score across income groups as we vary the assumption on unobserved confounding.
Robust Design and Evaluation of Predictive Algorithms
under Unobserved Confounding
Online Appendix
Ashesh Rambachan     Amanda Coston     Edward H. Kennedy
This online appendix contains additional results for “Robust Design and Evaluation of Predictive Algorithms under Unobserved Confounding” by Ashesh Rambachan, Amanda Coston, and Edward Kennedy. Section A contains an oracle inequality for the integrated mean square convergence rate of the pseudo-outcome regression procedures. Section B contains the proofs of all theoretical results presented in the main text. Section C contains additional theoretical results discussed in the main text. Section D extends our estimation results to alternative bounding functions, such as proxy outcome bounds and instrumental variable bounds. Section E discusses connections between existing sensitivity analysis frameworks in causal inference and observed outcome bounds in our framework. Section F contains Monte Carlo simulations, and Section G contains additional empirical results.
Appendix A Oracle Inequality for Pseudo-Outcome Regression
We provide an oracle inequality on the $L_{2}(\mathbb{P})$-error of nonparametric regression with estimated pseudo-outcomes. This generalizes the pointwise analysis in Kennedy(22)-towardsoptimal. In Section 3 of the main text, we apply this oracle inequality to analyze our estimators for the conditional probability bounds.
We first state a $L_{2}(\mathbb{P})$-stability condition on the second-stage nonparametric regression estimator. We then show that the $L_{2}(\mathbb{P})$-stability condition is satisfied by a variety of generic linear smoothers such as linear regression, series regression, nearest neighbor matching, random forest model and several others.
Assumption A.1.
Suppose $\mathcal{O}_{train}=\left(V_{01},\ldots,V_{0n}\right)$ and $\mathcal{O}_{test}=\left(V_{1},\ldots,V_{n}\right)$ are independent train and test sets with covariates $X_{i}\subseteq V_{i}$. Let (i) $\hat{f}(w)=\hat{f}(w;\mathcal{O}_{train})$ be an estimate of a function $f(w)$ using the training data; (ii) $\hat{b}(x)=\mathbb{E}[\hat{f}(V_{i})-f(V_{i})\mid X_{i}=x,\mathcal{O}_{train}]$ be the conditional bias of the estimator $\hat{f}$; and (iii) $\hat{\mathbb{E}}_{n}[V_{i}\mid X_{i}=x]$ be a generic nonparametric regression estimator that regresses outcomes $(V_{1},\ldots,V_{n})$ on covariates $(X_{1},\ldots,X_{n})$ in the test sample $\mathcal{O}_{test}$.
The regression estimator $\hat{\mathbb{E}}_{n}[\cdot]$ is $L_{2}(\mathbb{P})$-stable with respect to distance metric $d(\cdot,\cdot)$ if
$\frac{\int\Big[\widehat{\mathbb{E}}_{n}\{\widehat{f}(V_{i})\mid X_{i}=x\}-\widehat{\mathbb{E}}_{n}\{{f}(V_{i})\mid X_{i}=x\}-\widehat{\mathbb{E}}_{n}\{\widehat{b}(X_{i})\mid X_{i}=x\}\Big]^{2}\ d\mathbb{P}(x)}{\mathbb{E}\left(\int\Big[\widehat{\mathbb{E}}_{n}\{f(V_{i})\mid X_{i}=x\}-\mathbb{E}\{f(V_{i})\mid X_{i}=x\}\Big]^{2}d\mathbb{P}(x)\right)}\ \xrightarrow{p}0$ (8)
whenever $d(\widehat{f},f)\xrightarrow{p}0$.
Proposition A.1.
Linear smoothers of the form $\widehat{\mathbb{E}}_{n}\{\widehat{f}(V_{i})\mid X_{i}=x\}=\sum_{i}w_{i}(x;X^{n})\widehat{f}(V_{i})$ are $L_{2}(\mathbb{P})$-stable with respect to distance
$d(\widehat{f},f)=\|\widehat{f}-f\|_{w^{2}}\equiv\sum_{i=1}^{n}\left\{\frac{\|w_{i}(\cdot;X^{n})\|^{2}}{\sum_{j}\|w_{j}(\cdot;X^{n})\|^{2}}\right\}\int\left\{\widehat{f}(v)-f(v)\right\}^{2}\ d\mathbb{P}(v\mid X_{i}),$
whenever $1/\|\sigma\|_{w^{2}}=O_{\mathbb{P}}(1)$ for $\sigma(x)^{2}=Var\{f(V_{i})\mid X_{i}=x\}$.
Proof.
The proof follows an analogous argument as Theorem 1 of Kennedy(22)-towardsoptimal. Letting $T_{n}(x)=\widehat{m}(x)-\widetilde{m}(x)-\widehat{\mathbb{E}}_{n}\{\widehat{b}(X)\mid X=x\}$ denote the numerator of the left-hand side of (8), and $R_{n}^{2}=\mathbb{E}[\|\widetilde{m}-m\|]^{2}$ denote the oracle error, we will show that
$\|T_{n}\|=O_{\mathbb{P}}\left(\frac{\|\widehat{f}-f\|_{w^{2}}}{\|\sigma\|_{w^{2}}}R_{n}\right)$
which yields the result when $1/\|\sigma\|_{w^{2}}=O_{\mathbb{P}}(1)$. First, note that for linear smoothers
$T_{n}(x)=\widehat{\mathbb{E}}_{n}\{\widehat{f}(V_{i})-f(V_{i})-\widehat{b}(X_{i})\mid X_{i}=x\}=\sum_{i=1}^{n}w_{i}(x;X^{n})\Big\{\widehat{f}(V_{i})-f(V_{i})-\widehat{b}(X_{i})\Big\}$
and this term has mean zero since
$\mathbb{E}\Big\{\widehat{f}(V_{i})-f(V_{i})-\widehat{b}(X_{i})\mid\mathcal{O}_{train},X^{n}\Big\}=\mathbb{E}\Big\{\widehat{f}(V_{i})-f(V_{i})-\widehat{b}(X_{i})\mid\mathcal{O}_{train},X_{i}\Big\}=0$
by definition of $\widehat{b}$ and iterated expectation. Therefore,
$\mathbb{E}$ $(T_{n}(x)^{2}\mid\mathcal{O}_{train},X^{n})=Var\left[\sum_{i=1}^{n}w_{i}(x;X^{n})\Big\{\widehat{f}(V_{i})-f(V_{i})-\widehat{b}(X_{i})\Big\}\Bigm|\mathcal{O}_{train},X^{n}\right]$
$=\sum_{i=1}^{n}w_{i}(x;X^{n})^{2}\ Var\Big\{\widehat{f}(V_{i})-f(V_{i})\mid\mathcal{O}_{train},X_{i}\Big\}$ (9)
where the second line follows since $\widehat{f}(V_{i})-f(V_{i})$ are independent given the training data. Thus
$\mathbb{E}\left(\|T_{n}\|^{2}\Bigm|\mathcal{O}_{train},X^{n}\right)$ $=\int\sum_{i=1}^{n}w_{i}(x;X^{n})^{2}\ Var\Big\{\widehat{f}(V_{i})-f(V_{i})\mid\mathcal{O}_{train},X_{i}\Big\}\ d\mathbb{P}(x)$
$=\sum_{i=1}^{n}\|w_{i}(\cdot;X^{n})\|^{2}Var\Big\{\widehat{f}(V_{i})-f(V_{i})\mid\mathcal{O}_{train},X_{i}\Big\}$
$\leq\sum_{i=1}^{n}\|w_{i}(\cdot;X^{n})\|^{2}\int\Big\{\widehat{f}(v)-f(v)\Big\}^{2}\ d\mathbb{P}(v\mid X_{i})$
$=\|\widehat{f}-f\|_{w^{2}}\sum_{j}\|w_{j}(\cdot;X^{n})\|^{2}$
where the third line follows since $Var(\widehat{f}-f\mid\mathcal{O}_{train},X_{i})\leq\mathbb{E}\{(\widehat{f}-f)^{2}\mid\mathcal{O}_{train},X_{i}\}$, and the fourth by definition of $\|\cdot\|_{w^{2}}$.
Further note that $R_{n}^{2}$ equals
$\mathbb{E}$ $[\|\widetilde{m}-m\|]^{2}=\mathbb{E}\left(\int\left[\sum_{i=1}^{n}w_{i}(x;X^{n})\Big\{f(V_{i})-m(X_{i})\Big\}+\sum_{i=1}^{n}w_{i}(x;X^{n})m(X_{i})-m(x)\right]^{2}d\mathbb{P}(x)\right)$
$=\mathbb{E}\left(\int\left[\sum_{i=1}^{n}w_{i}(x;X^{n})\Big\{f(V_{i})-m(X_{i})\Big\}\right]^{2}d\mathbb{P}(x)\right)+\mathbb{E}\left[\int\left\{\sum_{i=1}^{n}w_{i}(x;X^{n})m(X_{i})-m(x)\right\}^{2}d\mathbb{P}(x)\right]$
$=\mathbb{E}\left\{\int\sum_{i=1}^{n}w_{i}(x;X^{n})^{2}\ \sigma(X_{i})^{2}\ d\mathbb{P}(x)\right\}+\mathbb{E}\left[\int\left\{\sum_{i=1}^{n}w_{i}(x;X^{n})m(X_{i})-m(x)\right\}^{2}d\mathbb{P}(x)\right]$
$\geq\mathbb{E}\sum_{i=1}^{n}\|w_{i}(\cdot;X^{n})\|^{2}\sigma(X_{i})^{2}=\mathbb{E}\left\{\|\sigma\|_{w^{2}}^{2}\sum_{j}\|w_{j}(\cdot;X^{n})\|^{2}\right\}$ (10)
where the second and third lines follow from iterated expectation and independence of the samples, and the fourth by definition of $\|\cdot\|_{w^{2}}$ (and since the integrated squared bias term from the previous line is non-negative). Therefore,
$\mathbb{P}\left\{\frac{\|\sigma\|_{w^{2}}\|T_{n}\|}{\|\widehat{f}-f\|_{w^{2}}R_{n}}\geq t\right\}$ $=\mathbb{E}\left[\mathbb{P}\left\{\frac{\|\sigma\|_{w^{2}}\|T_{n}\|}{\|\widehat{f}-f\|_{w^{2}}R_{n}}\geq t\Bigm|\mathcal{O}_{train},X^{n}\right\}\right]$
$\leq\left(\frac{1}{t^{2}R_{n}^{2}}\right)\mathbb{E}\left\{\|\sigma\|_{w^{2}}^{2}\mathbb{E}\left(\frac{\|T_{n}\|^{2}}{\|\widehat{f}-f\|_{w^{2}}^{2}}\Bigm|\mathcal{O}_{train},X^{n}\right)\right\}$
$\leq\left(\frac{1}{t^{2}R_{n}^{2}}\right)\mathbb{E}\left\{\|\sigma\|_{w^{2}}^{2}\sum_{i=1}^{n}\|w_{i}(\cdot;X^{n})\|^{2}\right\}\leq\frac{1}{t^{2}}$
where the second line follows by Markov’s inequality, the third from the bound in (9) and iterated expectation, and the last from the bound in (10). The result follows since we can always pick $t^{2}=1/\epsilon$ to ensure the above probability is no more than any $\epsilon$. ∎
[…]
Lemma A.1.
Under the same setup from Assumptions A.1, define (i) $m(x)=\mathbb{E}[f(V_{i})\mid X_{i}=x]$ the conditional expectation of $f(V_{i})$ given $X_{i}$; (ii) $\hat{m}(x):=\hat{\mathbb{E}}_{n}[\hat{f}(V_{i})\mid X_{i}=x]$ the regression of $\hat{f}(V_{i})$ on $X_{i}$ in the test samples; (iii) $\tilde{m}(x):=\hat{\mathbb{E}}_{n}[f(V_{i})\mid X_{i}=x]$ the oracle regression of $f(V_{i})$ on $X_{i}$ in the test samples. Furthermore, let $\tilde{b}(x):=\hat{\mathbb{E}}_{n}[b(V_{i})\mid X_{i}=x]$ be the $\hat{\mathbb{E}}_{n}$-smoothed bias and $R_{n}^{2}=E[\|\widetilde{m}-m\|]^{2}$ be the oracle $L_{2}$-error. If
  1. i.
    the regression estimator $\hat{\mathbb{E}}_{n}[\cdot]$ is $L_{2}(\mathbb{P})$-stable with respect to distance metric $d(\cdot,\cdot)$;
  2. ii.
    $d(\hat{f},f)\xrightarrow{p}0$,
then
$\|\hat{m}-\tilde{m}\|=\|\tilde{b}(\cdot)\|+o_{\mathbb{P}}(R_{n}).$
If further $\|\tilde{b}\|=o_{\mathbb{P}}\left(\sqrt{\mathbb{E}\|\widetilde{m}-m\|^{2}}\right)$, then $\widehat{m}$ is oracle efficient in the $L_{2}$-norm, i.e., asymptotically equivalent to the oracle estimator $\widetilde{m}$ in the sense that
$\frac{\|\widehat{m}-\widetilde{m}\|}{\sqrt{\mathbb{E}\|\widetilde{m}-m\|^{2}}}\ \xrightarrow{p}\ 0$
and
$\|\widehat{m}-m\|=\|\tilde{m}-m\|+o_{\mathbb{P}}(R_{n}).$
Proof.
Note
$\|\widehat{m}-\widetilde{m}\|$ $\leq\|\widehat{m}-\widetilde{m}-\widetilde{b}\|+\|\widetilde{b}\|=\|\widetilde{b}\|+o_{\mathbb{P}}\Big(\sqrt{\mathbb{E}\|\widetilde{m}-m\|^{2}}\Big)$
where the first equality follows by the triangle inequality, and the second equality by the $L_{2}(\mathbb{P})$-stability and $d(\cdot,\cdot)$-consistency of $\widehat{f}$. ∎
This generalizes Proposition 1 of Kennedy(22)-towardsoptimal, which shows that a pointwise stability condition and consistency of $\hat{f}$ implies an oracle inequality on the pointwise convergence of a feasible pseudo-outcome regression. In Section 3 of the main text, we apply Lemma A.1 to analyze the convergence of our proposed estimators for the conditional probability bounds.
Appendix B Proofs of Results in the Main Text
B.1 Proof of Lemma 2.1
The statements for $\mathcal{H}(\mu^{*}(x))$ and $\mathcal{H}(\mbox{perf}(s;\beta))$ follow immediately since (i) both $\mu^{*}(x)$ and $\mbox{perf}(s;\beta)$ are linear in $\delta(\cdot)$, and (ii) $\Delta$ is closed and convex. To prove the statement for $\mbox{perf}_{+}(s;\beta)$, define
$\mbox{perf}_{+}(s;\beta,\delta):=\frac{\mathbb{E}[\beta_{0,i}\mu_{1}(X_{i})+\beta_{0,i}\pi_{0}(X_{i})\delta(X_{i})]}{\mathbb{E}[\mu_{1}(X_{i})+\pi_{0}(X_{i})\delta(X_{i})]}.$
[…]
Towards this, we apply a change-of-variables. Let $U(\cdot)\colon\mathcal{X}\rightarrow[0,1]$ be defined as $U(x)=\frac{\delta(x)-\underline{\delta}(x)}{\overline{\delta}(x)-\underline{\delta}(x)}$. For any $\delta(\cdot)\in\Delta$, there exists $U(\cdot)\in[0,1]$ such that $\mbox{perf}_{+}(s;\beta,\delta)=\mbox{perf}_{+}(s;\beta,U)$, where
$\mbox{perf}_{+}(s;\beta,U):=\frac{\mathbb{E}[\beta_{0,i}\mu_{1}(X_{i})+\beta_{0,i}\pi_{0}(X_{i})\underline{\delta}(X_{i})+\beta_{0,i}\pi_{0}(X_{i})(\overline{\delta}(X_{i})-\underline{\delta}(X_{i}))U(X_{i})]}{\mathbb{E}[\mu_{1}(X_{i})+\pi_{0}(X_{i})\underline{\delta}(X_{i})+\pi_{0}(X_{i})(\overline{\delta}(X_{i})-\underline{\delta}(X_{i}))U(X_{i})]}.$
Conversely, for any $U(\cdot)\in[0,1]$, there exists a corresponding $\delta(\cdot)\in\Delta$ such that $\mbox{perf}_{+}(s;\beta,U)=\mbox{perf}_{+}(s;\beta,\delta)$, where $\delta(x)=\underline{\delta}(x)+(\overline{\delta}(x)-\underline{\delta}(x))U(x)$.
Next, apply the Charnes-Cooper transformation with
$\tilde{V}=\frac{1}{\mathbb{E}[\mu_{1}(X_{i})+\pi_{0}(X_{i})\underline{\delta}(X_{i})+\pi_{0}(X_{i})(\overline{\delta}(X_{i})-\underline{\delta}(X_{i}))U(X_{i})]}$
$\tilde{U}(\cdot)=\frac{U(\cdot)}{\mathbb{E}[\mu_{1}(X_{i})+\pi_{0}(X_{i})\underline{\delta}(X_{i})+\pi_{0}(X_{i})(\overline{\delta}(X_{i})-\underline{\delta}(X_{i}))U(X_{i})]}.$
[…]
$\mbox{perf}_{+}(s;\beta,\tilde{U},\tilde{V})=\mathbb{E}[\beta_{0}(X_{i})\mu_{1}(X_{i})+(1-D_{i})\beta_{0}(X_{i})\underline{\delta}(X_{i})]\tilde{V}+\mathbb{E}[\beta_{0}(X_{i})\pi_{0}(X_{i})(\overline{\delta}(X_{i})-\underline{\delta}(X_{i}))\tilde{U}(X_{i})].$
Conversely, for any such $\tilde{U}(\cdot),\tilde{V}$, there exists $U(\cdot)\in[0,1]$ such that $\mbox{perf}_{+}(s;\beta,\tilde{U},\tilde{V})=\mbox{perf}_{+}(s;\beta,U)$.
Now consider any $\tilde{p}\in[\underline{\mbox{perf}}_{+}(s;\beta),\overline{\mbox{perf}}_{+}(s;\beta)]$, which satisfies for some $\lambda\in[0,1]$
$\tilde{p}=\lambda\underline{\mbox{perf}}_{+}(s;\beta)+(1-\lambda)\overline{\mbox{perf}}_{+}(s;\beta).$
Let $\underline{\delta}(\cdot),\overline{\delta}(\cdot)$ be the functions achieving the infimum and supremum respectively
$\underline{\delta}(\cdot)\in\arg\min_{\delta\in\Delta}\mbox{perf}_{+}(s;\beta,\delta),\mbox{ }\overline{\delta}(\cdot)\in\arg\max_{\delta\in\Delta}\mbox{perf}_{+}(s;\beta,\delta).$
[…]
B.2 Proof of Lemma 2.2
Define the bounds
$\underline{U}(d):=\mathbb{E}[\left(u_{1,0,i}-(u_{1,1,i}+u_{1,0,i})\overline{\mu}^{*}(X_{i})\right)d(X_{i})+\left(-u_{0,0,i}+(u_{0,0,i}+u_{0,1,i})\underline{\mu}^{*}(X_{i})\right)(1-d(X_{i}))],$
$\overline{U}(d):=\mathbb{E}[\left(u_{1,0,i}-(u_{1,1,i}+u_{1,0,i})\underline{\mu}^{*}(X_{i})\right)d(X_{i})+\left(-u_{0,0,i}+(u_{0,0,i}+u_{0,1,i})\overline{\mu}^{*}(X_{i})\right)(1-d(X_{i}))].$
At each value $x\in\mathcal{X}$, notice that if $d(x)=1$, then
$(-u_{1,1,i}\mu^{*}(x)+u_{1,0,i}(1-\mu^{*}(x)))d(x)+(-u_{0,0,i}(1-\mu^{*}(x))+u_{0,1,i}\mu^{*}(x))(1-d(x))=u_{1,0,i}-(u_{1,1,i}+u_{1,0,i})\mu^{*}(x).$
This is minimized over $\mu^{*}(x)\in\mathcal{H}(\mu^{*}(x))$ at $\mu^{*}(x)=\overline{\mu}^{*}(x)$. If $d(x)=0$, then
$(-u_{1,1,i}\mu^{*}(x)+u_{1,0,i}(1-\mu^{*}(x)))d(x)+(-u_{0,0,i}(1-\mu^{*}(x))+u_{0,1,i}\mu^{*}(x))(1-d(x))=-u_{0,0,i}+(u_{0,0,i}+u_{0,1,i})\mu^{*}(x).$
This is minimized over $\mu^{*}(x)\in\mathcal{H}(\mu^{*}(x))$ at $\mu^{*}(x)=\underline{\mu}^{*}(x)$. The result for the lower bound immediately follows. The upper bound follows analogously. $\Box$
B.3 Proof of Proposition 3.1
We prove the result for our estimator of the upper bound, and the same argument applies for our estimator of the lower bound. We define $\widehat{\pi}_{0}(x)=1-\widehat{\pi}_{1}(x)$ throughout. Let $\mathcal{O}_{1}$ denote the observations in the first fold and $\mathcal{O}_{2}$ denote the observations in the second fold. We first observe that
$\|\widehat{\overline{\mu}}(\cdot;\Delta)-\overline{\mu}^{*}(\cdot;\Delta)\|$ $\leq\|\widehat{\overline{\mu}}(\cdot;\Delta)-\widehat{\overline{\mu}}_{oracle}(\cdot;\Delta)\|+\|\widehat{\overline{\mu}}_{oracle}(\cdot;\Delta)-\overline{\mu}^{*}(\cdot;\Delta)\|$
$\leq\|\widehat{\overline{\mu}}(\cdot;\Delta)-\widehat{\overline{\mu}}_{oracle}(\cdot;\Delta)-\tilde{b}(\cdot)\|+\|\tilde{b}(\cdot)\|+\|\widehat{\overline{\mu}}_{oracle}(\cdot;\Delta)-\overline{\mu}^{*}(\cdot;\Delta)\|$
for $\tilde{b}(x)=\widehat{\mathbb{E}}_{n}[\hat{b}(X_{i})\mid X_{i}=x]$ the smoothed bias and
$\hat{b}(x)=$ $\underbrace{\mathbb{E}[\phi_{\mu,i}(\hat{\eta})-\phi_{\mu,i}(\eta)\mid\mathcal{O}_{1},X_{i}=x]}_{(a)}+(\overline{\Gamma}-1)\underbrace{\mathbb{E}[\phi_{\pi\mu,i}(\hat{\eta})-\phi_{\pi\mu,i}(\eta)\mid\mathcal{O}_{1},X_{i}=x]}_{(b)}.$
Under Assumption A.1, $\|\widehat{\overline{\mu}}(\cdot)-\widehat{\overline{\mu}}_{oracle}(\cdot)-\tilde{b}(\cdot)\|=o_{\mathbb{P}}(R_{oracle})$ by Lemma A.1. Furthermore, $\hat{b}(x)^{2}\leq 2(a)^{2}+2(\bar{\Gamma}-1)^{2}(b)^{2}$, where
$(a)^{2}=\left\{\frac{\pi_{1}(x)-\hat{\pi}_{1}(x)}{\hat{\pi}_{1}(x)}(\mu_{1}(x)-\hat{\mu}_{1}(x))\right\}^{2}\leq\frac{1}{\epsilon^{2}}\left\{(\pi_{1}(x)-\hat{\pi}_{1}(x))(\mu_{1}(x)-\hat{\mu}_{1}(x))\right\}^{2},$
and
$(b)^{2}=\left\{(\pi_{0}(x)-\hat{\pi}_{0}(x))\hat{\mu}_{1}(x)+\frac{\pi_{1}(x)}{\hat{\pi}_{1}(x)}\left(\mu_{1}(x)-\hat{\mu}_{1}(x)\right)\hat{\pi}_{0}(x)+\hat{\pi}_{0}(x)\hat{\mu}_{1}(x)-\pi_{0}(x)\mu_{1}(x)\right\}^{2}=$
$\left\{(\pi_{0}(x)-\hat{\pi}_{0}(x))\hat{\mu}_{1}(x)+\frac{\pi_{1}(x)}{\hat{\pi}_{1}(x)}\left(\mu_{1}(x)-\hat{\mu}_{1}(x)\right)\hat{\pi}_{0}(x)+\hat{\pi}_{0}(x)(\hat{\mu}_{1}(x)-\mu_{1}(x))+\mu_{1}(x)(\hat{\pi}_{0}(x)-\pi_{0}(x))\right\}^{2}=$
$\left\{(\pi_{0}(x)-\hat{\pi}_{0}(x))(\hat{\mu}_{1}(x)-\mu_{1}(x))+\frac{\hat{\pi}_{0}(x)}{\hat{\pi}_{1}(x)}(\pi_{1}(x)-\hat{\pi}_{1}(x))(\mu_{1}(x)-\hat{\mu}_{1}(x))\right\}^{2}=$
$\left\{(\pi_{1}(x)-\hat{\pi}_{1}(x))(\mu_{1}(x)-\hat{\mu}_{1}(x))+\frac{\hat{\pi}_{0}(x)}{\hat{\pi}_{1}(x)}(\pi_{1}(x)-\hat{\pi}_{1}(x))(\mu_{1}(x)-\hat{\mu}_{1}(x))\right\}^{2}\leq$
$\frac{1}{\epsilon^{2}}\left\{(\pi_{1}(x)-\hat{\pi}_{1}(x))(\hat{\mu}_{1}(x)-\mu_{1}(x))\right\}^{2}$
by iterated expectations and the assumption of bounded propensity score. Putting this together yields
$\|\widehat{\overline{\mu}}(\cdot)-\overline{\mu}^{*}(\cdot)\|\leq\|\widehat{\overline{\mu}}_{oracle}(\cdot)-\overline{\mu}^{*}(\cdot)\|+\sqrt{2}\epsilon^{-1}(\overline{\Gamma}-1)\|\tilde{R}(\cdot)\|+o_{\mathbb{P}}(R_{oracle})$
as desired. $\Box$
B.4 Proof of Proposition 3.2
Recall from Lemma 2.2 that, for any decision rule $d(\cdot)\colon\mathcal{X}\rightarrow\{0,1\}$,
$\underline{U}(d):=\mathbb{E}[\left(u_{1,0,i}-(u_{1,1,i}+u_{1,0,i})\overline{\mu}^{*}(x)\right)d(X_{i})+\left(-u_{0,0,i}+(u_{0,0,i}+u_{0,1,i})\underline{\mu}^{*}(x)\right)(1-d(X_{i}))]=$
$\mathbb{E}[-u_{0,0,i}+(u_{0,0,i}+u_{0,1,i})\underline{\mu}^{*}(X_{i})]+\mathbb{E}[\left((u_{1,0,i}+u_{0,0,i})-\tilde{\mu}^{*}(x)\right)d(X_{i})]$
for $\widetilde{\mu}^{*}(x)=(u_{1,1,i}+u_{1,0,i})\overline{\mu}^{*}(x)+(u_{0,0,i}+u_{0,1,i})\underline{\mu}^{*}(x)$. Therefore, we can rewrite regret as
$R(\hat{d})=\underline{U}(d^{*})-\underline{U}(\hat{d})=\mathbb{E}[\left(c(X_{i})-\tilde{\mu}^{*}(X_{i})\right)(d^{*}(X_{i})-\hat{d}(X_{i})],$
defining $c(X_{i}):=u_{1,0}(X_{i})+u_{0,0}(X_{i})$. It then follows that
$R(\hat{d})=\int_{x\in\mathcal{X}}\left(c(x)-\tilde{\mu}^{*}(x)\right)\left(d^{*}(x)-\hat{d}(x)\right)dP(x)\leq\int_{x\in\mathcal{X}}\left|\tilde{\mu}^{*}(x)-c(x)\right|1\{d^{*}(x)\neq\hat{d}(x)\}dP(x).$
At any fixed $X_{i}=x$, $\hat{d}(X_{i})\neq d^{*}(X_{i})$ implies that $|\tilde{\mu}^{*}(x)-\widehat{\tilde{\mu}}(x)|\geq|\tilde{\mu}^{*}(x)-c(x)|$. Combining this with the previous display implies $R(\hat{d})\leq\int_{x\in\mathcal{X}}|\tilde{\mu}^{*}(x)-\widehat{\tilde{\mu}}(x)|dP(x).$ Substituting in the definition of $\tilde{\mu}^{*}(x)$ and $\widehat{\tilde{\mu}}(x)$, we have
$|\tilde{\mu}^{*}(x)-\widehat{\tilde{\mu}}(x)|=$
$|(u_{1,1}(x)+u_{1,0}(x))\overline{\mu}^{*}(x)+(u_{0,0}(x)+u_{0,1}(x))\underline{\mu}^{*}(x)-(u_{1,1}(x)+u_{1,0}(x))\widehat{\overline{\mu}}(x)-(u_{0,0}(x)+u_{0,1}(x))\widehat{\underline{\mu}}(x)|\leq$
$|\overline{\mu}^{*}(x)-\widehat{\overline{\mu}}(x)|+|\underline{\mu}^{*}(x)-\widehat{\underline{\mu}}(x)|$
by the triangle inequality and using $u_{0,0}(x),u_{0,1}(x),u_{1,0}(x),u_{1,1}(x)$ are non-negative and sum to one. Substituting back into the bound on R(d^)R($\hat${d}) delivers
$R(\hat{d})\leq\int_{x\in\mathcal{X}}|\overline{\mu}^{*}(x)-\widehat{\overline{\mu}}(x)|dP(x)+\int_{x\in\mathcal{X}}|\underline{\mu}^{*}(x)-\widehat{\underline{\mu}}(x)|dP(x)=\|\overline{\mu}^{*}(x)-\widehat{\overline{\mu}}(x)\|_{1}+\|\underline{\mu}^{*}(x)-\widehat{\underline{\mu}}(x)\|_{1}.$
Using the Cauchy-Schwarz inequality $\|\overline{\mu}^{*}(x)-\widehat{\overline{\mu}}(x)\|_{1}^{2}\leq\|\overline{\mu}^{*}(x)-\widehat{\overline{\mu}}(x)\|_{2}^{2}$ and $\|\underline{\mu}^{*}(x)-\widehat{\underline{\mu}}(x)\|_{1}^{2}\leq\|\underline{\mu}^{*}(x)-\widehat{\underline{\mu}}(x)\|_{2}^{2}$ and the inequality $(a+b)^{2}\leq 2(a^{2}+b^{2})$,
$R(\hat{d})^{2}\leq 2\|\overline{\mu}^{*}(x)-\widehat{\overline{\mu}}(x)\|_{2}^{2}+2\|\underline{\mu}^{*}(x)-\widehat{\underline{\mu}}(x)\|_{2}^{2}.$
The result then follows by applying Proposition 3.1. $\Box$
B.5 Proof of Proposition 4.1
To prove Proposition 4.1, we first state and prove various auxiliary lemmas. Let $\mathcal{O}_{k}$ denote the observations in the $k$-th fold and $\mathcal{O}_{-k}$ denote the observations not in the $k$-th fold.
Lemma B.1.
Let $\beta(\cdot)$ be some function of $X_{i}$ such that $\|\beta(\cdot)\|\leq M$ for some $M<\infty$ and define $R_{1,n}^{k}=\|\hat{\mu}_{1,-k}(\cdot)-\mu_{1}(\cdot)\|\|\hat{\pi}_{1,-k}(\cdot)-\pi_{1}(\cdot)\|$. Assume that there exists $\epsilon>0$ s.t. $\mathbb{P}(\hat{\pi}_{1,-k}(X_{i})\geq\epsilon)=1$. Then,
$\mathbb{E}[\beta(X_{i})\phi_{\mu,i}(\hat{\eta}_{-k})-\beta(X_{i})\phi_{\mu,i}(\eta)\mid\mathcal{O}_{-k}]=O_{\mathbb{P}}(R_{1,n}^{k}).$
Proof.
We follow the proof of Lemma 3 in MishlerEtAl(21). Suppressing the dependence on $\mathcal{O}_{-k}$ to ease notation, observe that
$\mathbb{E}[\beta(X_{i})\phi_{\mu,i}(\hat{\eta}_{-k})-\beta(X_{i})\phi_{\mu,i}(\eta)]=$
$\mathbb{E}\left[\beta(X_{i})\left(\frac{D_{i}}{\hat{\pi}_{1}(X_{i})}(Y_{i}-\hat{\mu}_{1}(X_{i}))-\frac{D}{\pi_{1}(X_{i})}(Y_{i}-\mu_{1}(X_{i}))+(\hat{\mu}_{1}(X_{i})-\mu_{1}(X_{i}))\right)\right]\overset{(1)}{=}$
$\mathbb{E}\left[\beta(X_{i})\left(\frac{\pi_{1}(X_{i})}{\hat{\pi}_{1}(X_{i})}(\mu_{1}(X_{i})-\hat{\mu}_{1}(X_{i}))+(\hat{\mu}_{1}(X_{i})-\mu_{1}(X_{i})\right)\right]=$
$\mathbb{E}\left[\beta(X_{i})\frac{(\hat{\mu}_{1}(X_{i})-\mu_{1}(X_{i}))(\hat{\pi}_{1}(X_{i})-\pi_{1}(X_{i}))}{\hat{\pi}_{1}(X_{i})}\right]\overset{(2)}{\leq}\\$
$\epsilon^{-1}\mathbb{E}[\beta(X_{i})(\hat{\mu}_{1}(X_{i})-\mu_{1}(X_{i}))(\hat{\pi}_{1}(X_{i})-\pi_{1}(X_{i}))],$
where (1) follows by iterated expectations and (2) by the assumption of a bounded propensity score estimator. The result follows by applying the Cauchy-Schwarz inequality and using $\|\beta(\cdot)\|\leq M$ to conclude that $\|\mathbb{E}[\beta(X_{i})\phi_{\mu,i}(\hat{\eta})-\beta(X_{i})\phi_{\mu,i}(\eta)]\|=O_{\mathbb{P}}(R_{1,n}^{k})$. ∎
Lemma B.2 (Lemma 2 in KennedyEtAl(2020)-SharpInstruments).
Let $\hat{\phi}(X_{i})$ be a function estimated from a sample $O_{i}:=(X_{i},D_{i},Y_{i})\sim P(\cdot)$ i.i.d. for $i=1,\ldots,N$ and let $\mathbb{E}_{n}[\cdot]$ denote the empirical average over another independent sample $O_{j}\sim P(\cdot)$ i.i.d. for $j=N+1,\ldots,n$. Then,
$\mathbb{E}_{n}[\hat{\phi}(O_{i})-\phi(O_{i})]-\mathbb{E}[\hat{\phi}(O_{i})-\phi(O_{i})]=O_{\mathbb{P}}\left(\frac{\|\hat{\phi}(\cdot)-\phi(\cdot)\|}{\sqrt{n}}\right).$
Lemma B.3.
Let $\beta(\cdot)$ be some function of $X_{i}$ such that $\|\beta(\cdot)\|\leq M$ for some $M<\infty$. Let $\hat{\phi}(O_{i})$ be a function estimated from a sample $O_{i}:=(X_{i},D_{i},Y_{i})\sim P(\cdot)$ i.i.d. for $i=1,\ldots,N$ and let $\mathbb{E}_{n}[\cdot]$ denote the empirical average over another independent sample $O_{j}:=(X_{j},D_{j},Y_{j})\sim P(\cdot)$ i.i.d. for $j=N+1,\ldots,n$. Then,
$\mathbb{E}_{n}[\beta(X_{i})\hat{\phi}(O_{i})-\beta(X_{i})\phi(O_{i})]-\mathbb{E}[\beta(X_{i})\hat{\phi}(O_{i})-\beta(X_{i})\phi(O_{i})]=O_{P}\left(\frac{\|\hat{\phi}(\cdot)-\phi(\cdot)\|}{\sqrt{n}}\right).$
Proof.
The proof follows the same argument as the proof of Lemma 2 in KennedyEtAl(2020)-SharpInstruments. ∎
[…]
Define $\|\hat{\mu}_{1,-k}(\cdot)-\mu_{1}(\cdot)\|\|\hat{\pi}_{1,-k}(\cdot)-\pi_{1}(\cdot)\|=R_{1,n}^{k}$. Assume (i) there exists $\delta>0$ such that $\mathbb{P}(\pi_{1}(X_{i})\geq\delta)=1$; and (ii) there exists $\epsilon>0$ such that $\mathbb{P}(\hat{\pi}_{1-k}(X_{i})\geq\epsilon)=1$. Then,
$\|\phi_{\mu,i}(\hat{\eta}_{-k})-\phi_{\mu,i}(\eta)\|=O_{\mathbb{P}}(R_{1,n}^{k}+\|\hat{\pi}_{1,-k}-\pi_{1}\|+\|\hat{\mu}_{1,-k}-\mu_{1}\|).$
Proof.
This result follows directly from the stated conditions after some algebra. Suppressing dependence on the subscript k-k to ease notation, observe that we can rewrite
$\left\|\phi_{\mu,i}(\hat{\eta})-\phi_{\mu,i}(\eta)\right\|=\left\|\frac{D_{i}}{\hat{\pi}_{1}(X_{i})}(Y_{i}-\hat{\mu}_{1}(X_{i}))-\frac{D_{i}}{\pi_{1}(X_{i})}(Y_{i}-\mu_{1}(X_{i}))+(\mu_{1}(X_{i})-\hat{\mu}_{1}(X_{i}))\right\|\overset{(1)}{=}$
$\left\|\frac{D_{i}}{\pi_{1}(X_{i})}\frac{\pi_{1}(X_{i})-\hat{\pi}_{1}(X_{i})}{\hat{\pi}_{1}(X_{i})}(Y_{i}-\hat{\mu}_{1}(X_{i}))-\frac{D_{i}}{\pi_{1}(X_{i})}(\hat{\mu}_{1}(X_{i})-\mu_{1}(X_{i}))+(\hat{\mu}_{1}(X_{i})-\mu_{1}(X_{i}))\right\|\overset{(2)}{\leq}$
$\left\|\frac{D_{i}}{\pi_{1}(X_{i})}\frac{\pi_{1}(X_{i})-\hat{\pi}_{1}(X_{i})}{\hat{\pi}_{1}(X_{i})}(Y_{i}-\mu_{1}(X_{i}))\right\|+\left\|\frac{D_{i}}{\pi_{1}(X_{i})}\frac{\pi_{1}(X_{i})-\hat{\pi}_{1}(X_{i})}{\hat{\pi}_{1}(X_{i})}(\mu_{1}(X_{i})-\hat{\mu}_{1}(X_{i}))\right\|+$
$\left\|\frac{D_{i}}{\pi_{1}(X_{i})}(\hat{\mu}_{1}(X_{i})-\mu_{1}(X_{i}))\right\|+\left\|(\hat{\mu}_{1}(X_{i})-\mu_{1}(X_{i}))\right\|\overset{(3)}{\leq}$
$\frac{\|D_{i}\|}{\delta}\frac{\|\pi_{1}-\hat{\pi}_{1}\|}{\epsilon}\|Y_{i}-\mu_{1}(X_{i})\|+\frac{\|D_{i}\|}{\delta}\frac{\|\pi_{1}-\hat{\pi}_{1}\|}{\epsilon}\|\hat{\mu}_{1}-\mu_{1}\|+\frac{\|D_{i}\|}{\delta}\|\hat{\mu}_{1}-\mu_{1}\|+\left\|\hat{\mu}_{1}-\mu_{1}\right\|$
where (1) follows by adding and subtracting $\frac{D_{i}}{\pi_{1}(X_{i})}(Y_{i}-\hat{\mu}_{1}(X_{i}))$, (2) follows by adding and subtracting $\frac{D_{i}}{\pi_{1}(X_{i})}\frac{\pi_{1}(X_{i})-\hat{\pi}_{1}(X_{i})}{\hat{\pi}_{1}(X_{i})}\mu_{1}(X_{i})$ and applying the triangle inequality, and (3) applies the assumption of strict overlap and bounded propensity score estimator. ∎
[…]
Define $\|\hat{\mu}_{1,-k}(\cdot)-\mu_{1}(\cdot)\|\|\hat{\pi}_{1,-k}(\cdot)-\pi_{1}(\cdot)\|=R_{1,n}^{k}$. Assume that (i) there exists $\delta>0$ such that $\mathbb{P}(\pi_{1}(X_{i})\geq\delta)=1$; and (ii) there exists $\epsilon>0$ such that $\mathbb{P}(\hat{\pi}_{1,-k}(X_{i})\geq\epsilon)=1$. Then,
$\|\phi_{\pi\mu,i}(\hat{\eta}_{-k})-\phi_{\pi\mu,i}(\eta)\|=O_{\mathbb{P}}(R_{1,n}^{k}+\|\hat{\pi}_{1,-k}-\pi_{1}\|+\|\hat{\mu}_{1,-k}-\mu_{1}\|)$
Proof.
This result follows directly from the stated conditions after some simple algebra. We omit the dependence on $X_{i}$ and k-k. Observe that we can rewrite
[…]
$\|((1-D_{i})-\hat{\pi}_{0})\hat{\mu}_{1}+\frac{D_{i}}{\hat{\pi}_{1}}(Y_{i}-\hat{\mu}_{1})\hat{\pi}_{0}+\hat{\pi}_{0}\hat{\mu}_{1}-((1-D_{i})-\pi_{0})\mu_{1}-\frac{D_{i}}{\pi_{1}}(Y_{i}-\mu_{1})\pi_{0}-\pi_{0}\mu_{1}\|\overset{(1)}{=}$
$\|((1-D_{i})-\hat{\pi}_{0})\hat{\mu}_{1}+\frac{D_{i}}{\pi_{1}}\frac{\pi_{1}-\hat{\pi}_{1}}{\hat{\pi}_{1}}(Y_{i}-\hat{\mu}_{1})+\frac{D_{i}}{\pi_{1}}\pi_{0}(\mu_{1}-\hat{\mu}_{1})+\hat{\pi}_{0}\hat{\mu}_{1}-\pi_{0}\mu_{1}\|\overset{(2)}{\leq}$
$\|\frac{D_{i}}{\pi_{1}}\frac{\pi_{1}-\hat{\pi}_{1}}{\hat{\pi}_{1}}(Y_{i}-\mu_{1})+\frac{D_{i}}{\pi_{1}}\frac{\pi_{1}-\hat{\pi}_{1}}{\hat{\pi}_{1}}(\mu_{1}-\hat{\mu}_{1})+\frac{D_{i}}{\pi_{1}}\pi_{0}(\mu_{1}-\hat{\mu}_{1})+((1-D_{i})-\hat{\pi}_{0})\hat{\mu}_{1}+\hat{\pi}_{0}\hat{\mu}_{1}-\pi_{0}\mu_{1}\|\overset{(3)}{=}$
$\|\frac{D_{i}}{\pi_{1}}\frac{\pi_{1}-\hat{\pi}_{1}}{\hat{\pi}_{1}}(Y_{i}-\mu_{1})+\frac{D_{i}}{\pi_{1}}\frac{\pi_{1}-\hat{\pi}_{1}}{\hat{\pi}_{1}}(\mu_{1}-\hat{\mu}_{1})+\frac{D_{i}}{\pi_{1}}\pi_{0}(\mu_{1}-\hat{\mu}_{1})+((1-D_{i})-\hat{\pi}_{0})(\hat{\mu}_{1}-\mu_{1})+(\pi_{0}-\hat{\pi}_{0})\mu_{1}+\hat{\pi}_{0}\hat{\mu}_{1}-\pi_{0}\mu_{1}\|\overset{(4)}{\leq}$
$\|\frac{D_{i}}{\pi_{1}}\frac{\pi_{1}-\hat{\pi}_{1}}{\hat{\pi}_{1}}(Y_{i}-\mu_{1})+\frac{D_{i}}{\pi_{1}}\frac{\pi_{1}-\hat{\pi}_{1}}{\hat{\pi}_{1}}(\mu_{1}-\hat{\mu}_{1})+\frac{D_{i}}{\pi_{1}}\pi_{0}(\mu_{1}-\hat{\mu}_{1})+((1-D_{i})-\pi_{0})(\hat{\mu}_{1}-\mu_{1})+(\pi_{0}-\hat{\pi}_{0})(\hat{\mu}_{1}-\mu_{1})+(\pi_{0}-\hat{\pi}_{0})\mu_{1}\|+$
$\|\hat{\pi}_{0}\hat{\mu}_{1}-\pi_{0}\mu_{1}\|\overset{(5)}{=}$
$\|\frac{D_{i}}{\pi_{1}}\frac{\pi_{1}-\hat{\pi}_{1}}{\hat{\pi}_{1}}(Y_{i}-\mu_{1})+\frac{D_{i}}{\pi_{1}}\frac{\pi_{1}-\hat{\pi}_{1}}{\hat{\pi}_{1}}(\mu_{1}-\hat{\mu}_{1})+\frac{D_{i}}{\pi_{1}}\pi_{0}(\mu_{1}-\hat{\mu}_{1})+((1-D_{i})-\pi_{0})(\hat{\mu}_{1}-\mu_{1})+(\pi_{0}-\hat{\pi}_{0})(\hat{\mu}_{1}-\mu_{1})+(\pi_{0}-\hat{\pi}_{0})\mu_{1}\|+$
$\|\hat{\pi}_{0}(\hat{\mu}_{1}-\mu_{1})-\mu_{1}(\pi_{0}-\hat{\pi}_{0})\|$
where (1) follows by adding/subtracting $\frac{D_{i}}{\pi_{1}}(Y_{i}-\hat{\mu}_{1})$, (2) follows by adding/subtracting $\frac{D_{i}}{\pi_{1}}\frac{\pi_{1}-\hat{\pi}_{1}}{\hat{\pi}_{1}}\mu_{1}$, (3) follows by adding/subtracting $\mu_{1}((1-D_{i})-\hat{\pi}_{0})$, (4) follows by adding/subtracting $\pi_{0}(\hat{\mu}_{1}-\mu_{1})$ and applying the triangle inequality once, and (5) follows by adding/subtracting $\hat{\pi}_{0}\mu_{1}$. We then again apply the triangle inequality and use the assumptions of strict overlap and bounded propensity score estimator to arrive at
$\leq\frac{1}{\epsilon\delta}\|\hat{\pi}_{1}-\pi_{1}\|\|Y_{i}-\mu_{1}\|+\frac{1}{\epsilon\delta}\|\hat{\pi}_{1}-\pi_{1}\|\|\mu_{1}-\hat{\mu}_{1}\|+\frac{1-\delta}{\delta}\|\mu_{1}-\hat{\mu}_{1}\|+\|(1-D_{i})-\pi_{0}\|\|\hat{\mu}_{1}-\mu_{1}\|+$
$\|\hat{\pi}_{1}-\pi_{1}\|\|\hat{\mu}_{1}-\mu_{1}\|+\|\hat{\pi}_{1}-\pi_{1}\|\|\mu_{1}\|+(1-\epsilon)\|\hat{\mu}_{1}-\mu_{1}\|+\|\mu_{1}\|\|\hat{\pi}_{1}-\pi_{1}\|.$
The result then follows. ∎
Lemma B.6.
Let $\beta(\cdot)$ be some function of $X_{i}$ such that $\|\beta(\cdot)\|\leq M$ for some $M<\infty$ and define $R_{1,n}^{k}=\|\hat{\mu}_{1,-k}(\cdot)-\mu_{1}(\cdot)\|\|\hat{\pi}_{1,-k}(\cdot)-\pi_{1}(\cdot)\|$. Assume that there exists $\epsilon>0$ s.t. $\mathbb{P}(\hat{\pi}_{1,-k}(X_{i})\geq\epsilon)=1$. Then,
$\mathbb{E}[\beta(X_{i})\left(\phi_{\pi\mu,i}(\hat{\eta}_{-k})-\phi_{\pi\mu,i}(\eta)\right)\mid\mathcal{O}_{-k}]=O_{\mathbb{P}}(R_{1,n}^{k})$
Proof.
For ease of notation, we omit the dependence on $X_{i}$ and the subscript k-k. The proof follows an analogous argument to Lemma B.1. Observe that
$\mathbb{E}[\beta(X_{i})\left(\phi_{\pi\mu,i}(\hat{\eta})-\phi_{\pi\mu,i}(\eta)\right)]=$
$\mathbb{E}[\beta(X_{i})\left\{((1-D_{i})-\hat{\pi}_{0})\hat{\mu}_{1}+\frac{D_{i}}{\hat{\pi}_{1}}(Y_{i}-\hat{\mu}_{1})\hat{\pi}_{0}+\hat{\pi}_{0}\hat{\mu}_{1}-((1-D_{i})-\pi_{0})\mu_{1}-\frac{D_{i}}{\pi_{1}}(Y_{i}-\mu_{1})\pi_{0}-\pi_{0}\mu_{1}\right\}]\overset{(1)}{=}$
$\mathbb{E}[\beta(X_{i})\left\{(\pi_{0}-\hat{\pi}_{0})\hat{\mu}_{1}+\frac{\pi_{1}}{\hat{\pi}_{1}}(\mu_{1}-\hat{\mu}_{1})\hat{\pi}_{0}\right\}+\hat{\pi}_{0}\hat{\mu}_{1}-\pi_{0}\mu_{1}\}]\overset{(2)}{=}$
$\mathbb{E}[\beta(X_{i})\left\{(\pi_{0}-\hat{\pi}_{0})\hat{\mu}_{1}+\frac{\pi_{1}}{\hat{\pi}_{1}}(\mu_{1}-\hat{\mu}_{1})\hat{\pi}_{0}+\hat{\pi}_{0}(\hat{\mu}_{1}-\mu_{1})+\mu_{1}(\hat{\pi}_{0}-\pi_{0})\right\}]=$
$\mathbb{E}[\beta(X_{i})\left\{(\hat{\mu}_{1}-\mu_{1})(\pi_{0}-\hat{\pi}_{0})+\frac{\pi_{1}-\hat{\pi}_{1}}{\hat{\pi}_{1}}(\mu_{1}-\hat{\mu}_{1})\hat{\pi}_{0}\right\}]$
where (1) applies iterated expectations, (2) adds/subtracts $\hat{\pi}_{0}\mu_{1}$, and the final equality re-arranges. The result then follows by applying the assumption of bounded propensity score estimator and applying the Cauchy-Schwarz inequality. ∎
We are now ready to prove Proposition 4.1. To prove the first claim, consider our proposed estimator of the upper bound $\widehat{\overline{\mbox{perf}}}(s;\beta)$. Let
$\overline{\mbox{perf}}_{i}=\beta_{0,i}+\beta_{1,i}\phi_{\mu,i}(\eta)+\beta_{1,i}\left(1\{\beta_{1,i}>0\}(\overline{\Gamma}-1)+1\{\beta_{1,i}\leq 0\}(\underline{\Gamma}-1)\right)\phi_{\pi\mu,i}(\eta)$
$\widehat{\overline{\mbox{perf}}}_{i}=\beta_{0,i}+\beta_{1,i}\phi_{\mu,i}(\hat{\eta}_{-K_{i}})+\beta_{1,i}\left(1\{\beta_{1,i}>0\}(\overline{\Gamma}-1)+1\{\beta_{1,i}\leq 0\}(\underline{\Gamma}-1)\right)\phi_{\pi\mu,i}(\hat{\eta}_{-K_{i}}).$
Observe $|\widehat{\overline{\mbox{perf}}}(s;\beta)-\overline{\mbox{perf}}(s;\beta)|$ can be written as
$|\mathbb{E}_{n}[\widehat{\overline{\mbox{perf}}}_{i}]-\mathbb{E}[\overline{\mbox{perf}}_{i}]|\leq|\underbrace{\mathbb{E}_{n}[\overline{\mbox{perf}}_{i}]-\mathbb{E}[\overline{\mbox{perf}}_{i}]}_{(a)}|+|\underbrace{\mathbb{E}_{n}\left[\left(\widehat{\overline{\mbox{perf}}}_{i}-\overline{\mbox{perf}}_{i}\right)\right]}_{(b)}|.$
By Chebyshev’s inequality, (a) is $O_{\mathbb{P}}(1/\sqrt{n})$. Next, we can further rewrite (b) as
$\left|\mathbb{E}_{n}\left[\left(\widehat{\overline{\mbox{perf}}}_{i}-\overline{\mbox{perf}}_{i}\right)\right]\right|=\left|\sum_{k=1}^{K}\mathbb{E}_{n}[1\{K_{i}=k\}]\mathbb{E}_{n}^{k}[\widehat{\overline{\mbox{perf}}}_{i,-k}-\overline{\mbox{perf}}_{i}]\right|\leq\sum_{k=1}^{K}\left|\mathbb{E}_{n}^{k}[\widehat{\overline{\mbox{perf}}}_{i,-k}-\overline{\mbox{perf}}_{i}]\right|.$
We will show that each term in the sum is $O_{\mathbb{P}}(1/\sqrt{n}+R_{1,n}^{k}+R_{1,n}^{k}/\sqrt{n})$. Observe that
$\mathbb{E}_{n}^{k}[\widehat{\overline{\mbox{perf}}}_{i,-k}-\overline{\mbox{perf}}_{i}]|\leq|\mathbb{E}_{n}^{k}[\widehat{\overline{\mbox{perf}}}_{i,-k}-\overline{\mbox{perf}}_{i}]-\mathbb{E}[\widehat{\overline{\mbox{perf}}}_{i,-k}-\overline{\mbox{perf}}_{i}\mid\mathcal{O}_{-k}]|+|\mathbb{E}[\widehat{\overline{\mbox{perf}}}_{i,-k}-\overline{\mbox{perf}}_{i}\mid\mathcal{O}_{-k}]|,$
where
$\widehat{\overline{\mbox{perf}}}_{i,-k}-\overline{\mbox{perf}}_{i}=\beta_{1,i}(\phi_{\mu,i}(\hat{\eta}_{-k})-\phi_{\mu,i}(\eta))+\tilde{\beta}_{1,i}\left(\phi_{\pi\mu,i}(\hat{\eta}_{-k})-\phi_{\pi\mu,i}(\eta)\right)$
for $\tilde{\beta}_{i}=\beta_{1,i}\left(1\{\beta_{1,i}>0\}(\overline{\Gamma}-1)+1\{\beta_{1,i}\leq 0\}(\underline{\Gamma}-1)\right)$. So $|\mathbb{E}_{n}^{k}[\widehat{\overline{\mbox{perf}}}_{i,-k}-\overline{\mbox{perf}}_{i}]-\mathbb{E}[\widehat{\overline{\mbox{perf}}}_{i,-k}-\overline{\mbox{perf}}_{i}\mid\mathcal{O}_{-k}]|$ is bounded by
$\underbrace{|\mathbb{E}_{n}^{k}[\beta_{1,i}\left(\phi_{\mu,i}(\hat{\eta}_{-k})-\phi_{\mu,i}(\eta)\right)]-\mathbb{E}[\beta_{1,i}\left(\phi_{\mu,i}(\hat{\eta}_{-k})-\phi_{\mu,i}(\eta)\right)\mid\mathcal{O}_{-k}]|}_{(c)}+$
$\underbrace{|\mathbb{E}_{n}^{k}[\tilde{\beta}_{1,i}\left(\phi_{\pi\mu,i}(\hat{\eta}_{-k})-\phi_{\pi\mu,i}(\eta)\right)]-\mathbb{E}[\tilde{\beta}_{1,i}\left(\phi_{\pi\mu,i}(\hat{\eta}_{-k})-\phi_{\pi\mu,i}(\eta)\right)\mid\mathcal{O}_{-k}]|}_{(d)},$
where (c) is $O_{\mathbb{P}}(\|\hat{\pi}_{1,-k}-\pi_{1}\|/\sqrt{n}+\|\hat{\mu}_{1,-k}+\mu_{1}\|/\sqrt{n}+R_{1,n}^{k}/\sqrt{n})$ by Lemma B.3 and Lemma B.4, and (d) is also $O_{\mathbb{P}}(\|\hat{\pi}_{1,-k}-\pi_{1}\|/\sqrt{n}+\|\hat{\mu}_{1,-k}+\mu_{1}\|/\sqrt{n}+R_{1,n}^{k}/\sqrt{n})$ Lemma B.3 and Lemma B.5. The second term $\mathbb{E}[\widehat{\overline{\mbox{perf}}}_{i,-k}-\overline{\mbox{perf}}_{i}\mid\mathcal{O}_{k}]|$ is bounded by
$\underbrace{|\mathbb{E}[\beta_{1,i}\left(\phi_{\mu,i}(\hat{\eta}_{-k})-\phi_{\mu,i}(\eta)\right)\mid\mathcal{O}_{-k}]|}_{(e)}+|\underbrace{\mathbb{E}[\tilde{\beta}_{1,i}\left(\phi_{\pi\mu,i}(\hat{\eta}_{-k})-\phi_{\pi\mu,i}(\eta)\right)\mid\mathcal{O}_{-k}]}_{(f)}|,$
where (e) is $O_{\mathbb{P}}(R_{1,n}^{k})$ by Lemma B.1 and (f) is $O_{\mathbb{P}}(R_{1,n}^{k})$ by Lemma B.6. Putting this together, we have shown the first claim
$\left|\widehat{\overline{\mbox{perf}}}(s;\beta)-\overline{\mbox{perf}}(s;\beta,\Delta)\right|=$
$O_{\mathbb{P}}\left(1/\sqrt{n}+\sum_{k=1}^{K}R_{1,n}^{k}+\sum_{k=1}^{K}R_{1,n}^{k}/\sqrt{n}+\sum_{k=1}^{K}\left(\|\hat{\pi}_{1,-k}-\pi_{1}\|+\|\hat{\mu}_{1,-k}+\mu_{1}\|\right)/\sqrt{n}\right).$
The result for $\widehat{\underline{\mbox{perf}}}(s;\beta)$ follows the same argument.
The second claim follows by noticing that the proof of the first claim showed that
$\sqrt{n}\left(\begin{pmatrix}\widehat{\overline{\mbox{perf}}}(s;\beta)\\ \widehat{\underline{\mbox{perf}}}(s;\beta)\end{pmatrix}-\begin{pmatrix}\overline{\mbox{perf}}(s;\beta)\\ \underline{\mbox{perf}}(s;\beta)\end{pmatrix}\right)=\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\begin{pmatrix}\overline{\mbox{perf}}_{i}-\mathbb{E}[\overline{\mbox{perf}}_{i}]\\ \underline{\mbox{perf}}_{i}-\mathbb{E}[\underline{\mbox{perf}}_{i}]\end{pmatrix}+o_{\mathbb{P}}(1)$
if $R_{1,n}=o_{\mathbb{P}}(1/\sqrt{n})$. By the central limit theorem,
$\sqrt{n}\left(\begin{pmatrix}\widehat{\overline{\mbox{perf}}}(s;\beta)\\ \widehat{\underline{\mbox{perf}}}(s;\beta)\end{pmatrix}-\begin{pmatrix}\overline{\mbox{perf}}(s;\beta)\\ \underline{\mbox{perf}}(s;\beta)\end{pmatrix}\right)\xrightarrow{d}N\left(0,Cov\left(\begin{pmatrix}\overline{\mbox{perf}}_{i}\\ \underline{\mbox{perf}}_{i}\end{pmatrix}\right)\right)$,
from which the result follows. $\Box$
B.6 Proof of Proposition 4.2
To prove Proposition 4.2, we first consider the case in which $\underline{\delta}(\cdot;\eta),\overline{\delta}(\cdot;\eta)$ are known as a stepping stone. We prove the result for the upper bound since the argument is identical for the lower bound. To simplify notation, we assume the data are split into two equal sized folds.
Define
$\widehat{\overline{\mbox{perf}}}_{+}(s;\beta,\Delta_{n}):=\max_{\tilde{\delta}\in\Delta_{n}}\frac{\mathbb{E}_{n}[\beta_{0,i}\phi_{\mu,i}(\hat{\eta})+\beta_{0,i}\tilde{\delta}_{i}]}{\mathbb{E}_{n}[\phi_{\mu,i}(\hat{\eta})+\tilde{\delta}_{i}]},$ (11)
where $\Delta_{n}=\left\{\tilde{\delta}\colon(1-D_{i})\underline{\delta}(X_{i};\eta)\leq\tilde{\delta}_{i}\leq(1-D_{i})\overline{\delta}(X_{i};\eta)\mbox{ for }i=1,\ldots,n/2\right\}$ and $\underline{\delta}(X_{i};\eta)=(\underline{\Gamma}-1)\mu_{1}(X_{i})$, $\overline{\delta}(X_{i};\eta)=(\overline{\Gamma}-1)\mu_{1}(X_{i})$. As shorthand, define $\underline{\delta}^{\prime}(D_{i},X_{i};\eta)=(1-D_{i})\underline{\delta}(X_{i};\eta)$ and $\overline{\delta}^{\prime}(D_{i},X_{i};\eta)=(1-D_{i})\overline{\delta}(X_{i};\eta)$.
Lemma B.7.
[…]
$\widehat{\overline{\mbox{perf}}}_{+}(s;\beta,\Delta_{n}):=\max_{\tilde{\delta}\in\Delta_{n}^{M}}\frac{\mathbb{E}_{n}[\beta_{0,i}\phi_{\mu,i}(\hat{\eta})+\beta_{0,i}\tilde{\delta}_{i}]}{\mathbb{E}_{n}[\phi_{\mu,i}(\hat{\eta})+\tilde{\delta}_{i}]}.$
Proof.
We first show this result for the fold-specific estimator $\widehat{\overline{\mbox{perf}}}_{+}(s;\beta,\Delta_{n})$ by using the proof strategy of Proposition 2 in KallusZhou(21) By Lemma C.5,
$\widehat{\overline{\mbox{perf}}}_{+}(s;\beta,\Delta_{n})=\max_{\tilde{U},\tilde{V}}$ $\hat{\alpha}^{\prime}\tilde{U}+\hat{c}\tilde{V}$
s.t. $0\leq\tilde{U}_{i}\leq\tilde{V}\mbox{ for }i=1,\ldots n/2,\mbox{ }0\leq\tilde{V},\mbox{ }\hat{\gamma}^{\prime}\tilde{U}+\tilde{V}\hat{d}=1.$
Next, define the dual program associated with this primal linear program. Let $P_{i}$ be the dual variables associated with the constraints $\tilde{U}_{i}\leq\tilde{V}$, $Q_{i}$ be the dual variables associated with the constraints $\tilde{U}_{i}\geq 0$, and $\lambda$ be the dual variable associated with the constraint $\hat{\gamma}^{\prime}\tilde{U}+\tilde{V}\hat{d}=1$. The dual linear program is
$\min_{\lambda,P,Q}$ $\lambda$
s.t. $P_{i}-Q_{i}+\lambda\hat{\gamma}_{i}=\hat{\alpha}_{i},\mbox{ }-\textbf{1}^{\prime}P+\lambda\hat{d}\geq\hat{c}$,
$P_{i}\geq 0,Q_{i}\geq 0\mbox{ for }i=1,\ldots,n/2,$
where 1 is the vector of all ones of appropriate dimension. By re-arranging the first constraint and substituting in the expressions for $\hat{\alpha}_{i},\hat{\gamma}_{i}$, observe that
$P_{i}-Q_{i}=(\beta_{0}-\lambda)(\overline{\delta}^{\prime}_{i}-\underline{\delta}^{\prime}_{i}).$
By complementary slackness, at most only one of $P_{i}$ or $Q_{i}$ will be non-zero at the optimum, and so combined with the previous display, this implies
$P_{i}=\max\{\beta_{0,i}-\lambda,0\}(\overline{\delta}^{\prime}_{i}-\underline{\delta}^{\prime}_{i}),$
$Q_{i}=\max\{\lambda-\beta_{0,i},0\}(\overline{\delta}^{\prime}_{i}-\underline{\delta}^{\prime}_{i}).$
Next, notice that the constraint $\textbf{1}^{\prime}P+\lambda\hat{d}\geq\hat{c}$ must be tight at the optimum. Plugging in the previous expression for $P_{i}$ and the expressions for $\hat{c},\hat{d}$, this implies that $\lambda$ satisfies
$-\mathbb{E}_{n}[\max\{\beta_{0,i}-\lambda,0\}(\overline{\delta}^{\prime}_{i}-\underline{\delta}^{\prime}_{i})]=\mathbb{E}_{n}[\left(\beta_{0,i}-\lambda\right)\left(\phi_{\mu,i}(\hat{\eta})+\underline{\delta}^{\prime}_{i}\right)].$
Finally, we consider three separate cases:
  1. 1.
    Suppose that $\lambda\geq\max_{i}\beta_{0,i}$. From the previous display, $\lambda$ must satisfy
    $0=\mathbb{E}_{n}[\left(\beta_{0,i}-\lambda\right)\left(\phi_{\mu,i}(\hat{\eta})+\underline{\delta}^{\prime}_{i}\right)]\implies\lambda=\frac{\mathbb{E}_{n}[\beta_{0,i}\left(\phi_{\mu,i}(\hat{\eta})+\underline{\delta}^{\prime}_{i}\right)]}{\mathbb{E}_{n}[\phi_{\mu,i}(\hat{\eta})+\underline{\delta}^{\prime}_{i}]}.$
    At this value for $\lambda$, the expressions for $P_{i},Q_{i}$ imply that $P_{i}=0$, $Q_{i}>0$ for all $i$. By complementary slackness, this in turn implies that $\tilde{U}_{i}=0$, or equivalently $U_{i}=0$ for all $i$.
  2. 2.
    Suppose that $\lambda\leq\min_{i}\beta_{0,i}$. From the previous display, $\lambda$ must satisfy
    $\mathbb{E}_{n}[(\beta_{0,i}-\lambda)(\overline{\delta}^{\prime}_{i}-\underline{\delta}^{\prime}_{i})]=\mathbb{E}_{n}[\left(\beta_{0,i}-\lambda\right)\left(\phi_{\mu,i}(\hat{\eta})+\underline{\delta}^{\prime}_{i}\right)]\implies\lambda=\frac{\mathbb{E}_{n}[\beta_{0,i}\left(\phi_{\mu,i}(\hat{\eta})+\overline{\delta}^{\prime}_{i}\right)]}{\mathbb{E}_{n}[\phi_{\mu,i}(\hat{\eta})+\overline{\delta}^{\prime}_{i}]}.$
    At this value for $\lambda$, the expressions for $P_{i},Q_{i}$ imply that $P_{i}>0$, $Q_{i}=0$ for all $i$. By complementary slackness, this implies that $\tilde{U}_{i}=\tilde{V}$, or equivalently $U_{i}=1$ for all $i$.
  3. 3.
    Suppose that $\min_{i}\beta_{0,i}<\lambda<\max_{i}\beta_{0,i}$. Then, $\beta_{0,(j)}<\lambda\leq\beta_{0,(j+1)}$ for some $j$ where $\beta_{0,(1)},\ldots,\beta_{0,(n_{k})}$ are the order statistics of the sample outcomes. The expressions for $P_{i},Q_{i}$ in turn imply that $Q_{i}>0$ only when $\beta_{0,i}\leq\beta_{0,(j)}$ (in which case $U_{i}=0$) and $P_{i}>0$ only when $\beta_{0,i}\geq\beta_{0,(j+1)}$ (in which case $U_{i}=1$).
Therefore, in all three cases, the optimal solution is such that there exists a non-decreasing function $u(\cdot)\colon\mathbb{R}\rightarrow[0,1]$ such that $U_{i}=u(\beta_{0,i})$ attains the upper bound.
We next prove the result for the population bound $\overline{\mbox{perf}}_{+}(s;\beta,\Delta)$ via a similar argument. We can first rewrite the bound as
$\overline{\mbox{perf}}_{+}(s;\beta,\Delta):=\sup_{\delta(\cdot)\in\Delta}\mathbb{E}[\mu_{1}(X_{i})+(1-D_{i})\delta(X_{i})]^{-1}\mathbb{E}[\beta_{0,i}\mu_{1}(X_{i})+\beta_{0,i}(1-D_{i})\delta(X_{i})]$
by iterated expectations. Then, applying the change-of-variables $\delta(X_{i})=\underline{\delta}(X_{i})+(\overline{\delta}(X_{i})-\underline{\delta}(X_{i}))U(X_{i})$, we further rewrite the population bound as
$\overline{\mbox{perf}}_{+}(s;\beta,\Delta):=\sup_{U(\cdot)\colon\mathcal{X}\rightarrow[0,1]}\frac{\mathbb{E}[\beta_{0,i}\mu_{1}(X_{i})+\beta_{0,i}\pi_{0}(X_{i})\underline{\delta}_{i}+\beta_{0,i}\pi_{0}(X_{i})(\overline{\delta}_{i}-\underline{\delta}_{i})U(X_{i})]}{\mathbb{E}[\mu_{1}(X_{i})+\pi_{0}(X_{i})\underline{\delta}_{i}+\pi_{0}(X_{i})(\overline{\delta}_{i}-\underline{\delta}_{i})U(X_{i})]}.$
Define $c:=\mathbb{E}[\beta_{0,i}\mu_{1}(X_{i})+\beta_{0,i}\pi_{0}(X_{i})\underline{\delta}_{i}]$, $d:=\mathbb{E}[\mu_{1}(X_{i})+\pi_{0}(X_{i})\underline{\delta}_{i}]$, and $\alpha(x):=\beta_{0}(x;s)\pi_{0}(x)(\overline{\delta}(x)-\underline{\delta}(x))$, $\gamma(x):=\pi_{0}(x)(\overline{\delta}(x)-\underline{\delta}(x))$. Letting $\langle f,g\rangle_{P(\cdot)}$ denote the inner product $\mathbb{E}[f(X_{i})g(X_{i})]$, we can further rewrite the population bound as
$\overline{\mbox{perf}}_{+}(s;\beta,\Delta):=\sup_{U(\cdot)\colon\mathcal{X}\rightarrow[0,1]}\frac{c+\langle\alpha,U\rangle_{P(\cdot)}}{d+\langle\gamma,U\rangle_{P(\cdot)}}.$
Define the change-of-variables $\tilde{U}(\cdot)=\frac{U(\cdot)}{d+\langle\gamma,U\rangle_{P(\cdot)}}$ and $\tilde{V}=\frac{1}{\langle\gamma,U\rangle_{P(\cdot)}}$. The previous linear-fractional optimization is equivalent to
$\sup_{\tilde{U}(\cdot),\tilde{V}}$ $\langle\alpha,\tilde{U}\rangle_{P(\cdot)}+c\tilde{V}\mbox{ s.t. }0\leq\tilde{U}(x)\leq\tilde{V}\mbox{ for all }x\in\mathcal{X},\mbox{ }\langle\gamma,\tilde{U}\rangle_{P(\cdot)}+\tilde{V}d=1.$
Define the dual associated with this primal program. Let $\tilde{P}(x)$ be the dual function associated with the constraint $\tilde{U}(x)\leq\tilde{V}$, $\tilde{Q}(x)$ be the dual variables associated with the constraints $\tilde{U}(x)\geq 0$, and $\lambda$ be the dual variable associated with the constraint $\langle\gamma,\tilde{U}\rangle_{P(\cdot)}+\tilde{V}d=1$. The dual is
$\inf_{\lambda,\tilde{P}(\cdot),\tilde{Q}(\cdot)}$ $\lambda$
s.t. $\tilde{P}(x)-\tilde{Q}(x)+\lambda\gamma(x)=\alpha(x)\mbox{ for all }x\in\mathcal{X}$
$-\langle\textbf{1},\tilde{P}\rangle_{P(\cdot)}+\lambda d\geq c,\mbox{ }\tilde{P}(x)\geq 0,\tilde{Q}(x)\geq 0\mbox{ for all }x\in\mathcal{X}.$
By complementary slackness, at most only one of $\tilde{P}(x)$ or $\tilde{Q}(x)$ can be non-zero at the optimum for all $x\in\mathcal{X}$. Therefore, by re-arranging the first constraint and substituting in for $\alpha(x),\gamma(x)$, observe
$\tilde{P}(x)-\tilde{Q}(x)=(\beta_{0}(x)-\lambda)\pi_{0}(x)(\overline{\delta}(x)-\underline{\delta}(x)),$
which in turn implies that
$\tilde{P}(x)=\max\{\beta_{0}(x)-\lambda,0\}\pi_{0}(x)(\overline{\delta}(x)-\underline{\delta}(x)),$
$\tilde{Q}(x)=\max\{\lambda-\beta_{0}(x),0\}\pi_{0}(x)(\overline{\delta}(x)-\underline{\delta}(x)).$
[…]
$-\mathbb{E}[\max\{\beta_{0}(X_{i})-\lambda,0\}\pi_{0}(X_{i})(\overline{\delta}(X_{i})-\underline{\delta}(X_{i}))]=\mathbb{E}[(\beta_{0}(X_{i})-\lambda)(\mu_{1}(X_{i})+\pi_{0}(X_{i})\underline{\delta}(X_{i})].$
As in the proof for the estimator, we can consider three cases: (i) $\lambda\geq\overline{\beta}_{0}$, (ii) $\lambda\leq\underline{\beta}_{0}$ and (iii) $\underline{\beta}_{0}<\lambda<\overline{\beta}_{0}$ for $\underline{\beta}_{0}:=\inf_{x\in\mathcal{X}}\beta_{0}(x)$, $\overline{\beta}_{0}=\sup_{x\in\mathcal{X}}\beta_{0}(x)$. In each case, the optimal solution is such that there exists a non-decreasing function $u(\cdot)\colon\mathbb{R}\rightarrow[0,1]$ such that $U(x)=u(\beta_{0}(x))$ attains the upper bound. The result follows by applying the definitions of $\underline{\delta}^{\prime}(d,x;\eta),\overline{\delta}^{\prime}(d,x;\eta)$. ∎
Lemma B.8.
Define $R_{1,n}=\|\hat{\mu}_{1}(\cdot)-\mu_{1}(\cdot)\|\|\hat{\pi}_{1}(\cdot)-\pi_{1}(\cdot)\|$. Assume that (i) there $\delta>0$ such that $\mathbb{P}(\pi_{1}(X_{i})\geq\delta)=1$; (ii) there exists $\epsilon>0$ such that $\mathbb{P}(\hat{\pi}(X_{i})\geq\epsilon)=1$; and (iii) $\|\hat{\mu}_{1}(\cdot)-\mu_{1}(\cdot)\|=o_{P}(1)$ and $\|\hat{\pi}_{1}(\cdot)-\pi_{1}(\cdot)\|=o_{P}(1)$. Then,
$\left\|\widehat{\overline{\mbox{perf}}}_{+}(s;\beta,\Delta_{n})-\overline{\mbox{perf}}_{+}(s;\beta,\Delta)\right\|=O_{\mathbb{P}}\left(1/\sqrt{n}+R_{1,n}\right).$
Proof.
Let $\widehat{\mbox{perf}}_{+}(s;\beta,\tilde{\delta}):=\mathbb{E}_{n}[\beta_{0,i}\phi_{\mu,i}(\hat{\eta})+\beta_{0,i}\tilde{\delta}_{i}]/\mathbb{E}_{n}[\phi_{\mu,i}(\hat{\eta})+\tilde{\delta}_{i}]$ for $\tilde{\delta}_{i}=(1-D_{i})\delta_{i}$. To prove this result, we first observe that
[…]
where the first equality uses Lemma B.7. For any $\tilde{\delta}\in\Delta^{M}$, we have that
[…]
$\frac{\mathbb{E}_{n}[\beta_{0,i}\phi_{\mu,i}(\hat{\eta})+\beta_{0,i}\tilde{\delta}_{i}]}{\mathbb{E}_{n}[\phi_{\mu,i}(\hat{\eta})+\tilde{\delta}_{i}]}-\frac{\mathbb{E}[\beta_{0,i}\phi_{\mu,i}(Y_{i};\hat{\eta})+\beta_{0,i}\tilde{\delta}_{i}]}{\mathbb{E}[\phi_{\mu,i}(\hat{\eta})+\tilde{\delta}_{i}]}=$
$\frac{\mathbb{E}_{n}[(\text{\#1})]}{\mathbb{E}_{n}[(\text{\#2})]}-\frac{\mathbb{E}[(\text{\#3})]}{\mathbb{E}[(\text{\#4})]}=\mathbb{E}_{n}[(\text{\#2})]^{-1}\left\{\mathbb{E}_{n}[(\text{\#1})]-\mathbb{E}[(\text{\#3})]-\frac{\mathbb{E}[(\text{\#3})]}{\mathbb{E}[(\text{\#4})]}(\mathbb{E}_{n}[(\text{\#2})]-\mathbb{E}[(\text{\#4})])\right\},$
where
$\mathbb{E}_{n}[(\text{\#1})]-\mathbb{E}[(\text{\#3})]$ $=\mathbb{E}_{n}[\beta_{0,i}\phi_{\mu,i}(\hat{\eta})+\beta_{0,i}\tilde{\delta}_{i}]-\mathbb{E}[\beta_{0,i}\phi_{\mu,i}(\eta)+\beta_{0,i}\tilde{\delta}_{i}]$
$=\left(\mathbb{E}_{n}[\beta_{0,i}\phi_{\mu,i}(\hat{\eta})]-\mathbb{E}[\beta_{0,i}\phi_{\mu,i}(\eta)]\right)+(\mathbb{E}_{n}-\mathbb{E})[\beta_{0,i}\tilde{\delta}_{i}]$
$\mathbb{E}_{n}[(\text{\#2})]-\mathbb{E}[(\text{\#4})]$ $=\mathbb{E}_{n}[\phi_{\mu,i}(\hat{\eta})+\tilde{\delta}_{i}]-\mathbb{E}[\phi_{\mu,i}(\eta)+\tilde{\delta}_{i}]$
$=\left(\mathbb{E}_{n}[\phi_{\mu,i}(\hat{\eta})]-\mathbb{E}[\phi_{\mu,i}(\eta)]\right)+(\mathbb{E}_{n}-\mathbb{E})[\tilde{\delta}_{i}].$
Observe that
$\mathbb{E}_{n}[(\text{\#2})]=\mathbb{E}_{n}[\phi_{\mu,i}(\hat{\eta})+\tilde{\delta}_{i}]\geq\mathbb{E}_{n}[\phi_{\mu,i}(\hat{\eta})+(1-D_{i})\underline{\delta}_{i}]$
$\mathbb{E}[(\text{\#3})]=\mathbb{E}[\beta_{0,i}\phi_{\mu,i}(\eta)+\beta_{0,i}\tilde{\delta}_{i}]\leq\mathbb{E}[\beta_{0,i}\phi_{\mu,i}(\eta)+\beta_{0,i}(1-D_{i})\overline{\delta}_{i}]$
$\mathbb{E}[(\text{\#4})]=\mathbb{E}[\mu_{1}(X_{i})+\tilde{\delta}_{i}]\geq\mathbb{E}[\mu_{1}(X_{i})+(1-D_{i})\underline{\delta}_{i}].$
Therefore, there exists $C_{1}>0$ such that $\mathbb{E}_{n}[(\text{\#2})]>C_{1}$ for all $n$ under the assumption of bounded nuisance parameter estimates. There also exists constants $C_{2}<\infty$, $C_{3}>0$ such that $\mathbb{E}[(\text{\#3})]<C_{2}$ and $\mathbb{E}[(\text{\#4})]>C_{3}$. Putting this together, we therefore have
$\left\|\widehat{\overline{\mbox{perf}}}(s;\beta,\Delta_{n})-\overline{\mbox{perf}}_{+}(s;\beta,\Delta)\right\|\leq$
$C_{1}\left\|\underbrace{\mathbb{E}_{n}[\beta_{0,i}\phi_{\mu,i}(\hat{\eta})]-\mathbb{E}[\beta_{0,i}\phi_{\mu,i}(\eta)]}_{(a)}\right\|+C_{1}+\left\|\underbrace{\mathbb{E}_{n}[\beta_{0,i}\underline{\delta}_{i}^{\prime}]-\mathbb{E}[\beta_{0,i}\underline{\delta}_{i}^{\prime}]}_{(b)}+C_{1}\right\|\underbrace{\sup_{U\in\mathcal{U}}\left\|(\mathbb{E}_{n}-\mathbb{E})[\beta_{0,i}(\overline{\delta}_{i}^{\prime}-\underline{\delta}_{i}^{\prime})U(\beta_{0,i})]\right\|}_{(c)}+$
$C_{1}\frac{C_{2}}{C_{3}}\left\|\underbrace{\mathbb{E}_{n}[\phi_{\mu,i}(\hat{\eta})]-\mathbb{E}[\phi_{\mu,i}(\eta)]}_{(d)}\right\|+C_{1}\frac{C_{2}}{C_{3}}\left\|\underbrace{\mathbb{E}_{n}[\underline{\delta}_{i}]-\mathbb{E}[\underline{\delta}_{i}]}_{(e)}\right\|+C_{1}\frac{C_{2}}{C_{3}}\underbrace{\sup_{U\in\mathcal{U}}\left\|(\mathbb{E}_{n}-\mathbb{E})[(\overline{\delta}_{i}^{\prime}-\underline{\delta}_{i}^{\prime})U(\beta_{0,i})\right\|}_{(f)}.$
We analyze each term separately. The proof of Proposition 4.1 establishes that (a), (d) are $O_{\mathbb{P}}(1/\sqrt{n}+R_{1,n})$. Moreover, (b), (e) are $O_{\mathbb{P}}(1/\sqrt{n})$ by standard arguments. Consider (c), which we may write out as
$\sup_{U\in\mathcal{U}}\left|(n/2)^{-1}\sum_{i}\beta_{0,i}(1-D_{i})(\overline{\delta}_{i}-\underline{\delta}_{i})U(\beta_{0,i})-\mathbb{E}[\beta_{0,i}(1-D_{i})(\overline{\delta}_{i}-\underline{\delta}_{i})U(\beta_{0,i})]\right|.$
Define f(w,x,y,z)=w(yx)zf(w,x,y,z)=w(y-x)z and $\mathcal{F}=\{f_{U}\}_{U\in\mathcal{U}}$ to be the class of functions $f_{U}\colon(d,\underline{\delta},\overline{\delta},\beta)\rightarrow(1-d)(\overline{\delta}-\underline{\delta})\beta u(\beta)$. Observe $f$ is a contraction over its final argument on $[0,1]$. We can then rewrite (c) as
$\sup_{f_{U}\in\mathcal{F}}\left|(n/2)^{-1}\sum_{i}f_{U}(D_{i},\underline{\delta}_{i},\overline{\delta}_{i},\beta_{0,i})-\mathbb{E}[f_{\tilde{\delta}}(D_{i},\underline{\delta}_{i},\overline{\delta}_{i},\beta_{0,i})]\right|.$
Applying a standard concentration inequality (e.g., Theorem 4.10 in Wainwright_textbook), observe that, with probability at least $1-\delta$,
$\sup_{f_{U}\in\mathcal{F}}\left|(n/2)^{-1}\sum_{i}f_{U}(D_{i},\underline{\delta}_{i},\overline{\delta}_{i},\beta_{0,i})-\mathbb{E}[f_{U}(D_{i},\underline{\delta}_{i},\overline{\delta}_{i},\beta_{0,i})]\right|\leq R_{n}(\mathcal{F})+\sqrt{\frac{2\log(1/\delta)}{n/2}},$
where $R_{n}(\mathcal{F})$ is the Rademacher complexity of $\mathcal{F}$. Now we relate $R_{n}(\mathcal{F})$ to $R_{n}(\mathcal{U})$. For any fixed $(d_{1},\underline{\delta}_{1},\overline{\delta}_{i},\beta_{0,1}),\ldots,(d_{n/2},\underline{\delta}_{n/2},\overline{\delta}_{n/2},\beta_{0,n/2})$, observe that
$\mathbb{E}_{\epsilon}[\sup_{U\in\mathcal{U}}\left|\sum_{i=1}^{n/2}\epsilon_{i}f_{U}(d_{i},\underline{\delta}_{i},\overline{\delta}_{i},\beta_{0,i})\right|]$ $=\mathbb{E}_{\epsilon}[\sup_{U\in\mathcal{U}}\left|\sum_{i=1}^{n/2}\epsilon_{i}\beta_{0,i}(1-d_{i})(\overline{\delta}_{i}-\underline{\delta}_{i})U(\beta_{0,i})\right|]$
$\leq 2(\overline{\Gamma}-\underline{\Gamma})\mathbb{E}_{\epsilon}[\sup_{U\in\mathcal{U}}\left|\sum_{i=1}^{n/2}\epsilon_{i}U(\beta_{0,i})\right|]$
where we used that $(1-d_{i})(\overline{\delta}_{i}-\underline{\delta}_{i})\leq(\overline{\Gamma}-\underline{\Gamma})$ and the Ledoux-Talagrand contraction inequality (e.g., Eq. (5.61) in Wainwright_textbook). Dividing by $n/2$ and averaging over the observations yields $R_{n}(\mathcal{F})\leq 2(\overline{\Gamma}-\underline{\Gamma})R_{n}(\mathcal{U})$. Finally, we can bound the Rademacher complexity of $\mathcal{U}$ using Dudley’s entropy integral (e.g., Theorem 5.22 in Wainwright_textbook) as
$R_{n}(\mathcal{U})\leq\frac{C}{\sqrt{n/2}}\int_{0}^{1}\sqrt{\log(N(\xi,\mathcal{U},\|\cdot\|_{\mathbb{P}_{n}}))}d\xi\leq\frac{C}{\sqrt{n/2}}\int_{0}^{1}\sqrt{\log(N_{[]}(2\xi,\mathcal{U},\|\cdot\|_{\mathbb{P}_{n}}))}d\xi,$
for some constant $C$, where $N(\xi,\mathcal{U},\|\cdot\|_{\mathbb{P}_{n}})$ is the covering number and $N_{[]}(2\xi,\mathcal{U},\|\cdot\|_{p_{n}})$ is the bracketing number. But, Theorem 2.7.5 of VDVWellner establishes that the bracketing entropy $\log(N_{[]}(\xi,\mathcal{U},\|\cdot\|_{\mathbb{P}_{n}})$ of the class of monotone non-decreasing functions is bounded by $(1/\xi)\log(1/\xi)$, and so $\int_{0}^{1}\sqrt{\log(N_{[]}(\xi,\mathcal{U},\|\cdot\|_{\mathbb{P}_{n}}))}d\xi=\sqrt{2\pi}$. It follows that, for any $\delta>0$,
$\sup_{U\in\mathcal{U}}\left\|(\mathbb{E}_{n}^{k}-\mathbb{E})[\beta_{0,i}(\overline{\delta}_{i}^{\prime}-\underline{\delta}_{i}^{\prime})U(\beta_{0,i})]\right\|\leq\frac{C}{\sqrt{n/2}}+\sqrt{\frac{2\log(1/\delta)}{2n/2}}$
holds with probability $1-\delta$. We therefore conclude that (c) is $O_{\mathbb{P}}(1/\sqrt{n})$. Similarly, (f) is $O_{\mathbb{P}}(1/\sqrt{n})$ by the same argument. ∎
[…]
$\widehat{\overline{\mbox{perf}}}_{+}(s;\beta,\hat{\Delta}_{n}):=\max_{\tilde{\delta}\in\hat{\Delta}_{n}}\frac{\mathbb{E}_{n}[\beta_{0,i}\phi_{\mu,i}(\hat{\eta})+\beta_{0,i}\tilde{\delta}_{i}]}{\mathbb{E}_{n}[\phi_{\mu,i}(\hat{\eta})+\tilde{\delta}_{i}]},$
where $\hat{\Delta}_{n}:=\left\{\delta\in\mathbb{R}^{n}\colon(1-D_{i})\hat{\underline{\delta}}(X_{i})\leq\delta_{i}\leq(1-D_{i})\hat{\overline{\delta}}(X_{i})\mbox{ for }i=1,\ldots,n_{k}\right\}$. Observe that
$\|\widehat{\overline{\mbox{perf}}}_{+}(s;\beta,\hat{\Delta}_{n})-\overline{\mbox{perf}}_{+}(s;\beta,\Delta)\|\leq\|\widehat{\overline{\mbox{perf}}}_{+}(s;\beta,\hat{\Delta}_{n})-\widehat{\overline{\mbox{perf}}}_{+}(s;\beta,\Delta_{n})\|+$
$\|\widehat{\overline{\mbox{perf}}}_{+}(s;\beta,\Delta_{n})-\overline{\mbox{perf}}_{+}(s;\beta,\Delta)\|.$
Lemma B.8 analyzed the convergence rate of the second term. It is therefore sufficient to bound the first term.
Lemma B.9.
Assume the same conditions as Lemma B.8. Then,
$\|\widehat{\overline{\mbox{perf}}}_{+}(s;\beta,\hat{\Delta}_{n})-\widehat{\overline{\mbox{perf}}}_{+}(s;\beta,\Delta_{n})\|\lesssim\sqrt{\frac{1}{n/2}\sum_{i=1}^{n/2}(\widehat{\underline{\delta}}_{i}-\underline{\delta}_{i})^{2}}+\sqrt{\frac{1}{n/2}\sum_{i=1}^{n/2}(\widehat{\overline{\delta}}_{i}-\overline{\delta}_{i})^{2}},$
where $a\lesssim b$ means $a\leq Cb$ for some constant $C$.
Proof.
Applying the change-of-variables in the Proof of Lemma C.3,
$\widehat{\overline{\mbox{perf}}}_{+}(s;\beta,\Delta_{n}):=\max_{0\leq U\leq 1}$ $\frac{\sum_{i=1}^{n/2}\beta_{0,i}\phi_{\mu,i}(\hat{\eta})+\beta_{0,i}(1-D_{i})\underline{\delta}_{i}+\beta_{0,i}(1-D_{i})(\overline{\delta}_{i}-\underline{\delta}_{i})U_{i}}{\sum_{i=1}^{n_{k}}\phi_{\mu,i}(\hat{\eta})+(1-D_{i})\underline{\delta}_{i}+(1-D_{i})(\overline{\delta}_{i}-\underline{\delta}_{i})U_{i}}$
$\widehat{\overline{\mbox{perf}}}_{+}(s;\beta,\hat{\Delta}_{n}):=\max_{0\leq U\leq 1}$ $\frac{\sum_{i=1}^{n/2}\beta_{0,i}\phi_{\mu,i}(\hat{\eta})+\beta_{0,i}(1-D_{i})\hat{\underline{\delta}}_{i}+\beta_{0,i}(1-D_{i})(\hat{\overline{\delta}}_{i}-\hat{\underline{\delta}}_{i})U_{i}}{\sum_{i=1}^{n/2}\phi_{\mu,i}(\hat{\eta})+(1-D_{i})\hat{\underline{\delta}}_{i}+(1-D_{i})(\hat{\overline{\delta}}_{i}-\hat{\underline{\delta}}_{i})U_{i}}.$
We can therefore rewrite
$\|\widehat{\overline{\mbox{perf}}}_{+}(s;\beta,\hat{\Delta}_{n})-\widehat{\overline{\mbox{perf}}}_{+}(s;\beta,\Delta_{n})\|\leq$
$\max_{0\leq U\leq 1}\left\|\frac{\sum_{i=1}^{n/2}\beta_{0,i}\phi_{\mu,i}(\hat{\eta})+\beta_{0,i}(1-D_{i})\hat{\underline{\delta}}_{i}+\beta_{0,i}(1-D_{i})(\hat{\overline{\delta}}_{i}-\hat{\underline{\delta}}_{i})U_{i}}{\sum_{i=1}^{n/2}\phi_{\mu,i}(\hat{\eta})+(1-D_{i})\hat{\underline{\delta}}_{i}+(1-D_{i})(\hat{\overline{\delta}}_{i}-\hat{\underline{\delta}}_{i})U_{i}}-\frac{\sum_{i=1}^{n/2}\beta_{0,i}\phi_{\mu,i}(\hat{\eta})+\beta_{0,i}(1-D_{i})\underline{\delta}_{i}+\beta_{0,i}(1-D_{i})(\overline{\delta}_{i}-\underline{\delta}_{i})U_{i}}{\sum_{i=1}^{n/2}\phi_{\mu,i}(\hat{\eta})+(1-D_{i})\underline{\delta}_{i}+(1-D_{i})(\overline{\delta}_{i}-\underline{\delta}_{i})U_{i}}\right\|=$
$\max_{0\leq U\leq 1}\|\frac{\mathbb{E}_{n}[(\text{\#1})]}{\mathbb{E}_{n}[(\text{\#2})]}-\frac{\mathbb{E}_{n}[(\text{\#3})]}{\mathbb{E}_{n}[(\text{\#4})]}\|=\max_{0\leq U\leq 1}\mathbb{E}_{n}[(\text{\#2})]^{-1}\left\{\underbrace{\left(\mathbb{E}_{n}[(\text{\#1})]-\mathbb{E}_{n}[(\text{\#3})]\right)}_{(a)}-\frac{\mathbb{E}_{n}[(\text{\#3})]}{\mathbb{E}_{n}[(\text{\#4})]}\underbrace{\left(\mathbb{E}_{n}[(\text{\#2})]-\mathbb{E}_{n}[(\text{\#4})]\right)}_{(b)}\right\}.$
Notice that we can rewrite (a), (b) as
$(a)=\mathbb{E}_{n}[\beta_{0,i}(1-D_{i})(\hat{\underline{\delta}}_{i}-\underline{\delta}_{i})(1-U_{i})+\beta_{0,i}(1-D_{i})(\hat{\overline{\delta}}_{i}-\overline{\delta}_{i})U_{i}]$
$(b)=\mathbb{E}_{n}[(1-D_{i})(\hat{\underline{\delta}}_{i}-\underline{\delta}_{i})(1-U_{i})+(1-D_{i})(\hat{\overline{\delta}}_{i}-\overline{\delta}_{i})U_{i}].$
Notice that
$\mathbb{E}_{n}[(\text{\#2})]\geq\mathbb{E}_{n}[\phi_{\mu,i}(\hat{\eta})+(1-D_{i})\underline{\delta}_{i}]\mbox{ for all }n,$
$\mathbb{E}_{n}[(\text{\#3})]\leq\mathbb{E}_{n}[\beta_{0,i}\phi_{\mu,i}(\hat{\eta})+\beta_{0,i}(1-D_{i})\overline{\delta}_{i}]\mbox{ for all }n,$
$\mathbb{E}_{n}[(\text{\#4})]\leq\mathbb{E}_{n}[\phi_{\mu,i}(\hat{\eta})+(1-D_{i})\underline{\delta}_{i}]\mbox{ for all }n.$
So, there exists a constant $0<C_{1}$ such that $\mathbb{E}_{n}[(\text{\#2})]>C_{1}$ for all $n$ under the assumption of bounded nuisance parameter estimators and the assumption on the estimated bounds and there exists constants $0<C_{2}<\infty,C_{3}>0$ such that $\mathbb{E}[(\text{\#3})]<C_{2}$, $\mathbb{E}[(\text{\#4})]>C_{3}$ under the assumption of bounded nuisance parameter estimators. Putting this together implies that
$\|\widehat{\overline{\mbox{perf}}}_{+}(s;\beta,\hat{\Delta}_{n})-\widehat{\overline{\mbox{perf}}}_{+}(s;\beta,\Delta_{n})\|\lesssim$
$\max_{0\leq U\leq 1}\|\mathbb{E}_{n}[\beta_{0,i}(1-D_{i})(\hat{\underline{\delta}}_{i}-\underline{\delta}_{i})(1-U_{i})+\beta_{0,i}(1-D_{i})(\hat{\overline{\delta}}_{i}-\overline{\delta}_{i})U_{i}]\|+\|\mathbb{E}_{n}[(1-D_{i})(\hat{\underline{\delta}}_{i}-\underline{\delta}_{i})(1-U_{i})+(1-D_{i})(\hat{\overline{\delta}}_{i}-\overline{\delta}_{i})U_{i}]\|\leq$
$\max_{0\leq U\leq 1}(n/2)^{-1}\sum_{i=1}^{n/2}\|\beta_{0,i}(1-D_{i})\{(\hat{\underline{\delta}}_{i}-\underline{\delta}_{i})(1-U_{i})+(\hat{\overline{\delta}}_{i}-\overline{\delta}_{i})U_{i}\}\|+(n/2)^{-1}\sum_{i=1}^{n/2}\|(1-D_{i})\{(\hat{\underline{\delta}}_{i}-\underline{\delta}_{i})(1-U_{i})+(\hat{\overline{\delta}}_{i}-\overline{\delta}_{i})U_{i}\}\|\leq$
$\max_{0\leq U\leq 1}(n/2)^{-1}\sum_{i=1}^{n/2}\|(\hat{\underline{\delta}}_{i}-\underline{\delta}_{i})(1-U_{i})+(\hat{\overline{\delta}}_{i}-\overline{\delta}_{i})U_{i}\|+(n/2)^{-1}\sum_{i=1}^{n/2}\|(\hat{\underline{\delta}}_{i}-\underline{\delta}_{i})(1-U_{i})+(\hat{\overline{\delta}}_{i}-\overline{\delta}_{i})U_{i}\}\|\lesssim$
$\mathbb{E}_{n}[|\hat{\underline{\delta}}_{i}-\underline{\delta}_{i}|]+\mathbb{E}_{n}[|\hat{\overline{\delta}}_{i}-\overline{\delta}_{i}|].$
Then, using the inequality $\|v\|_{1}\leq\sqrt{n}\|v\|_{2}$ for $v\in\mathbb{R}^{n}$, it follows that
$\|\widehat{\overline{\mbox{perf}}}_{+}(s;\beta,\hat{\Delta}_{n})-\widehat{\overline{\mbox{perf}}}_{+}(s;\beta,\Delta_{n})\|\lesssim\sqrt{\frac{1}{n/2}\sum_{i=1}^{n/2}(\widehat{\underline{\delta}}_{i}-\underline{\delta}_{i})^{2}}+\sqrt{\frac{1}{n/2}\sum_{i=1}^{n/2}(\widehat{\overline{\delta}}_{i}-\overline{\delta}_{i})^{2}}.$
Putting together Lemma B.8 and Lemma B.9, the result follows. $\Box$
B.7 Proof of Corollary 4.1
To prove this result, we first observe that
$\sqrt{\mathbb{E}_{n}\left[\left(\widehat{\delta}(X_{i})-\mu_{1}(X_{i})\right)^{2}\right]}=\underbrace{\sqrt{\mathbb{E}\left[\left(\widehat{\delta}(X_{i})-\mu_{1}(X_{i})\right)^{2}\right]}}_{:=\left\|\widehat{\delta}(\cdot)-\mu_{1}(\cdot)\right\|}+O_{\mathbb{P}}(1/\sqrt{n})$
by the continuous mapping theorem, since the estimated nuisance functions and estimated bounds are constructed on separate folds. We then apply the same argument as Proposition 3.1 to arrive at
$\left\|\widehat{\delta}(\cdot)-\mu_{1}(\cdot)\right\|\leq\|\widehat{\delta}(\cdot)-\widehat{\delta}_{oracle}(\cdot)-\tilde{b}(\cdot)\|+\|\tilde{b}(\cdot)\|+\|\widehat{\delta}_{oracle}(\cdot)-\mu_{1}(\cdot)\|$
for $\tilde{b}(x)=\widehat{\mathbb{E}}_{n}[\hat{b}(X_{i})\mid X_{i}=x]$ the smoothed bias and $\hat{b}(x)=\mathbb{E}[\phi_{\mu,i}(\hat{\eta})-\phi_{\mu,i}(\eta)\mid\mathcal{O}_{1},X_{i}=x]$. Under Assumption A.1, $\|\widehat{\delta}(\cdot)-\widehat{\delta}_{oracle}(\cdot)-\tilde{b}(\cdot)\|=o_{\mathbb{P}}(R_{oracle})$ by Lemma A.1. Furthermore, $\hat{b}(x)^{2}=(a)^{2}$, where
$\hat{b}(x)^{2}=\left\{\frac{\pi_{1}(x)-\hat{\pi}_{1}(x)}{\hat{\pi}_{1}(x)}(\mu_{1}(x)-\hat{\mu}_{1}(x))\right\}^{2}\leq\frac{1}{\epsilon^{2}}\left\{(\pi_{1}(x)-\hat{\pi}_{1}(x))(\mu_{1}(x)-\hat{\mu}_{1}(x))\right\}^{2},$
by iterated expectations and the assumption of bounded propensity score. Putting this together yields
$\|\widehat{\delta}(\cdot)-\mu_{1}(\cdot)\|\leq\|\widehat{\delta}_{oracle}(\cdot)-\mu_{1}(\cdot)\|+\epsilon^{-1}\|\tilde{R}(\cdot)\|+o_{\mathbb{P}}(R_{oracle})$
as desired. $\Box$
Appendix C Additional Theoretical Results
C.1 Identification and Estimation of Predictive Disparities
A central question in algorithmic fairness is whether predictive algorithms perform differently across groups defined by sensitive attributes such as race, gender, or income. As noted in Remark 1 of the main text, differences in predictive performance across groups formalize violations of popular fairness criteria. Consider a binary sensitive attribute $G_{i}\in\{0,1\}$ (e.g., ethnicity, gender, race, etc.), define the overall disparity of $s(\cdot)$ as $\mbox{disp}(s;\beta):=\mbox{perf}_{1}(s;\beta)-\mbox{perf}_{0}(s;\beta)$ for $\mbox{perf}_{g}(s;\beta):=\mathbb{E}[\beta_{0}(X_{i},s)+\beta_{1}(X_{i};s)Y_{i}^{*}\mid G_{i}=g]$ the overall performance on group $G_{i}=g$. The positive class and negative class disparities $\mbox{disp}_{+}(s;\beta)$ and $\mbox{disp}_{-}(s;\beta)$ are defined analogously. In this section, we show that predictive disparities are partially identified under Assumption 2.1. We propose estimators for the bounds on predictive disparities, extending our analysis of overall and positive class performance.
C.1.1 Overall Predictive Disparities
Define $\mathcal{H}(\mbox{disp}(s;\beta))$ to be the set of overall predictive disparities consistent with Assumption 2.1. Further let $\beta_{0,i}^{g}:=\beta_{0,i}/P(G_{i}=g)$, $\beta_{1,i}^{g}:=\beta_{1,i}/P(G_{i}=g)$ for $g\in\{0,1\}$, and $\tilde{\beta}_{0,i}:=G_{i}\beta_{0,i}^{1}-(1-G_{i})\beta_{0,i}^{0}$, $\tilde{\beta}_{1,i}:=G_{i}\beta_{1,i}^{1}-(1-G_{i})\beta_{1,i}^{0}$.
Lemma C.1.
$\mathcal{H}(\mbox{disp}(s;\beta))=[\underline{\mbox{disp}}(s;\beta),\overline{\mbox{disp}}(s;\beta)],$
where
$\overline{\mbox{disp}}(s;\beta):=\mathbb{E}[\tilde{\beta}_{0,i}+\tilde{\beta}_{1,i}\mu_{1}(X_{i})+\tilde{\beta}_{1,i}\pi_{0}(X_{i})(\overline{\nu}_{i}\overline{\delta}_{i}+\underline{\nu}_{i}\underline{\delta}_{i})]$
$\underline{\mbox{disp}}(s;\beta):=\mathbb{E}[\tilde{\beta}_{0,i}+\tilde{\beta}_{1,i}\mu_{1}(X_{i})+\tilde{\beta}_{1,i}\pi_{0}(X_{i})(\overline{\nu}_{i}\underline{\delta}_{i}+\underline{\nu}_{i}\overline{\delta}_{i})].$
for $\overline{\nu}_{i}=G_{i}1\{\beta_{1,i}\geq 0\}+(1-G_{i})1\{\beta_{1,i}\leq 0\}$ and $\underline{\nu}_{i}=G_{i}1\{\beta_{1,i}<0\}+(1-G_{i})1\{\beta_{i,1}>0\}$.
Proof.
For $g\in\{0,1\}$ and $\alpha_{g}=P(G_{i}=g)$, observe that
$\mbox{perf}_{g}(s;\beta)=\alpha_{g}^{-1}\mathbb{E}[\beta_{0,i}1\{G_{i}=g\}+\beta_{1,i}1\{G_{i}=g\}\mu_{1}(X_{i})+\beta_{1,i}1\{G_{i}=g\}\pi_{0}(X_{i})\delta(X_{i})].$
We rewrite $\mbox{disp}(s;\beta)$ as
$\mathbb{E}[\tilde{\beta}_{0,i}+\tilde{\beta}_{1,i}\mu_{1}(X_{i})+\tilde{\beta}_{1,i}\pi_{0}(X_{i})\delta(X_{i})]$
using the definitions of $\tilde{\beta}_{0,i},\tilde{\beta}_{1,i}$. As in the proof of Lemma 2.1, it follows that $\mathcal{H}(\mbox{disp}(s;\beta))$ equals the closed interval
$\tilde{\beta}_{0,i}+\tilde{\beta}_{1,i}\mu_{1}(X_{i})+\tilde{\beta}_{1,i}\pi_{0}(X_{i})\left(1\{\tilde{\beta}_{1,i}\geq 0\}\underline{\delta}_{i}+1\{\tilde{\beta}_{1,i}<0\}\overline{\delta}_{i}\right),$
$\tilde{\beta}_{0,i}+\tilde{\beta}_{1,i}\mu_{1}(X_{i})+\tilde{\beta}_{1,i}\pi_{0}(X_{i})\left(1\{\tilde{\beta}_{1,i}\geq 0\}\overline{\delta}_{i}+1\{\tilde{\beta}_{1,i}<0\}\underline{\delta}_{i}\right)].$
The result then follows by noticing that
$1\{\tilde{\beta}_{1,i}\geq 0\}=1\{(G_{i}-\alpha_{1})\beta_{i,1}\geq 0\}=G_{i}1\{\beta_{1,i}\geq 0\}+(1-G_{i})1\{\beta_{1,i}\leq 0\}.$
$1\{\tilde{\beta}_{1,i}<0\}=1\{(G_{i}-\alpha_{1})\beta_{i,1}<0\}=G_{i}1\{\beta_{1,i}<0\}+(1-G_{i})1\{\beta_{i,1}>0\}.$
Since the bounds on overall disparities are linear functionals of known functions of the data and identified nuisance parameters, we construct estimators as in Section 4.1 of the main text. For simplicity, we develop the estimators assuming $\mathbb{P}(G_{i}=1)$ is known and for observed outcome bounds, but they can be easily extended to the case where this is estimated and for alternative choices of bounding functions.
Estimation Procedure for Observed Outcome Bounds:
We randomly split the data into $K$ disjoint folds. For each fold $k$, we construct estimators of the nuisance functions $\hat{\eta}_{-k}=\left(\hat{\pi}_{1,-k},\hat{\mu}_{1,-k}\right)$ using only the sample of observations not in the $k$-th fold. For each observation in the $k$-th fold, we construct
$\overline{\mbox{disp}}_{i}(\hat{\eta}_{-k}):=\tilde{\beta}_{0,i}+\tilde{\beta}_{1,i}\phi_{\mu,i}(\hat{\eta}_{-k})+\tilde{\beta}_{1,i}(\overline{\nu}_{i}(\overline{\Gamma}-1)+\underline{\nu}_{i}(\underline{\Gamma}-1))\phi_{\pi\mu,i}(\hat{\eta}_{-k}),$ (12)
$\underline{\mbox{disp}}_{i}(\hat{\eta}_{-k}):=\tilde{\beta}_{0,i}+\tilde{\beta}_{1,i}\phi_{\mu,i}(\hat{\eta}_{-k})+\tilde{\beta}_{1,i}(\overline{\nu}_{i}(\underline{\Gamma}-1)+\underline{\nu}_{i}(\overline{\Gamma}-1))\phi_{\pi\mu,i}(\hat{\eta}_{-k}).$ (13)
We then estimate the bounds on overall disparities by $\widehat{\overline{\mbox{disp}}}(s;\beta):=\mathbb{E}_{n}[\overline{\mbox{disp}}_{i}(\hat{\eta}_{-k})]$ and $\widehat{\underline{\mbox{disp}}}(s;\beta):=\mathbb{E}_{n}[\underline{\mbox{disp}}_{i}(\hat{\eta}_{-k})]$. As in Appendix C.2, we estimate the asymptotic covariance matrix as
$n^{-1}\sum_{i=1}^{n}\begin{pmatrix}\hat{\sigma}_{i,11}&\hat{\sigma}_{i,12}\\ \hat{\sigma}_{i,12}&\hat{\sigma}_{i,22}\end{pmatrix}$
for $\hat{\sigma}_{i,11}=(\overline{\mbox{disp}}(O_{i};\hat{\eta}_{-K(i)})-\widehat{\overline{\mbox{disp}}}(s;\beta))^{2}$, $\hat{\sigma}_{i,12}=(\overline{\mbox{disp}}(O_{i};\hat{\eta}_{-K(i)})-\widehat{\overline{\mbox{disp}}}(s;\beta))(\underline{\mbox{disp}}(O_{i};\hat{\eta}_{-K(i)})-\widehat{\underline{\mbox{disp}}}(s;\beta))$, and $\hat{\sigma}_{i,22}=(\underline{\mbox{disp}}(O_{i};\hat{\eta}_{-K(i)})-\widehat{\underline{\mbox{disp}}}(s;\beta))^{2}$. Under regularity conditions and the same arguments as the proof of Proposition 4.1, we can derive the rate of convergence of our proposed estimators and provide conditions under which they are jointly asymptotically normal.
C.1.2 Positive Class Predictive Disparities
We provide non-sharp bounds for the positive class disparity since the positive class disparity is the difference of two linear-fractional functions. Define $\mathcal{H}(\mbox{disp}_{+}(s;\beta))$ to be the set of all positive class disparities that are consistent with Assumption 2.1.
Lemma C.2.
$\mathcal{H}(\mbox{disp}_{+}(s;\beta))\subseteq[\underline{\mbox{disp}}_{+}(s;\beta),\overline{\mbox{disp}}_{+}(s;\beta)],$
where $\overline{\mbox{disp}}_{+}(s;\beta)=\overline{\mbox{perf}}_{+,1}(s;\beta)-\underline{\mbox{perf}}_{+,0}(s,\beta)$, $\underline{\mbox{disp}}_{+}(s;\beta)=\underline{\mbox{perf}}_{+,1}(s;\beta)-\overline{\mbox{perf}}_{+,0}(s,\beta)$ for, $g\in\{0,1\}$,
$\overline{\mbox{perf}}_{+,g}(s;\beta)=\sup_{\delta\in\Delta}\frac{\mathbb{E}[\beta_{0,i}\mu_{1}(X_{i})+\beta_{0,i}\pi_{0}(X_{i})\delta(X_{i})\mid G_{i}=g]}{\mathbb{E}[\mu_{1}(X_{i})+\pi_{0}(X_{i})\delta(X_{i})\mid G_{i}=g]},$
$\underline{\mbox{perf}}_{+,g}(s;\beta)=\inf_{\delta\in\Delta}\frac{\mathbb{E}[\beta_{0,i}\mu_{1}(X_{i})+\beta_{0,i}\pi_{0}(X_{i})\delta(X_{i})\mid G_{i}=g]}{\mathbb{E}[\mu_{1}(X_{i})+\pi_{0}(X_{i})\delta(X_{i})\mid G_{i}=g]}.$
Proof.
Observe that, for $g\in\{0,1\}$ and $\delta(\cdot)\in\Delta$,
$\mbox{perf}_{+,g}(s;\beta,\delta)=\frac{\mathbb{E}[\beta_{0,i}\mu_{1}(X_{i})+\beta_{0,i}\pi_{0}(X_{i})\delta(X_{i})\mid G_{i}=g]}{\mathbb{E}[\mu_{1}(X_{i})+\pi_{0}(X_{i})\delta(X_{i})\mid G_{i}=g]},$
and so the positive class predictive disparity $\mbox{disp}_{+}(s;\beta)$ can be written as $\mbox{disp}_{+}(s;\beta,\delta)=\mbox{perf}_{+,1}(s;\beta,\delta)-\mbox{perf}_{+,0}(s;\beta,\delta)$. The result then follows since
$\sup_{\delta\in\Delta}\mbox{disp}_{+}(s;\beta,\delta)\leq\sup_{\delta\in\Delta}\mbox{perf}_{+,1}(s;\beta,\delta)-\inf_{\delta\in\Delta}\mbox{perf}_{+,0}(s;\beta,\delta)$
and
$\inf_{\delta\in\Delta}\mbox{disp}_{+}(s;\beta,\delta)\geq\inf_{\delta\in\Delta}\mbox{perf}_{+,1}(s;\beta,\delta)-\sup_{\delta\in\Delta}\mbox{perf}_{+,0}(s;\beta,\delta).$
We can therefore construct estimators for the non-sharp bounds on positive class disparities by directly applying our estimator in Section 4.2 of the main text. We separately estimate the bounds on group-specific positive class performance $\overline{\mbox{perf}}_{+,g}(s;\beta)$ and $\underline{\mbox{perf}}_{+,g}(s;\beta)$ and then take the appropriate difference. The same arguments as the proof of Proposition 4.2 can be applied to show that this estimator inherits that a partial double robustness property.
C.2 Variance Estimation for Overall Performance Bounds
As mentioned in Section 4.1 of the main text, we develop a consistent estimator of the asymptotic covariance matrix of the estimators for the bounds on overall performance under observed outcome bounds. Recall from Proposition 4.1, if $R_{1,n}^{k}=o_{\mathbb{P}}(1/\sqrt{n})$ for all folds $k$, then
$\sqrt{n}\left(\begin{pmatrix}\widehat{\overline{\mbox{perf}}}(s;\beta)\\ \widehat{\underline{\mbox{perf}}}(s;\beta)\end{pmatrix}-\begin{pmatrix}\overline{\mbox{perf}}(s;\beta)\\ \underline{\mbox{perf}}(s;\beta)\end{pmatrix}\right)\xrightarrow{d}N\left(0,\Sigma\right),$
where $\Sigma=Cov\left((\overline{\mbox{perf}}_{i},\underline{\mbox{perf}}_{i})^{\prime}\right)$ for
$\overline{\mbox{perf}}_{i}(\eta)=\beta_{0,i}+\beta_{1,i}\phi_{\mu,i}(\eta)+\beta_{1,i}\left(1\{\beta_{1,i}>0\}(\overline{\Gamma}-1)+1\{\beta_{1,i}\leq 0\}(\underline{\Gamma}-1)\right)\phi_{\pi\mu,i}(\eta)$
$\underline{\mbox{perf}}_{i}(\eta)=\beta_{0,i}+\beta_{1,i}\phi_{\mu,i}(\eta)+\beta_{1,i}\left(1\{\beta_{1,i}>0\}(\underline{\Gamma}-1)+1\{\beta_{1,i}\leq 0\}(\overline{\Gamma}-1)\right)\phi_{\pi\mu,i}(\eta)$
and $\mathbb{E}[\overline{\mbox{perf}}_{i}]=\overline{\mbox{perf}}(s;\beta)$, $\mathbb{E}[\underline{\mbox{perf}}_{i}]=\underline{\mbox{perf}}(s;\beta)$. Consider the estimator
$\widehat{\Sigma}=\frac{1}{n}\sum_{i=1}^{n}\begin{pmatrix}\widehat{\overline{\mbox{perf}}}_{i}(\hat{\eta}_{-K_{i}})-\widehat{\overline{\mbox{perf}}}(s;\beta)\\ \widehat{\underline{\mbox{perf}}}_{i}(\hat{\eta}_{-K_{i}})-\widehat{\underline{\mbox{perf}}}(s;\beta)\end{pmatrix}\begin{pmatrix}\widehat{\overline{\mbox{perf}}}_{i}(\hat{\eta}_{-K_{i}})-\widehat{\overline{\mbox{perf}}}(s;\beta)\\ \widehat{\underline{\mbox{perf}}}_{i}(\hat{\eta}_{-K_{i}})-\widehat{\underline{\mbox{perf}}}(s;\beta)\end{pmatrix}^{\prime}$.
To show $\widehat{\Sigma}\xrightarrow{p}\Sigma$, it suffices to show convergence in probability for each entry. We prove this directly by extending Lemma 1 in DornGuoKallus(21).
Lemma C.3.
Let $\phi_{1},\phi_{2}$ be any two square integrable functions. Let $\hat{\phi}_{1,n}=\left(\hat{\phi}_{1}(O_{i}),\ldots,\hat{\phi}_{1}(O_{n})\right),\hat{\phi}_{2,n}=\left(\hat{\phi}_{2}(O_{i}),\ldots,\hat{\phi}_{2}(O_{n})\right)$ be random vectors satisfying
$\|\hat{\phi}_{1,n}-\phi_{1,n}\|_{L_{2}(\mathbb{P}_{n})}:=\sqrt{n^{-1}\sum_{i=1}^{n}(\hat{\phi}_{1}(O_{i})-\phi_{1}(O_{i}))^{2}}=o_{\mathbb{P}}(1),$
$\|\hat{\phi}_{2,n}-\phi_{2,n}\|_{L_{2}(\mathbb{P}_{n})}:=\sqrt{n^{-1}\sum_{i=1}^{n}(\hat{\phi}_{2}(O_{i})-\phi_{2}(O_{i}))^{2}}=o_{\mathbb{P}}(1),$
where $\phi_{1,n}=\left(\phi_{1}(O_{i}),\ldots,\phi_{1}(O_{n})\right)$ and $\phi_{2,n}=\left(\phi_{2}(O_{i}),\ldots,\phi_{2}(O_{n})\right)$. Define $\mathbb{P}_{n}$ to be the empirical distribution. Then, the second moments of $\mathbb{P}_{n}$ converge in probability to the respective second moments of $\left(\phi_{1}(O_{i}),\phi_{2}(O_{i})\right)\sim P$
Proof.
Let $\hat{\phi}_{i,1}=\hat{\phi}_{i}(O_{i})$ and define $\phi_{i,1},\hat{\phi}_{i,1},\phi_{i,2}$ analogously. To prove this result, we first show that $\mathbb{E}_{n}[\hat{\phi}_{1,i}^{2}]=\mathbb{E}[\phi_{1,i}^{2}]+o_{\mathbb{P}}(1)$ since the same argument applies for $\phi_{2,i}$. Observe that
$n^{-1}\sum_{i=1}^{n}\hat{\phi}_{1,i}-\mathbb{E}[\phi_{i,1}^{2}]=n^{-1}\sum_{i=1}^{n}(\hat{\phi}_{i,1}^{2}-\phi_{i,1}^{2})+(\mathbb{E}_{n}-\mathbb{E})[\phi_{i,1}^{2}],$
where $(\mathbb{E}_{n}-\mathbb{E})[\phi_{i,1}^{2}]=o_{\mathbb{P}}(1)$. Furthermore, we can rewrite the first term as
$n^{-1}\sum_{i=1}^{n}(\hat{\phi}_{i,1}^{2}-\phi_{i,1}^{2})=n^{-1}\sum_{i=1}^{n}\left(\hat{\phi}_{i,1}-\phi_{i,1}\right)\left(\hat{\phi}_{i,1}+\phi_{i,1}\right)=$
$n^{-1}\sum_{i=1}^{n}\left(\hat{\phi}_{i,1}-\phi_{i,1}\right)\left(\hat{\phi}_{i,1}-\phi_{i,1}+2\phi_{i,1}\right)\leq\|\hat{\phi}_{i,n}-\phi_{i,1}\|\left(\|\hat{\phi}_{i,1}-\phi_{i,1}\|+2\|\phi_{i,1}\|\right)=o_{\mathbb{P}}(1),$
where the last inequality applies the Cauchy-Schwarz inequality and triangle inequality. We next show that $\mathbb{E}_{n}[\hat{\phi}_{i,1}\hat{\phi}_{i,2}]=\mathbb{E}[\phi_{i,1}\phi_{i,2}]+o_{\mathbb{P}}(1)$. Observe that
$n^{-1}\sum_{i=1}^{n}\hat{\phi}_{i,1}\hat{\phi}_{i,2}-\mathbb{E}[\phi_{i,1}\phi_{i,2}]=n^{-1}\sum_{i=1}^{n}\left(\hat{\phi}_{i,1}\hat{\phi}_{i,2}-\phi_{i,1}\phi_{i,2}\right)+(\mathbb{E}_{n}-\mathbb{E})[\phi_{i,1}\phi_{i,2}],$
where $(\mathbb{E}_{n}-\mathbb{E})[\phi_{i,1}\phi_{i,2}]=o_{\mathbb{P}}(1)$. We can further rewrite the first term as
$n^{-1}\sum_{i=1}^{n}\left(\hat{\phi}_{i,1}\hat{\phi}_{i,2}-\phi_{i,1}\phi_{i,2}\right)=n^{-1}\sum_{i=1}^{n}\left(\hat{\phi}_{i,1}(\hat{\phi}_{i,2}-\phi_{i,2})+\phi_{i,2}(\hat{\phi}_{i,1}-\phi_{i,1})\right)=$
$n^{-1}\sum_{i=1}^{n}\phi_{i,1}(\hat{\phi}_{i,2}-\phi_{i,2})+n^{-1}\sum_{i=1}^{n}(\hat{\phi}_{i,1}-\phi_{i,1})(\hat{\phi}_{i,2}-\phi_{i,2})+n^{-1}\sum_{i=1}^{n}\phi_{i,2}(\hat{\phi}_{i,1}-\phi_{i,1})\leq$
$\|\phi_{1,n}\|\|\hat{\phi}_{2,n}-\phi_{2,n}\|+\|\hat{\phi}_{1,n}-\phi_{1,n}\|\|\hat{\phi}_{2,n}-\phi_{2,n}\|+\|\phi_{2,n}\|\|\hat{\phi}_{1,n}-\phi_{1,n}\|=o_{\mathbb{P}}(1),$
where the last inequality applies Cauchy-Schwarz inequality. ∎
We show that the conditions of Lemma C.3 are satisfied for $\widehat{\overline{\mbox{perf}}}_{i}(\hat{\eta}_{-K_{i}})$ and $\widehat{\underline{\mbox{perf}}}_{i}(\hat{\eta}_{-K_{i}})$. The convergence of $\hat{\Sigma}$ follows by the continuous mapping theorem since we already established the convergence of the first moments in Proposition 4.1.
Lemma C.4.
Under the same assumptions as Proposition 4.1, for each fold $k$,
$\|\widehat{\overline{\mbox{perf}}}_{i}(\hat{\eta}_{-k})-\overline{\mbox{perf}}_{i}(\eta)\|_{L_{2}(\mathbb{P}_{n}^{k})}=o_{\mathbb{P}}(1)$
$\|\widehat{\underline{\mbox{perf}}}_{i}(\hat{\eta}_{-k})-\underline{\mbox{perf}}_{i}(\eta)\|_{L_{2}(\mathbb{P}_{n}^{k})}=o_{\mathbb{P}}(1)$
conditionally on $\mathcal{O}_{-k}$.
Proof.
We prove the result for $\widehat{\overline{\mbox{perf}}}_{i}(\hat{\eta}_{-k})$ since the analogous argument applies for $\widehat{\underline{\mbox{perf}}}_{i}(\hat{\eta}_{-k})$. Following the proof of Lemma B.4 and Lemma B.5, we observe that
$\|\widehat{\overline{\mbox{perf}}}_{i}(\hat{\eta}_{-k})-\overline{\mbox{perf}}_{i}(\eta)\|_{L_{2}(\mathbb{P}_{n}^{k})}\leq(a)+(b)$
for
$(a)=M(\frac{\|D_{i}\|_{L_{2}(\mathbb{P}_{n}^{k})}}{\delta}\frac{\|\pi_{1}-\hat{\pi}_{1}\|_{L_{2}(\mathbb{P}_{n}^{k})}}{\epsilon}\|Y_{i}-\mu_{1}(X_{i})\|_{L_{2}(\mathbb{P}_{n}^{k})}+$
$\frac{\|D_{i}\|_{L_{2}(\mathbb{P}_{n}^{k})}}{\delta}\frac{\|\pi_{1}-\hat{\pi}_{1}\|_{L_{2}(\mathbb{P}_{n}^{k})}}{\epsilon}\|\hat{\mu}_{1}-\mu_{1}\|_{L_{2}(\mathbb{P}_{n}^{k})}+\frac{\|D_{i}\|_{L_{2}(\mathbb{P}_{n}^{k})}}{\delta}\|\hat{\mu}_{1}-\mu_{1}\|_{L_{2}(\mathbb{P}_{n}^{k})}+\left\|\hat{\mu}_{1}-\mu_{1}\right\|_{L_{2}(\mathbb{P}_{n}^{k})})$
and
$(b)=\tilde{M}(\frac{1}{\epsilon\delta}\|\hat{\pi}_{1}-\pi_{1}\|_{L_{2}(\mathbb{P}_{n}^{k})}\|Y_{i}-\mu_{1}\|_{L_{2}(\mathbb{P}_{n}^{k})}+\frac{1}{\epsilon\delta}\|\hat{\pi}_{1}-\pi_{1}\|_{L_{2}(\mathbb{P}_{n}^{k})}\|\mu_{1}-\hat{\mu}_{1}\|_{L_{2}(\mathbb{P}_{n}^{k})}+\frac{1-\delta}{\delta}\|\mu_{1}-\hat{\mu}_{1}\|_{L_{2}(\mathbb{P}_{n}^{k})}+$
$\|(1-D_{i})-\pi_{0}\|_{L_{2}(\mathbb{P}_{n}^{k})}\|\hat{\mu}_{1}-\mu_{1}\|_{L_{2}(\mathbb{P}_{n}^{k})}+\|\hat{\pi}_{1}-\pi_{1}\|_{L_{2}(\mathbb{P}_{n}^{k})}\|\hat{\mu}_{1}-\mu_{1}\|_{L_{2}(\mathbb{P}_{n}^{k})}+\|\hat{\pi}_{1}-\pi_{1}\|_{L_{2}(\mathbb{P}_{n}^{k})}\|\mu_{1}\|_{L_{2}(\mathbb{P}_{n}^{k})}+$
$(1-\epsilon)\|\hat{\mu}_{1}-\mu_{1}\|_{L_{2}(\mathbb{P}_{n}^{k})}+\|\mu_{1}\|_{L_{2}(\mathbb{P}_{n}^{k})}\|\hat{\pi}_{1}-\pi_{1}\|_{L_{2}(\mathbb{P}_{n}^{k})})$
and constants $M,\tilde{M}$. However, by Markov’s Inequality,
$(a)=\mathcal{O}_{\mathbb{P}}(\|\pi_{1}-\hat{\pi}_{1}\|_{L_{2}(\mathbb{P})}+\|\pi_{1}-\hat{\pi}_{1}\|_{L_{2}(\mathbb{P})}\|\hat{\mu}_{1}-\mu_{1}\|_{L_{2}(\mathbb{P})}+\|\hat{\mu}_{1}-\mu_{1}\|_{L_{2}(\mathbb{P})})$
and
$(b)=\mathcal{O}_{\mathbb{P}}(\|\pi_{1}-\hat{\pi}_{1}\|_{L_{2}(\mathbb{P})}+\|\pi_{1}-\hat{\pi}_{1}\|_{L_{2}(\mathbb{P})}\|\hat{\mu}_{1}-\mu_{1}\|_{L_{2}(\mathbb{P})}+\|\hat{\mu}_{1}-\mu_{1}\|_{L_{2}(\mathbb{P})}).$
The result is then immediate. ∎
In our Monte Carlo simulations and empirical application, we use this formula for calculating standard errors. In the proof of Proposition 4.1, we show that our estimator is asymptotically equivalent to the centered average of the true bounds, so bootstrap-based inference would also be valid.
C.3 Linear Programming Reduction for Bounds on Positive Class Performance
As mentioned in Section 4.2 of the main text, the optimization program defining $\widehat{\overline{\mbox{perf}}}_{+}(s;\beta)$ can be characterized as a linear program by applying the Charnes-Cooper transformation (CharnesCooper(62)). Lemma C.5 states this reduction for arbitrary bounding functions $\underline{\delta}_{i}:=\underline{\delta}(X_{i};\eta)$, $\overline{\delta}_{i}:=\overline{\delta}(X_{i};\eta)$ and for any fold of observations.
Lemma C.5.
Let $n^{k}$ denote the number of observations in any fold $k$ and $\mathbb{E}_{n}^{k}[\cdot]$ the sample average over all observations in the $k$-th fold. Define $\widehat{c}^{k}=\mathbb{E}_{n}^{k}[\beta_{0,i}\phi_{\mu,i}(\hat{\eta})+\beta_{0,i}(1-D_{i})\widehat{\underline{\delta}}(X_{i})]$, $\hat{d}^{k}=\mathbb{E}_{n}^{k}[(\phi_{\mu,i}(\hat{\eta})+(1-D_{i})\widehat{\underline{\delta}}(X_{i})]$, $\widehat{\alpha}_{i}=n_{k}^{-1}\beta_{0,i}(1-D_{i})(\widehat{\overline{\delta}}(X_{i})-\widehat{\underline{\delta}}(X_{i}))$, $\widehat{\gamma}_{i}=n_{k}^{-1}(1-D_{i})(\widehat{\overline{\delta}}(X_{i})-\widehat{\underline{\delta}}(X_{i}))$, and $\widehat{\alpha}=(\hat{\alpha}_{1},\ldots,\hat{\alpha}_{n})$, $\widehat{\gamma}=(\hat{\gamma}_{1},\ldots,\hat{\gamma}_{n})$. Then,
$\widehat{\overline{\mbox{perf}}}_{+}^{k}(s;\beta)=\max_{\tilde{U}\in\mathbb{R}^{n^{k}},\tilde{V}\in\mathbb{R}}$ $\hat{\alpha}^{\prime}\tilde{U}+\hat{c}^{k}\tilde{V}$
s.t. $0\leq\tilde{U}_{i}\leq\tilde{V}\mbox{ for }i=1,\ldots n_{k},$
$0\leq\tilde{V},\mbox{ }\hat{\gamma}^{\prime}\tilde{U}+\tilde{V}\hat{d}^{k}=1.$
$\widehat{\underline{\mbox{perf}}}_{+}^{k}(s;\beta,\Delta_{n})$ is optimal value of the corresponding minimization problem.
Proof.
We first use the change-of-variables $\delta(X_{i})=\widehat{\underline{\delta}}(X_{i})+(\widehat{\overline{\delta}}(X_{i})-\widehat{\underline{\delta}}(X_{i}))U_{i}$ for $U_{i}\in[0,1]$ to rewrite $\widehat{\overline{\mbox{perf}}}_{+}^{k}(s;\beta)$ as
$\widehat{\overline{\mbox{perf}}}_{+}^{k}(s;\beta):=\max_{U}$ $\frac{\mathbb{E}_{n}^{k}[\beta_{0,i}\phi_{\mu,i}(\hat{\eta})+\beta_{0,i}(1-D_{i})\widehat{\underline{\delta}}(X_{i})+\beta_{0,i}(1-D_{i})(\widehat{\overline{\delta}}(X_{i})-\widehat{\underline{\delta}}(X_{i}))U_{i}]}{\mathbb{E}_{n}^{k}[\phi_{\mu,i}(\hat{\eta})+(1-D_{i})\widehat{\underline{\delta}}(X_{i})+(1-D_{i})(\widehat{\overline{\delta}}(X_{i})-\widehat{\underline{\delta}}(X_{i}))U_{i}]}$
s.t. $0\leq U_{i}\leq 1\mbox{ for }i=1,\ldots,n_{k},$
where $U=\left(U_{1},\ldots,U_{n}\right)^{\prime}$.
Next, define $\hat{c}^{k}=\mathbb{E}_{n}^{k}[\beta_{0,i}\phi_{\mu,i}(\hat{\eta})+\beta_{0,i}(1-D_{i})\widehat{\underline{\delta}}(X_{i})]$, $\hat{d}=\mathbb{E}_{n}^{k}[\phi_{\mu,i}(\hat{\eta})+(1-D_{i})\widehat{\underline{\delta}}(X_{i})]$, $\hat{\alpha}_{i}:=n_{k}^{-1}\beta_{0,i}(1-D_{i})(\overline{\delta}(X_{i})-\widehat{\underline{\delta}}(X_{i}))$, $\hat{\gamma}_{i}:=n_{k}^{-1}(1-D_{i})(\widehat{\overline{\delta}}(X_{i})-\widehat{\underline{\delta}}(X_{i}))$. We can further rewrite the estimator as
$\widehat{\overline{\mbox{perf}}}_{+}^{k}(s;\beta)=\max_{U}\frac{\hat{\alpha}^{\prime}U+\hat{c}^{k}}{\hat{\gamma}^{\prime}U+\hat{d}^{k}}\mbox{ s.t. }0\leq U_{i}\leq 1\mbox{ for }i=1,\ldots,n_{k},$
where $\hat{\alpha}=\left(\hat{\alpha}_{1},\ldots,\hat{\alpha}_{n}\right)^{\prime}$, $\hat{\gamma}=\left(\hat{\gamma}_{1},\ldots,\hat{\gamma}_{n}\right)^{\prime}$. Finally, applying the Charnes-Cooper transformation with $\tilde{U}=\frac{U}{\hat{\gamma}^{\prime}U+\hat{d}^{k}}$, $\tilde{V}=\frac{1}{\hat{\gamma}^{\prime}U+\hat{d}^{k}}$, this linear-fractional program is equivalent to the linear program
$\widehat{\overline{\mbox{perf}}}_{+}^{k}(s;\beta)=\max_{\tilde{U},\tilde{V}}\,$ $\hat{\alpha}^{\prime}\tilde{U}+\hat{c}^{k}\tilde{V}$
s.t. $0\leq\tilde{U}_{i}\leq\tilde{V}\mbox{ for }i=1,\ldots n_{k},$
$0\leq\tilde{V},\mbox{ }\hat{\gamma}^{\prime}\tilde{U}+\tilde{V}\hat{d}^{k}=1.$
Appendix D Estimation under Alternative Bounding Functions
In the main text, we developed estimation procedures for observed outcome bounds (Section 2.3.1). In this appendix, we extend our estimation framework to two additional classes of bounding functions: proxy outcome bounds (Section 2.3.2) and instrumental variable bounds (Section 2.3.3). Both approaches leverage additional data — proxy outcomes or quasi-experimental variation — to sharpen inference about unobserved confounding. While this complicates notation, we nonetheless can extend our estimators for the bounds on the conditional likelihood and predictive performance measures.
D.1 Estimation under Proxy Outcome Bounds
Suppose the proxy outcome $\tilde{Y}_{i}\in\{0,1\}$ is always observed (for both selected and unselected units) and statistically related to the true outcome $Y_{i}^{*}$. We assume that the proxy outcome is equally informative about the true outcome for selected and unselected units: that is, for all for all $x\in\mathcal{X}$,
$\mathbb{P}(Y_{i}^{*}=\tilde{Y}_{i}\mid D_{i}=0,X_{i}=x)=\gamma_{1}(x).$ (14)
Equation 14 states that the probability the true and proxy outcomes agree is the same regardless of selection status, conditional on covariates. In consumer lending, this says that default on credit cards is equally predictive of default on personal loans for both funded and unfunded applicants with the same credit profile.
Our next result characterizes the sharp bounds on the confounding function $\delta(x)$ under Equation 14.
Lemma D.1.
Under Equation (14), the confounding function satisfies, for all $x\in\mathcal{X}$,
$\underline{\delta}(x;\eta)\leq\delta(x)\leq\overline{\delta}(x;\eta)$
for $\underline{\delta}(x;\eta)=|1-\gamma_{1}(x)-\tilde{\mu}_{0}(x)|-\mu_{1}(x)$ and $\overline{\delta}(x;\eta)=1-|\gamma_{1}(x)-\tilde{\mu}_{0}(x)|-\mu_{1}(x)$.
Proof.
We consider a particular $x\in\mathcal{X}$ and drop the dependence on $x$ in our notation. Let $p(y^{*},\tilde{y})=\mathbb{P}(Y_{i}^{*}=y^{*},\tilde{Y}_{i}=\tilde{y}\mid D_{i}=0)$ for all $(y^{*},\tilde{y})\in\{0,1\}^{2}$. Observe that $\mathbb{P}(Y_{i}^{*}=1\mid D_{i}=0)=p(1,1)+p(1,0)$ and $\mathbb{P}(\tilde{Y}_{i}^{*}=1\mid D_{i}=0)=p(1,1)+p(0,1)$. The upper bound on $\mathbb{P}(Y_{i}^{*}=1\mid D_{i}=0)$ can be expressed as the optimal value of the following linear program
$\max_{p(0,0),p(0,1),p(1,0),p(1,1)}$ $\,p(1,1)+p(1,0)$
s.t. $p(1,1)+p(0,0)=\gamma_{1},$
$p(1,1)+p(0,1)=\tilde{\mu}_{0},$
$p(0,0)+p(0,1)+p(1,0)+p(1,1)=1,$
$p(0,0),p(0,1),p(1,0),p(1,1)\geq 0.$
The constraints $p(1,1)+p(0,1)=\tilde{\mu}_{0}$ and $p(1,1)+p(0,0)=\gamma_{1}$ imply $p(0,1)=\tilde{\mu}_{0}-p(1,1)$ and $p(0,0)=\gamma_{1}-p(1,1)$ respectively. The non-negativity constraints then imply $0\leq p(1,1)\leq\min\{\gamma_{1},\tilde{\mu}_{0}\}$. Substituting in, the constraint $p(0,0)+p(0,1)+p(1,0)+p(1,1)=1$ then implies
$p(1,0)=1-\gamma_{1}-\tilde{\mu}_{0}+p(1,1),$
and the non-negativity constraint implies $\max\{0,\gamma_{1}+\tilde{\mu}_{0}-1\}\leq p(1,1)$. We can therefore rewrite the linear program as
$\max_{p(1,1)}$  $\,2p(1,1)+(1-\gamma_{1}-\tilde{\mu}_{0})$
s.t. $\max\{0,\gamma_{1}+\tilde{\mu}_{0}-1\}\leq p(1,1)\leq\min\{\gamma_{1},\tilde{\mu}_{0}\}.$
The upper bound on $\mathbb{P}(Y_{i}^{*}=1\mid D_{i}=0)$ is therefore
$2\min\{\gamma_{1},\tilde{\mu}_{0}\}+(1-\gamma_{1}-\tilde{\mu}_{0})=\min\{1+\gamma_{1}-\tilde{\mu}_{0},1-\gamma_{1}+\tilde{\mu}_{1}\}=1-|\gamma_{1}-\tilde{\mu}_{0}|.$
Analogously, the lower bound is
$2\max\{0,\gamma_{1}+\tilde{\mu}_{1}-1\}+(1-\gamma_{1}-\tilde{\mu}_{0})=\max\{1-\gamma_{1}-\tilde{\mu}_{0},\tilde{\mu}_{0}+\gamma_{1}-1\}=|1-\gamma_{1}-\tilde{\mu}_{0}|.$
The result is then immediate. ∎
In many applications, the proxy satisfies additional restrictions that simplify the bounding functions. The bounding functions stated in Equation 6 of the main text immediately follow from Lemma D.1 if further $\tilde{\mu}_{0}(x)\leq\gamma_{1}(x)$ and $\tilde{\mu}_{0}(x)+\gamma_{1}(x)\leq 1$ for all $x\in\mathcal{X}$. These restrictions are satisfied by the proxy outcomes used in BlattnerNelson(21) (see their Table 3) and MullainathanObermeyer(21) (see their Table 2) on average over the covariates.
For the remainder of this section, we extend our estimators for the conditional probability, positive class performance, and overall performance to the simplified bounding functions based on a proxy outcome $\tilde{Y}_{i}$ satisfying $\tilde{\mu}_{0}(x)\leq\gamma_{1}(x)$ and $\tilde{\mu}_{0}(x)+\gamma_{1}(x)\leq 1$ for all $x\in\mathcal{X}$. The simplified bounding functions are:
$\underline{\delta}(x;\eta)=1-\gamma_{1}(x)-\tilde{\mu}_{0}(x)-\mu_{1}(x),$
$\overline{\delta}(x;\eta)=1-\gamma_{1}(x)+\tilde{\mu}_{0}(x)-\mu_{1}(x).$
We observe that
$\phi_{\tilde{\mu},i}(\eta)=\tilde{\mu}_{0}(X_{i})+\frac{1-D_{i}}{1-\pi_{1}(X_{i})}(\tilde{Y}_{i}-\tilde{\mu}_{0}(X_{i})),$
$\phi_{\gamma,i}(\eta)=\gamma_{1}(X_{i})-\frac{D_{i}}{\pi_{1}(X_{i})}(Y_{i}\tilde{Y}_{i}+(1-Y_{i})(1-\tilde{Y}_{i})-\gamma_{1}(X_{i}))$
$\phi_{\pi\tilde{\mu},i}(\eta)=\pi_{0}(X_{i})\tilde{\mu}_{0}(X_{i})+((1-D_{i})-\pi_{0}(X_{i}))\tilde{\mu}_{0}(X_{i})+\frac{1-D_{i}}{1-\pi_{1}(X_{i})}(\tilde{Y}_{i}-\tilde{\mu}_{0}(X_{i}))\pi_{0}(X_{i}),$
$\phi_{\pi\gamma,i}(\eta)=\pi_{0}(X_{i})\gamma_{1}(X_{i})+((1-D_{i})-\pi_{0}(X_{i}))\gamma_{1}(X_{i})+\frac{D_{i}}{\pi_{1}(X_{i})}(Y_{i}\tilde{Y}_{i}+(1-Y_{i})(1-\tilde{Y}_{i})-\gamma_{1}(X_{i}))\pi_{0}(X_{i})$
are influence functions for $\mathbb{E}[\tilde{\mu}_{0}(X_{i})],\mathbb{E}[\gamma_{1}(X_{i})]$, $\mathbb{E}[\pi_{i}(X_{i})\tilde{\mu}_{0}(X_{i})]$, and $\mathbb{E}[\pi_{0}(X_{i})\gamma_{1}(X_{i})]$ respectively by standard calculations (e.g., Kennedy(22)-IFreview; Hines_2022).
To extend our estimators to the general proxy outcome bounds in Lemma D.1, we use smooth approximations to the absolute value functions. The lower bounding function can also be written as $\max\{1-\gamma_{1}(x)-\tilde{\mu}_{0}(x),\tilde{\mu}_{0}(x)+\gamma_{1}(x)-1\}-\mu_{1}(x)$ and the upper bounding function as $\min\{1-\gamma_{1}(x)+\tilde{\mu}_{0}(x),1+\gamma_{1}(x)-\tilde{\mu}_{0}(x)\}$. We apply the log-sum-exponential function to approximate these pointwise maximum and minimum operations as we discuss below in Appendix D.2. Alternatively, for the overall performance estimator, we could invoke a margin condition as discussed in LevisEtAl(23)-CovariateAssistedIV; semenova2023adaptive.
D.1.1 Bounds on the Conditional Likelihood
We extend our nonparametric regression estimator for the proxy outcome bounds on the conditional probability. The construction is identical to Section 3, except we modify the pseudo-outcomes that are constructed in the second-stage. We again illustrate the procedure using two folds for simplicity.
Estimation Procedure for Proxy Outcome Bounds:
[…]
Bound on Integrated Mean Square Error Convergence:
For modified definitions of the smoothed doubly robust residuals and smoothed bias, we again derive a bound on the integrated mean square error of our feasible estimators relative to that of an infeasible oracle nonparametric regression that has access to the true nuisance functions using the same arguments as Proposition 3.1. We state the result for completeness, but skip the proof for brevity.
Proposition D.1.
Let $\widehat{\mathbb{E}}_{n}[\cdot\mid X_{i}=x]$ denote the second-stage pseudo-outcome regression estimator. Suppose $\widehat{\mathbb{E}}_{n}[\cdot\mid X_{i}=x]$ satisfies the $L_{2}(\mathbb{P})$-stability condition (Assumption A.1), and $\mathbb{P}(\epsilon\leq\hat{\pi}_{1}(X_{i})\leq 1-\epsilon)=1$ for some $\epsilon>0$. Define $\tilde{R}_{1}(x)=\widehat{\mathbb{E}}_{n}[(\pi_{1}(X_{i})-\hat{\pi}_{1}(X_{i}))(\mu_{1}(X_{i})-\hat{\mu}_{1}(X_{i}))\mid X_{i}=x]$, $\tilde{R}_{2}(x)=\widehat{\mathbb{E}}_{n}[(\pi_{1}(X_{i})-\hat{\pi}_{1}(X_{i}))(\hat{\tilde{\mu}}_{0}(X_{i})-\tilde{\mu}_{0}(X_{i}))\mid X_{i}=x]$, $\tilde{R}_{3}(x)=\widehat{\mathbb{E}}_{n}[(\pi_{1}(X_{i})-\hat{\pi}_{1}(X_{i}))(\hat{\gamma}_{1}(X_{i})-\gamma_{1}(X_{i}))\mid X_{i}=x]$, and $R_{oracle}^{2}=\mathbb{E}[\|\widehat{\overline{\mu}}_{oracle}(\cdot)-\overline{\mu}^{*}(\cdot)\|^{2}]$. Then,
$\|\widehat{\overline{\mu}}(\cdot)-\overline{\mu}^{*}(\cdot)\|\leq\|\widehat{\overline{\mu}}_{oracle}(\cdot)-\overline{\mu}^{*}(\cdot)\|+\epsilon^{-1}\left(\|\tilde{R}_{1}(\cdot)\|+\|\tilde{R}_{2}(\cdot)\|+\|\tilde{R}_{3}(\cdot)\|\right)+o_{\mathbb{P}}(R_{oracle})$
The analogous result holds for $\widehat{\underline{\mu}}(x)$.
D.1.2 Bounds on Overall Performance
Under proxy outcome bounds, the upper bound on overall performance can be written as
$\overline{\mbox{perf}}(s;\beta)=\mathbb{E}[\beta_{0,i}+\beta_{1,i}\mu_{1}(X_{i})+\beta_{1,i}(1-D_{i})-\beta_{1,i}(\pi_{0}(X_{i})\gamma_{1}(X_{i})+\pi_{0}(X_{i})\mu_{1}(X_{i})+\pi_{0}(X_{i})\tilde{\mu}_{0}(X_{i}))].$
The lower bound can be written analogously. Like in Section 4.1 of the main text, both bounds are linear functionals of known functions of the data and identified nuisance parameters. This linearity enables us to construct debiased estimators using standard arguments.
Estimation Procedure for Proxy Outcome Bounds:
We sketch the construction of our estimators based on $K$-fold cross-fitting. We randomly split the data into $K$ disjoint folds. For each fold $k$, we estimate the nuisance functions $\hat{\eta}_{-k}$ using all observations not in the $k$-th fold and construct
$\overline{\mbox{perf}}_{i}(\hat{\eta}_{-k})=\beta_{0,i}+\beta_{1,i}\phi_{\mu,i}(\hat{\eta}_{-k})+\beta_{1,i}(1-D_{i})-\beta_{1,i}(\phi_{\pi\gamma,i}(\hat{\eta}_{-k})+\phi_{\pi\mu,i}(\hat{\eta}_{-k})+\phi_{\pi\tilde{\mu},i}(\hat{\eta}_{-k}))$
for each observation $i$ in the $k$-th fold. We take the average across all observations and return $\widehat{\overline{\mbox{perf}}}(s;\beta)=\mathbb{E}_{n}[\widehat{\overline{\mbox{perf}}}(s;\beta)(\hat{\eta}_{-K_{i}})]$. The estimator for the lower bound is defined analogously.
Consistency and Asymptotic Normality:
Using the same arguments as in the proof of Proposition 4.1, we can again show that these estimators converge to the proxy outcome bounds on overall performance and are jointly asymptotically normal. We next state the result, but skip the proof for brevity.
Proposition D.2.
Define $R_{1,n}^{k}=\|\hat{\mu}_{1,-k}(\cdot)-\mu_{1}(\cdot)\|\|\hat{\pi}_{1,-k}(\cdot)-\pi_{1}(\cdot)\|$, $R_{2,n}^{k}=\|\tilde{\hat{\mu}}_{0,-k}(\cdot)-\tilde{\mu}_{0}(\cdot)\|\|\hat{\pi}_{1,-k}(\cdot)-\pi_{1}(\cdot)\|$, and $R_{3,n}^{k}=\|\hat{\gamma}_{1,-k}(\cdot)-\gamma_{1}(\cdot)\|\|\hat{\pi}_{1,-k}(\cdot)-\pi_{1}(\cdot)\|$ for each fold $k=1,\ldots,K$. Assume (i) there exists some $M<\infty$ such that $\|\beta_{1}(\cdot)\|\leq M$; (ii) $\mathbb{P}(\pi_{1}(X_{i})\geq\delta)=1$ for some $\delta>0$, (iii) there exists $\epsilon>0$ such that $\mathbb{P}(\hat{\pi}_{1,-k}(X_{i})\geq\epsilon)=1$ for each fold $k=1,\ldots,K$, and (iv) $\|\hat{\pi}_{1,-k}(\cdot)-\pi_{1}(\cdot)\|=o_{\mathbb{P}}(1),\|\hat{\mu}_{1,-k}(\cdot)-\mu_{1}(\cdot)\|=o_{\mathbb{P}}(1),\|\tilde{\hat{\mu}}_{0,-k}(\cdot)-\tilde{\mu}_{0}(\cdot)\|=o_{\mathbb{P}}(1)$, and $\|\hat{\gamma}_{1,-k}(\cdot)-\gamma_{1}(\cdot)\|=o_{\mathbb{P}}(1)$ for each fold $k=1,\ldots,K$. Then,
$\left|\widehat{\overline{\mbox{perf}}}(s;\beta)-\overline{\mbox{perf}}(s;\beta)\right|=O_{\mathbb{P}}\left(1/\sqrt{n}+\sum_{k=1}^{K}(R_{1,n}^{k}+R_{2,n}^{k}+R_{3,n}^{k})\right)$
and the analogous result holds for $\widehat{\underline{\mbox{perf}}}(s;\beta)$. If further $R_{1,n}^{k},R_{2,n}^{k},R_{3,n}^{k}=o_{\mathbb{P}}(1/\sqrt{n})$ for all folds $k=1,\ldots,K$, then
$\sqrt{n}\left(\begin{pmatrix}\widehat{\overline{\mbox{perf}}}(s;\beta)\\ \widehat{\underline{\mbox{perf}}}(s;\beta)\end{pmatrix}-\begin{pmatrix}\overline{\mbox{perf}}(s;\beta)\\ \underline{\mbox{perf}}(s;\beta)\end{pmatrix}\right)\xrightarrow{d}N\left(0,\Sigma\right)$
for covariance matrix $\Sigma=Cov\left((\overline{\mbox{perf}}_{i}(\eta),\underline{\mbox{perf}}_{i}(\eta))^{\prime}\right)$.
D.1.3 Bounds on Positive Class Performance
We analogously estimate the proxy outcome bounds on positive class performance by solving sample linear fractional programs as in Section 4.2 of the main text. Building on Proposition 4.2, all that needs to be modified is the choice estimator for the bounding functions. As an illustration, we sketch out how the sample linear fractional program can be combined with our pseudo-regression procedure for proxy outcome bounds discussed earlier in Appendix D.1.1.
Estimation Procedure for Proxy Outcome Bounds:
[…]
$\widehat{\overline{\mbox{perf}}}_{+}(s;\beta)=\max_{\tilde{\delta}\in\widehat{\Delta}_{n}}\,\frac{\mathbb{E}_{n}[\beta_{0,i}\phi_{\mu,i}(\hat{\eta})+\beta_{0,i}\tilde{\delta}_{i}]}{\mathbb{E}_{n}[\phi_{\mu,i}(\hat{\eta})+\tilde{\delta}_{i}]},$
where now $\widehat{\Delta}_{n}=\{(1-D_{i})\widehat{\underline{\delta}}(X_{i})\leq\tilde{\delta}_{i}\leq(1-D_{i})\widehat{\overline{\delta}}(X_{i})\}$. These estimators are again equivalent to solving linear programs by the Charnes-Cooper transformation. Proposition 4.2 can be applied, and this estimator has the same partial double robustness property. The error of this estimator again depends on the root mean square error of the estimated bounding functions. If the nonparametric regression procedure further satisfies the $L_{2}(\mathbb{P})$-stability condition, then we can further control these errors using the results in Appendix D.1.1.
D.2 Estimation under Instrumental Variable Bounds
Suppose we observe an instrumental variable $Z_{i}\in\mathcal{Z}$ with finite support that generates quasi-random variation in selection decisions but does not directly affect outcomes as discussed in Section 2.3.3. The instrument provides identifying power under two assumptions: (i) $Z_{i}$ exogenously shifts the selection decision but not the true outcome, $Y_{i}^{*}\perp\!\!\!\perp Z_{i}\mid X_{i}$; and (ii) $Z_{i}$ is relevant, meaning there exists $z,z^{\prime}\in\mathcal{Z}$ with $\mathbb{P}(D_{i}=1\mid X_{i}=x,Z_{i}=z)\neq\mathbb{P}(D_{i}=1\mid X_{i}=x,Z_{i}=z^{\prime})$. Under these conditions, classic results by Manski(94)-SampleSelectionBounds imply bounds on the confounding function. As notation, let $p=|\mathcal{Z}|$ denote the number of unique instrument values, $\lambda_{z}(x):=\mathbb{E}[Y_{i}D_{i}\mid X_{i}=x,Z_{i}=z]$ and $\kappa_{z}(x):=\mathbb{P}(D_{i}=0\mid X_{i}=x,Z_{i}=z)$.
Recall that for any value of the instrument $z\in\mathcal{Z}$, the confounding function satisfies, for all $x\in\mathcal{X}$,
$\underline{\delta}_{z}(x;\eta)\leq\delta(x)\leq\overline{\delta}_{z}(x;\eta),$
where $\underline{\delta}_{z}(x):=(\lambda_{z}(x)-\mu_{1}(x))/\pi_{0}(x)$ and $\overline{\delta}_{z}(x):=(\kappa_{z}(x)+\lambda_{z}(x)-\mu_{1}(x))/\pi_{0}(x)$. We will first extend our estimators for instrumental variable bounds at a particular instrument value $z\in\mathcal{Z}$. We will then discuss how to extend our estimators to a smooth approximation of the intersection bounds, for all $x\in\mathcal{X}$,
$\max_{z\in\mathcal{Z}}\underline{\delta}_{z}(x;\eta)\leq\delta(x)\leq\min_{z\in\mathcal{Z}}\overline{\delta}_{z}(x;\eta).$
To extend our estimators, it will be convenient to rewrite the instrumental variable bounds as, for each $z\in\mathcal{Z}$,
$\lambda_{z}(x)-\mu_{1}(x)\leq\pi_{0}(x)\delta(x)\leq\kappa_{z}(x)+\lambda_{z}(x)-\mu_{1}(x).$
We also observe that
$\phi_{\lambda,z,i}(\eta):=\frac{1\{Z_{i}=z\}}{\mathbb{P}(Z_{i}=z\mid X_{i})}\left(Y_{i}D_{i}-\lambda_{z}(X_{i})\right)+\lambda_{z}(X_{i}),$
$\phi_{\kappa,z,i}(\eta):=\frac{1\{Z_{i}=z\}}{\mathbb{P}(Z_{i}=z\mid X_{i})}\left((1-D_{i})-\kappa_{z}(X_{i})\right)+\kappa_{z}(X_{i})$
are the influence functions for $\mathbb{E}[\lambda_{z}(X_{i})]$ and $\mathbb{E}[\kappa_{z}(X_{i})]$ respectively.
D.2.1 Bounds on the Conditional Likelihood
We briefly extend our nonparametric regression estimator for for the instrumental variable bounds at a particular instrument value $z\in\mathcal{Z}$. The construction is identical to Section 3, except we modify the pseudo-outcomes that are constructed in the second-stage. We again illustrate the procedure using two folds for simplicity.
Estimation Procedure for Single-Instrument Bounds:
[…]
Bound on Integrated Mean Square Error Convergence:
For modified definitions of the smoothed doubly robust residuals and smoothed bias, we can again derive a bound on the integrated mean square error of our feasible estimators relative to that of an infeasible oracle nonparametric regression that has access to the true nuisance functions using the same arguments as Proposition 3.1. We state the result for completeness, but skip the proof for brevity.
Proposition D.3.
Suppose $\widehat{\mathbb{E}}_{n}[\cdot\mid X_{i}=x]$ satisfies the $L_{2}(\mathbb{P})$-stability condition (Assumption A.1) and $\mathbb{P}(\epsilon\leq\widehat{\mathbb{P}}(Z_{i}=z\mid X_{i}=x))=1$ for some $\epsilon>0$. Define $\tilde{R}_{2}(x)=\widehat{\mathbb{E}}_{n}[(\mathbb{P}(Z_{i}=z\mid X_{i}=x)-\widehat{\mathbb{P}}(Z_{i}=z\mid X_{i}=x))(\hat{\lambda}_{z}(x)-\lambda_{z}(x))\mid X_{i}=x]$, $\tilde{R}_{3}(x)=\widehat{\mathbb{E}}_{n}[(\mathbb{P}(Z_{i}=z\mid X_{i}=x)-\widehat{\mathbb{P}}(Z_{i}=z\mid X_{i}=x))(\hat{\kappa}_{z}(x)-\kappa_{z}(x))\mid X_{i}=x]$, and $R^{2}_{oracle}(z)=\mathbb{E}[\|\widehat{\overline{\mu}}(\cdot)-\overline{\mu}^{*}(\cdot)\|^{2}]$. Then,
$\|\widehat{\overline{\mu}}(\cdot)-\overline{\mu}(\cdot)\|\leq\|\widehat{\overline{\mu}}_{oracle}(\cdot)-\overline{\mu}(\cdot)\|+\epsilon^{-1}\left(\|\tilde{R}_{2}(\cdot)\|+\|\tilde{R}_{3}(\cdot)\|\right)+o_{\mathbb{P}}(R_{oracle}(z)),$
$\|\widehat{\underline{\mu}}(\cdot)-\underline{\mu}(\cdot)\|\leq\|\widehat{\underline{\mu}}_{oracle}(\cdot)-\underline{\mu}(\cdot)\|+\epsilon^{-1}\|\tilde{R}_{2}(\cdot)\|+o_{\mathbb{P}}(R_{oracle}(z)).$
D.2.2 Bounds on Overall Performance
Under instrumental variable bounds at a single instrument value $z\in\mathcal{Z}$, the upper bound on overall performance can be written as
$\overline{\mbox{perf}}(s;\beta)=\mathbb{E}[\beta_{0,i}+\beta_{i,1}\lambda_{z}(X_{i})+\beta_{i,1}1\{\beta_{i,1}>0\}\kappa_{z}(X_{i})].$
Like in Section 4.1, the bounds are again linear functionals of known functions of the data and identified nuisance parameters. We construct debiased estimators using standard arguments.
Estimation Procedure for Single-Instrument Bounds:
We randomly split the data into $K$ disjoint folds. For each fold $k$, we estimate the nuisance functions $\hat{\eta}_{-k}$ using all observations not in the $k$-th fold and construct
$\overline{\mbox{perf}}_{i}(\hat{\eta}_{-k})=\beta_{0,i}+\beta_{i,1}\phi_{\lambda,z,i}(\hat{\eta}_{-k})+\beta_{i,1}1\{\beta_{i,1}>0\}\phi_{\kappa,z,i}(\hat{\eta}_{-k})$
for each observation $i$ in the $k$-th fold. We take the average across all observations and return $\widehat{\overline{\mbox{perf}}}(s;\beta)=\mathbb{E}_{n}[\widehat{\overline{\mbox{perf}}}(s;\beta)(\hat{\eta}_{-K_{i}})]$. The estimator for the lower bound is defined analogously.
Consistency and Asymptotic Normality:
Using the same arguments as in the proof of Proposition 4.1, we can again show that these estimators converge quickly to the proxy outcome bounds on overall performance and are jointly asymptotically normal. We next state the result, but skip the proof for brevity.
Proposition D.4.
Define $R_{2,n}^{k}=\|\hat{\lambda}_{z,-k}(\cdot)-\lambda_{z}(\cdot)\|\|\hat{\mathbb{P}}(Z_{i}=z\mid X_{i}=\cdot)-\mathbb{P}(Z_{i}=z\mid X_{i}=\cdot)\|$ and $R_{3,n}^{k}=\|\hat{\kappa}_{z,-k}(\cdot)-\kappa_{z}(\cdot)\|\|\hat{\mathbb{P}}(Z_{i}=z\mid X_{i}=\cdot)-\mathbb{P}(Z_{i}=z\mid X_{i}=\cdot)\|$. Assume (i) there exists some $M<\infty$ such that $\|\beta_{1}(\cdot)\|\leq M$; (ii) $\mathbb{P}\{\mathbb{P}(Z_{i}=z\mid X_{i})\geq\delta\}=1$; (iii) there exists $\epsilon>0$ such that $\mathbb{P}\{\hat{\mathbb{P}}_{-k}(Z_{i}=z\mid X_{i})\geq\epsilon\}=1$ for all folds $k$; and (iv) $\|\hat{\lambda}_{z}(\cdot)-\lambda_{z}(\cdot)\|=o_{\mathbb{P}}(1)$, $\|\hat{\kappa}_{z}(\cdot)-\kappa_{z}(\cdot)\|=o_{\mathbb{P}}(1)$, and $\|\hat{\mathbb{P}}_{-k}(Z_{i}=z\mid X_{i})-\mathbb{P}(Z_{i}=z\mid X_{i})\|=o_{\mathbb{P}}(1)$ for all folds $k$. Then,
$\left|\widehat{\overline{\mbox{perf}}}(s;\beta)-\overline{\mbox{perf}}(s;\beta)\right|=O_{\mathbb{P}}\left(1/\sqrt{n}+\sum_{k=1}^{K}(R_{2,n}^{k}+R_{3,n}^{k})\right),$
and the analogous result holds for $\widehat{\underline{\mbox{perf}}}(s;\beta)$. If further $R_{2,n}^{k}=o_{\mathbb{P}}(1/\sqrt{n})$ and $R_{3,n}^{k}=o_{\mathbb{P}}(1/\sqrt{n})$ for all folds $k$, then
$\sqrt{n}\left(\begin{pmatrix}\widehat{\overline{\mbox{perf}}}(s;\beta)\\ \widehat{\underline{\mbox{perf}}}(s;\beta)\end{pmatrix}-\begin{pmatrix}\overline{\mbox{perf}}(s;\beta)\\ \underline{\mbox{perf}}(s;\beta)\end{pmatrix}\right)\xrightarrow{d}N\left(0,\Sigma\right)$
for covariance matrix $\Sigma=Cov\left((\overline{\mbox{perf}}_{i},\underline{\mbox{perf}}_{i})^{\prime}\right)$.
D.2.3 Bounds on Positive Class Performance
We analogously estimate the instrumental variable bounds on positive class performance by solving sample linear fractional programs as in Section 4.2 of the main text. All that needs to be modified is the choice estimator for the bounding functions. As an illustration, we sketch out how the sample linear fractional program can be combined with our pseudo-regression procedure discussed earlier in Appendix D.2.1.
Estimation Procedure under Single-Instrument Bounds:
[…]
$\widehat{\overline{\mbox{perf}}}_{+}(s;\beta)=\max_{\tilde{\delta}\in\widehat{\Delta}_{n}}\,\frac{\mathbb{E}_{n}[\beta_{0,i}\phi_{\mu,i}(\hat{\eta})+\beta_{0,i}\tilde{\delta}_{i}]}{\mathbb{E}_{n}[\phi_{\mu,i}(\hat{\eta}_{-k})+\tilde{\delta}_{i}]},$
where now $\widehat{\Delta}_{n}=\{(1-D_{i})\widehat{\underline{\delta}}(X_{i})\leq\tilde{\delta}_{i}\leq(1-D_{i})\widehat{\overline{\delta}}(X_{i})\}$. Proposition 4.2 can be applied, and this estimator has the same partial double robustness property. The error of this estimator again depends on the root mean square error of the estimated bounding functions. If the nonparametric regression procedure further satisfies the $L_{2}(\mathbb{P})$-stability condition, then we can further control these errors using the results in Appendix D.2.1.
D.2.4 Smooth Approximations to Intersection Bounds
We now discuss how our estimators can be extended to the intersection bounds
$\max_{z\in\mathcal{Z}}\underline{\delta}_{z}(x;\eta)\leq\delta(x)\leq\min_{z\in\mathcal{Z}}\overline{\delta}_{z}(x;\eta)$
for all $x\in\mathcal{X}$, where $\underline{\delta}_{z}(x):=(\lambda(x,z)-\mu_{1}(x))/\pi_{0}(x)$ and $\overline{\delta}_{z}(x):=(\pi_{0}(x,z)+\lambda(x,z)-\mu_{1}(x))/\pi_{0}(x)$. Rather than working directly with these intersection bounds, we instead consider bounding functions based on smooth approximations to the pointwise minimum and pointwise maximum functions, following LevisEtAl(23)-CovariateAssistedIV’s analysis of covariate-assisted Balke-Pearl bounds on the average treatment effects. We focus on the log-sum-exponential function but other choices are possible such as the Boltzmann operator. Exploring different choices of smooth approximations is an interesting question for future work.
For a researcher-specified $\alpha>0$, the log-sum-exponential function $g_{\alpha}(\cdot)\colon\mathbb{R}^{p}\rightarrow\mathbb{R}$ is
$g_{\alpha}(v)=\frac{1}{\alpha}\log\left(\sum_{j=1}^{p}\exp(\alpha v_{j})\right).$
The log-sum-exponential function approximates the pointwise maximum function and satisfies the inequality
$\max\,\{v_{1},\ldots,v_{p}\}\leq g_{\alpha}(v)\leq\max\,\{v_{1},\ldots,v_{p}\}+\frac{\log(p)}{\alpha}.$
The choice of parameter $\alpha>0$ therefore determines the quality of the approximation $g_{\alpha}(\cdot)$ provides to the pointwise maximum function. Furthermore, $\nabla g_{\alpha}(v)=\frac{v}{\ell^{T}v}$ for $\ell$ the $p$-dimensional vector of ones and $\nabla^{2}g_{\alpha}(v)=\frac{\alpha}{(\ell^{T}v)^{2}}((\ell^{T}v)diag(v)-vv^{T})$. Analogously, we can approximate the pointwise minimum with the function $g_{-\alpha}(v)$, which satisfies
$\min\,\{v_{1},\ldots,v_{p}\}-\frac{\log(p)}{\alpha}\leq g_{-\alpha}(v)\leq\min\,\{v_{1},\ldots,v_{p}\}.$
We will consider bounding functions that apply the log-sum-exponential function to the $\underline{\delta}_{z}(x;\eta)$ and $\overline{\delta}_{z}(x;\eta)$ respectively. We define the pointwise bounding functions on $\pi_{0}(x)\delta(x)$ as
[…]
$\overline{\delta}(x;\eta)=g_{-\alpha}\left(\kappa_{z_{1}}(x)+\lambda_{z_{1}}(x)-\mu_{1}(x),\ldots,\kappa_{z_{p}}(x)+\lambda_{z_{p}}(x)-\mu_{1}(x)\right)$
for some choice $\alpha>0$ such that there exists constants $C_{1},C_{2}$ such that $\|\nabla g_{\alpha}(v)\|_{\infty},\|\nabla g_{-\alpha}(v)\|_{\infty}\leq C_{1}$and $\|\nabla^{2}g_{\alpha}(v)\|_{op}\leq C_{2}$ uniformly over $v$. To then extend our estimators, we notice that
$\phi_{\lambda_{z},i}(\eta):=\frac{1\{Z_{i}=z\}}{\mathbb{P}(Z_{i}=z\mid X_{i})}(Y_{i}D_{i}-\lambda_{z}(X_{i}))+\lambda_{z}(X_{i}),$
$\phi_{\kappa_{z},i}(\eta):=\frac{1\{Z_{i}=z\}}{\mathbb{P}(Z_{i}=z\mid X_{i})}((1-D_{i})-\kappa_{z}(X_{i}))+\kappa_{z}(X_{i})$
are the influence functions for $\mathbb{E}[\lambda_{z}(X_{i})]$ and $\mathbb{E}[\kappa_{z}(X_{i})]$ respectively. Furthermore,
$\phi_{\underline{\delta},i}(\eta):=g_{\alpha}(\underline{\delta}_{z_{1}}(X_{i};\eta),\ldots,\underline{\delta}_{z_{p}}(X_{i};\eta))+\sum_{j=1}^{p}\frac{\partial g_{\alpha}(\underline{\delta}_{z_{1}}(X_{i};\eta),\ldots,\underline{\delta}_{z_{p}}(X_{i};\eta))}{\partial\underline{\delta}_{z_{j}}(X_{i};\eta)}\left(\phi_{\lambda_{z_{j}},i}(X_{i})-\phi_{\mu,i}(X_{i})\right)$
$\phi_{\overline{\delta},i}(\eta):=g_{\alpha}(\overline{\delta}_{z_{1}}(X_{i};\eta),\ldots,\overline{\delta}_{z_{p}}(X_{i};\eta))+\sum_{j=1}^{p}\frac{\partial g_{\alpha}(\overline{\delta}_{z_{1}}(X_{i};\eta),\ldots,\overline{\delta}_{z_{p}}(X_{i};\eta))}{\partial\overline{\delta}_{z_{j}}(X_{i};\eta)}\left(\phi_{\kappa_{z_{j}},i}(\eta)+\phi_{\lambda_{z_{j}},i}(X_{i})-\phi_{\mu,i}(X_{i})\right)$
are the influence functions for $\mathbb{E}[\underline{\delta}(X_{i};\eta)]$, $\mathbb{E}[\overline{\delta}(X_{i};\eta)]$ respectively, by Theorem 3 in LevisEtAl(23)-CovariateAssistedIV and standard calculations involving influence functions. With this in hand, we then extend our estimators for the bounds on overall performance, positive class performance, and the conditional likelihood by substituting in the appropriate influence functions for the upper bounding and lower bounding functions.
Alternative Approach under Margin Conditions.
An alternative is to invoke a margin condition as in LevisEtAl(23)-CovariateAssistedIV and semenova2023adaptive. The margin condition assumes that, for each $x\in\mathcal{X}$, the maximal and minimal values in the intersection bounds are sufficiently well-separated from the remaining instrument values. Under such margin conditions, LevisEtAl(23)-CovariateAssistedIV and semenova2023adaptive show that one can construct estimators directly targeting the averages of intersection bounds. For our overall performance bounds, these results apply directly. Extending this approach to estimate the conditional likelihood bounds and the positive class performance bounds is non-trivial. Developing these extensions is an interesting direction for future work.
Appendix E Connections to Sensitivity Analysis Models
A substantial literature in causal inference develops sensitivity analysis frameworks for assessing robustness to unobserved confounding. While these frameworks are often presented in treatment effect settings, they can be translated to our selective labels context. This appendix establishes formal connections between our framework—specifically Assumption 2.1 with observed outcome bounds discussed in Section 2.3.1 of the main text—and several influential sensitivity analysis models.
This is valuable for three reasons. First, pedagogically, researchers familiar with sensitivity analysis in causal inference may find it easier to interpret observed outcome bounds by relating them to established frameworks. Second, these connections show that our estimation results extend to settings where researchers invoke these alternative sensitivity models. Rather than developing separate estimation procedures for each model, researchers could apply our results by translating their sensitivity assumptions into the corresponding values of $\underline{\Gamma},\overline{\Gamma}$ (although the resulting bounds may not be sharp). Finally, most sensitivity analysis models in causal inference place restrictions on selection propensities—how unobservables affect the probability of treatment/selection. In contrast, Assumption 2.1 with observed outcome bounds restricts outcome differences—how outcomes differ between selected and unselected units conditional on observables. While conceptually distinct, Bayes’ rule establishes precise mathematical connections between these two approaches.
E.1 Marginal Sensitivity Model
The marginal sensitivity model (MSM) has received substantial recent attention in the causal inference literature (Tan(06); ZhaoSmallBhattacharya(19); KallusMaoZhou(18)-IntervalEstimation; DornGuo(21); DornGuoKallus(21); KallusZhou(21); JinRenCandes(21)). The MSM specifies that the selection mechanism satisfies a multiplicative bound on odds ratios: that is, for some $\Lambda\geq 1$,
$\Lambda^{-1}\leq\frac{\mathbb{P}(D_{i}=1\mid X_{i},Y_{i}^{*})}{\mathbb{P}(D_{i}=0\mid X_{i},Y_{i}^{*})}\frac{\mathbb{P}(D_{i}=0\mid X_{i})}{\mathbb{P}(D_{i}=1\mid X_{i})}\leq\Lambda\mbox{ holds with probability }1.$ (15)
The MSM bounds the ratio of conditional-to-marginal odds of selection. When $\Lambda=1$, this reduces to $\mathbb{P}(D_{i}=1\mid X_{i},Y^{*}_{i})=\mathbb{P}(D_{i}=1\mid X_{i})$ i.e., unconfoundedness. Larger values of $\Lambda$ allow greater departure from unconfoundedness. In other words, the parameter $\Lambda$ directly quantifies the strength of confounding on the odds ratio scale.
We can relate the MSM to Assumption 2.1 under observed outcome bounds via Bayes’ rule.
Proposition E.1.
  1. i.
    Suppose $(X_{i},D_{i},Y_{i}^{*})\sim\mathbb{P}(\cdot)$ satisfies the MSM (15) for some $\Lambda\geq 1$. Then, $\mathbb{P}(\cdot)$ satisfies Assumption 2.1 with $\underline{\delta}(x;\eta)=(\Lambda^{-1}-1)\mu_{1}(x)$ and $\overline{\delta}(x;\eta)=(\Lambda-1)\mu_{1}(x)$.
  2. ii.
    Suppose $(X_{i},D_{i},Y_{i}^{*})\sim\mathbb{P}(\cdot)$ satisfies Assumption 2.1 with observed outcome bounds for some $\underline{\Gamma},\overline{\Gamma}>0$. Then, $\mathbb{P}(\cdot)$ satisfies
    $\overline{\Gamma}^{-1}\leq\frac{\mathbb{P}(D_{i}=1\mid Y_{i}^{*}=1,X_{i})\mathbb{P}(D_{i}=0\mid X_{i})}{\mathbb{P}(D_{i}=0\mid Y_{i}^{*}=1,X_{i})\mathbb{P}(D_{i}=1\mid X_{i})}\leq\underline{\Gamma}^{-1}.$
Proof.
For brevity, we omit the conditioning on $X_{i}$. Consider the first claim. By Bayes’ rule, $\frac{\mathbb{P}(D_{i}=1\mid Y_{i}^{*})\mathbb{P}(D_{i}=0)}{\mathbb{P}(D_{i}=0\mid Y_{i}^{*})\mathbb{P}(D_{i}=1)}=\frac{\mathbb{P}(Y_{i}^{*}\mid D_{i}=1)}{\mathbb{P}(Y_{i}^{*}\mid D_{i}=0)}$. The MSM therefore implies bounds $\Lambda^{-1}\leq\frac{\mathbb{P}(Y_{i}^{*}\mid D_{i}=1)}{\mathbb{P}(Y_{i}^{*}\mid D_{i}=0)}\leq\Lambda$, which can be equivalently written as $\Lambda^{-1}\mathbb{P}(Y_{i}^{*}\mid D_{i}=1)\leq\mathbb{P}(Y_{i}^{*}\mid D_{i}=0)\leq\Lambda\mathbb{P}(Y_{i}^{*}\mid D_{i}=1)$. Adding and subtracting $\mathbb{P}(Y_{i}^{*}=1\mid D_{i}=1)$ then delivers the first claim.
Consider the second claim. Observed outcome bounds implies $\overline{\Gamma}^{-1}\leq\frac{\mathbb{P}(Y_{i}^{*}=1\mid D_{i}=1)}{\mathbb{P}(Y_{i}^{*}=1\mid D_{i}=0)}\leq\underline{\Gamma}^{-1}$. But, by Bayes’ rule, $\frac{\mathbb{P}(Y_{i}^{*}=1\mid D_{i}=1)}{\mathbb{P}(Y_{i}^{*}=1\mid D_{i}=0)}=\frac{\mathbb{P}(D_{i}=1\mid Y_{i}^{*}=1)\mathbb{P}(D_{i}=0)}{\mathbb{P}(D_{i}=0\mid Y_{i}^{*}=1)\mathbb{P}(D_{i}=1)}$, and so $\overline{\Gamma}^{-1}\leq\frac{\mathbb{P}(D_{i}=1\mid Y_{i}^{*}=1)\mathbb{P}(D_{i}=0)}{\mathbb{P}(D_{i}=0\mid Y_{i}^{*}=1)\mathbb{P}(D_{i}=1)}\leq\underline{\Gamma}^{-1}$ holds. ∎
Proposition E.1 shows that any analysis conducted under the MSM with parameter $\Lambda$ can be conducted using observed outcome bounds with $\underline{\Gamma}=\Lambda^{-1}$ and $\overline{\Gamma}=\Lambda$. In the opposite direction, observed outcome bounds imply a specific MSM-style restriction on selection propensities, but only for the positive class $Y_{i}^{*}=1$.
E.2 Rosenbaum’s $\Gamma$-Sensitivity Model
Rosenbaum’s $\Gamma$-sensitivity model (e.g., Rosenbaum(87); Rosenbaum(02)) is perhaps the most widely used framework for sensitivity analysis in observational studies. It bounds how much unobservables can differentially affect selection propensities across different outcome values: specifically, for some $\Gamma\geq 1$, $(X_{i},D_{i},Y_{i}^{*})\sim\mathbb{P}(\cdot)$ satisfies
$\Gamma^{-1}\leq\frac{\mathbb{P}(D_{i}=1\mid X_{i},Y_{i}^{*}=y^{*}}{\mathbb{P}(D_{i}=0\mid X_{i},Y_{i}^{*}=y^{*})}\frac{\mathbb{P}(D_{i}=0\mid X_{i},Y_{i}^{*}=\tilde{y}^{*})}{\mathbb{P}(D_{i}=1\mid X_{i},Y_{i}^{*}=\tilde{y}^{*})}\leq\Gamma$ (16)
for all $y^{*},\tilde{y}^{*}\in\{0,1\}$ and with probability one. Notice that $\Gamma=1$ again nests unconfoundedess.
AronowLee(2013) and MiratrixWagerZubizaretta(18) use a version of Rosenbaum’s $\Gamma$-sensitivity model to construct bounds on a finite-population from a random sample with unknown selection probabilities. yadlowsky2018bounds applies Rosenbaum’s $\Gamma$-sensitivity analysis model to derive bounds on the conditional average treatment effect and the average treatment effect. We refer the reader to Section 7 of ZhaoSmallBhattacharya(19) for an in-depth comparison of the marginal sensitivity model and Rosenbaum’s $\Gamma$-sensitivity model.
We can relate Rosenbaum’s $\Gamma$-sensitivity model to Assumption 2.1 under observed outcome bounds.
Proposition E.2.
Suppose $(X_{i},D_{i},Y_{i}^{*})\sim\mathbb{P}(\cdot)$ satisfies Rosenbaum’s sensitivity analysis model (16) for some $\Gamma>1$. Then $P(\cdot)$ satisfies Assumption 2.1 with $\underline{\delta}(x;\eta)=(\Gamma^{-1}-1)\mu_{1}(x)$ and $\overline{\delta}(x;\eta)=(\Gamma-1)\mu_{1}(x)$.
Proof.
For brevity, we omit the conditioning on $X_{i}$ throughout the proof. As a first step, we again apply Bayes’ rule and observe that $\frac{\mathbb{P}(Y_{i}^{*}\mid D_{i}=1)}{\mathbb{P}(Y_{i}^{*}\mid D_{i}=0)}=\frac{\mathbb{P}(D_{i}=1\mid Y_{i}^{*})\mathbb{P}(D_{i}=0)}{\mathbb{P}(D_{i}=0\mid Y_{i}^{*})\mathbb{P}(D_{i}=1)}$. Then, further notice that $\frac{\mathbb{P}(D_{i}=0)}{\mathbb{P}(D_{i}=1)}=\frac{\sum_{y\in\{0,1\}}P(D_{i}=0\mid Y_{i}^{*}=y)P(Y_{i}^{*}=y)}{\sum_{y\in\{0,1\}}P(D_{i}=1\mid Y_{i}^{*}=y)P(Y_{i}^{*}=y)}.$ Letting $y^{*}=\arg\max_{y^{*}\in\{0,1\}}\frac{P(D_{i}=0\mid Y_{i}(0)=y_{0},Y_{i}(1)=y_{1})}{P(D_{i}=1\mid Y_{i}(0)=y_{0},Y_{i}(1)=y_{1})}$, the quasi-linearity of the ratio function implies that
$\frac{\sum_{y\in\{0,1\}}P(D_{i}=0\mid Y_{i}^{*}=y)P(Y_{i}^{*}=y)}{\sum_{y\in\{0,1\}}P(D_{i}=1\mid Y_{i}^{*}=y)P(Y_{i}^{*}=y)}\leq\frac{P(D_{i}=0\mid Y_{i}^{*}=y^{*})}{P(D_{i}=1\mid Y_{i}^{*}=y^{*})}$
This implies, for any $y\in\{0,1\}$,
$\frac{\mathbb{P}(Y_{i}^{*}=y\mid D_{i}=1)}{\mathbb{P}(Y_{i}^{*}=y\mid D_{i}=0)}\leq\frac{\mathbb{P}(D_{i}=1\mid Y_{i}^{*}=y)}{\mathbb{P}(D_{i}=0\mid Y_{i}^{*}=y)}\frac{P(D_{i}=0\mid Y_{i}^{*}=y^{*})}{P(D_{i}=1\mid Y_{i}^{*}=y^{*})}\leq\Gamma,$
where the last inequality is implied by Rosenbaum’s sensitivity analysis model (16). From this, we follow the same argument as the proof of Proposition E.1 to show that $\overline{\delta}(x;\eta)=(\Gamma-1)\mu_{1}(x)$. The proof for the lower bound follows an analogous argument. ∎
Rosenbaum’s $\Gamma$-sensitivity model with parameter implies observed outcome bounds with $\underline{\Gamma}=\Gamma^{-1}$ and $\overline{\Gamma}=\Gamma$. However, the reverse direction does not hold — observed outcome bounds are weaker than Rosenbaum’s model.
E.3 Partial $c$-Dependence
A recent strand of the econometrics literature proposes the partial c-dependence model for conducting sensitivity analyses on unobserved confounding in causal inference (e.g., MastenPoirier(18); MastenPoirier(20)). Cast into our setting, partial $c$-dependence specifies that, for some $c>0$,
$\sup_{y\in\{0,1\}}\,|\mathbb{P}(D_{i}=1\mid Y_{i}^{*}=y,X_{i}=x)-\mathbb{P}(D=1\mid X_{i}=x)|\leq c$ (17)
for all $x\in\mathcal{X}$. Partial $c$-dependence bounds how much observing the outcome changes the selection problem. When $c=0$, we have unconfoundedness. The parameter $c$ represents the maximum percentage point difference in selection probabilities between different outcome values. The next result shows that Assumption 2.1 and partial c-dependence model can be related to one another through Bayes’ rule.
Proposition E.3.
  1. (i)
    Suppose $(X_{i},D_{i},Y_{i}^{*})\sim\mathbb{P}(\cdot)$ satisfies $\left|\mathbb{P}(Y_{i}^{*}=1\mid D_{i}=0,X_{i}=x)-\mathbb{P}(Y_{i}^{*}=1\mid D_{i}=1,X_{i}=x)\right|\leq C$ for all $x\in\mathcal{X}$. Then, $\sup_{y\in\{0,1\}}\,|\mathbb{P}(D_{i}=1\mid Y_{i}^{*}=y,X_{i}=x)-\mathbb{P}(D=1\mid X_{i}=x)|\leq c(x)$ for all $x\in\mathcal{X}$, where $c(x)=C\ \max\{\frac{V(D_{i}\mid X_{i}=x)}{\mathbb{P}(Y_{i}^{*}=1\mid X_{i}=x)},\frac{V(D_{i}\mid X_{i}=x)}{\mathbb{P}(Y_{i}^{*}=0\mid X_{i}=x)}\}$.
  2. (ii)
    Suppose $(X_{i},D_{i},Y_{i}^{*})\sim\mathbb{P}(\cdot)$ satisfies $\sup_{y\in\{0,1\}}\,|\mathbb{P}(D_{i}=1\mid Y_{i}^{*}=y,X_{i}=x)-\mathbb{P}(D=1\mid X_{i}=x)|\leq c$ for all $x\in\mathcal{X}$. Then, $\left|\mathbb{P}(Y_{i}^{*}=1\mid D_{i}=0,X_{i}=x)-\mathbb{P}(Y_{i}^{*}=1\mid D_{i}=1,X_{i}=x)\right|\leq c\ \min\{\frac{V(D_{i}\mid X_{i}=x)}{\mathbb{P}(Y_{i}^{*}=1\mid X_{i}=x)},\frac{V(D_{i}\mid X_{i}=x)}{\mathbb{P}(Y_{i}^{*}=0\mid X_{i}=x)}\}$ for all $x\in\mathcal{X}$.
Proof.
For brevity, we omit the conditioning on $X_{i}$ throughout the proof. To show the first result, we first rewrite $|\mathbb{P}(Y_{i}^{*}=1\mid D_{i}=0)-\mathbb{P}(Y_{i}^{*}=1\mid D_{i}=1)|\leq C$ as $|\mathbb{P}(D_{i}=0\mid Y_{i}^{*}=1)\mathbb{P}(D_{i}=1)-\mathbb{P}(D_{i}=1\mid Y_{i}^{*}=1)\mathbb{P}(D_{i}=0)|\leq C\frac{Var(D_{i})}{\mathbb{P}(Y_{i}^{*}=1)}$. We can further rewrite the left hand side to arrive at $|\mathbb{P}(D_{i}=1\mid Y_{i}^{*}=1,X_{i}=x)-\mathbb{P}(D_{i}=1\mid X_{i}=x)|\leq C\frac{Var(D_{i})}{\mathbb{P}(Y_{i}^{*}=1)}$. Analogously, we can rewrite the bound as $|\mathbb{P}(Y_{i}^{*}=0\mid D_{i}=1,X_{i}=x)-\mathbb{P}(Y_{i}^{*}=0\mid D_{i}=0,X_{i}=x)|\leq C$, which in turn implies $|\mathbb{P}(D_{i}=1\mid Y_{i}^{*}=0,X_{i}=x)-\mathbb{P}(D_{i}=1\mid X_{i}=x)\leq C\frac{Var(D_{i})}{\mathbb{P}(Y_{i}^{*}=0)}$ by the same argument. The second result follows by the same argument. ∎
As mentioned, partial $c$-dependence has an interpretation in terms of percentage point changes in selection probabilities, while observed outcome bounds has an interpretation in terms of changes in outcome probabilities. In applications like pretrial release or consumer lending, the appropriate value of $c$ may be challenging to calibrate — it requires articulating assumptions about how judge and lender decision rates differ across defendants and applicants who would versus would not experience adverse outcomes. Observed outcome bounds ask researchers to directly articulate assumptions about outcome rates directly (e.g., ”rejected applicants are at most twice as likely to default”). This may align more naturally with available domain knowledge or proxy data.
Of course, researchers should choose the parameterization that best aligns with their substantive knowledge and available evidence. Moreover, Proposition E.3 provides explicit formulas that allow researchers to translate between the two frameworks, enabling those who find it easier to specify partial $c$-dependence to map to implied observed outcome bounds (and vice versa).
E.4 Tukey’s Factorization and Outcome Models
A alternative approach to sensitivity analysis specifies outcome models via Tukey’s factorization that directly parameterize the relationship between $\mathbb{P}(Y^{*}_{i}\mid D_{i}=0,X_{i})$ and $\mathbb{P}(Y^{*}_{i}\mid D_{i}=1,X_{i})$ or $\mathbb{P}(Y^{*}_{i}\mid X_{i})$ (e.g., RotnitzkyEtAl(01); BirminghamEtAl(03); brumback2004sensitivity). For example, RobinsEtAl(00); FranksEtAl(19); Scharfstein(21) assume $\mathbb{P}(Y_{i}^{*}\mid D_{i}=0,X_{i})=\mathbb{P}(Y_{i}^{*}\mid D_{i}=1,X_{i})\frac{\exp(\gamma_{t}s_{t}(Y_{i}^{*}))}{C(\gamma_{t};X_{i})}$, where $\gamma_{t}$ is a chosen parameter and $s_{t}(\cdot)$ is a chosen “tilting function.” For particular fixed choices of $\gamma_{t}$, $s_{t}(\cdot)$, such a model is sufficient to point identify various quantities of interest such as the conditional probability and the predictive performance estimands we consider.
The key difference relative to our framework is how uncertainty is communicated. Researchers in this literature report how conclusions vary for alternative choices of $\gamma_{t}$ or $s_{t}(\cdot)$. By contrast, we suggest that researchers should articulate bounding functions that encode their beliefs about unobserved confounding, and then report the fully identified set consistent with those bounds. These bounding functions can formalize common heuristics without requiring a fully specified outcome model that can be difficult to justify.
An alternative approach places bounds on the mean difference in potential outcomes under treatment and control luedtke2015statistics; diaz2013sensitivity; diaz2018sensitivity. This is analogous to the “approximate mean independence” assumption in Manski(03). Our identification framework extends this approach by placing bounds on the covariate-conditional difference in means.
Appendix F Monte Carlo Simulations
In this section, we illustrate that the finite sample behavior of our proposed estimators. We simulate training and evaluation datasets satisfying Assumption 2.1 under observed outcome bounds. We draw $X_{i}=\left(X_{i,1},\ldots,X_{i,d}\right)^{\prime}\sim N(0,I_{d})$ and $D_{i}$ conditional on $X_{i}$ according to $\mathbb{P}(D_{i}=1\mid X_{i}=x)=\sigma\left(\frac{1}{2\sqrt{d_{\pi}}}\sum_{d=1}^{d_{\pi}}X_{i,d}\right)$ for some $d_{\pi}\in\{1,\ldots,d\}$ and $\sigma(a)=\frac{\exp(a)}{1+\exp(a)}$. We finally draw $Y_{i}^{*}$ conditional on $(D_{i},X_{i})$ according to
$\mathbb{P}(Y_{i}^{*}=1\mid D_{i}=1,X_{i}=x)=\sigma\left(\frac{1}{2\sqrt{d_{\mu}}}\sum_{d=1}^{d_{\mu}}X_{i,d}\right),\mbox{ and }\mathbb{P}(Y_{i}^{*}=1\mid D_{i}=0,X_{i}=x)=\Gamma_{true}\sigma\left(\frac{1}{2\sqrt{d_{\mu}}}\sum_{d=1}^{d_{\mu}}X_{i,d}\right)$
for some $d_{\mu}\in\{1,\ldots,d\}$ and $0<\Gamma_{true}<1$ We set $d=50$, $d_{\pi}=20$, $d_{\mu}=25$, and $\Gamma_{true}=0.75$.
F.1 Behavior of Estimated Conditional Probability Bounds
We compare three estimators for the upper bound on the conditional likelihood $\overline{\mu}^{*}(\cdot)$ with $\underline{\Gamma}=2/3$, $\overline{\Gamma}=3/2$: first, our proposed estimators; second, the infeasible oracle that uses the true observed conditional probability and propensity score; and finally, a plug-in learner that does not use sample splitting nor pseudo-outcomes. Our estimator and the oracle procedure are constructed using a single split of the evaluation data. The first-stage nuisance functions $\eta=(\pi_{1}(\cdot),\mu_{1}(\cdot))$ are estimated using cross-validated Lasso logistic regressions, and the second-stage nonparametric regression uses cross-validated Lasso. Proposition 3.1 established that the integrated mean-square error of our estimators converge at fast rates to the integrated mean-square error of the infeasible oracle. By contrast, the plug-in estimator will inherit errors from the estimation of the nuisance functions.
Figure A1: Average integrated mean square of our estimator, the oracle learner, and the plug-in learner for the upper bound on the conditional probability $\overline{\mu}^{*}(\cdot)$.
[…]
F.2 Behavior of Estimated Predictive Performance Bounds
To evaluate our estimators for the bounds on positive class performance and overall performance, we simulate a training dataset $(X_{i},D_{i},Y_{i})$ for $i=1,\ldots,n_{train}$ and estimate a predictive algorithm $s(\cdot)$ that predicts $Y_{i}=1$ only on the selected data $D_{i}=1$. We evaluate the constructed predictive algorithm’s positive class performance $\mbox{perf}_{+}(s;\beta)$ and overall performance $\mbox{perf}(s;\beta)$ using evaluation datatsets $(X_{i},D_{i},Y_{i})$ for $i=1,\ldots,n$ simulated from the same data generating process and observed outcome bounds for alternative choices of $\underline{\Gamma},\overline{\Gamma}$.
$n$ Estimate SD Bias
500 0.545 0.016 0.001
1000 0.544 0.010 0.000
2500 0.544 0.007 0.000
Table A1: Bias properties for estimator for the upper bound on the true positive rate.
We explore how well our proposed estimator recover the bound on the true positive rate $\overline{\mbox{perf}}_{+}(s;\beta)$ across $1000$ simulated datasets of varying size $n\in\{500,1000,2500\}$ and a fixed choice $\underline{\Gamma}=2/3$ and $\overline{\Gamma}=3/2$. We construct our estimator using three splits of the evaluation data. We estimate the first-stage nuisance functions using random forests in the first fold, and the estimated bounding functions using cross-validated Lasso in the second fold. Online Appendix Table A1 reports the average bias, and Online Appendix Figure A2 visualizes the density of our estimates across simulations. Their bias quickly converges to zero as the size of the evaluation data grows large. We further explore how the performance of proposed estimators varies as we vary the researcher’s assumptions on the strength of unobserved confounding. Online Appendix Figure A3 reports the simulation distribution of our proposed estimators for both the upper and lower bounds as we vary $\overline{\Gamma}=\widetilde{\Gamma}$ for $\widetilde{\Gamma}\in\{1,\ldots,2.5\}$ and set $\underline{\Gamma}=1/\widetilde{\Gamma}$.
Figure A2: Distribution of estimator for the upper bound on the true positive rate across Monte Carlo simulations with observed outcome bounds
Figure A3: Estimated bounds on the true positive rate as $\underline{\Gamma},\overline{\Gamma}$ varies.
Our proposed estimators for the overall performance bounds follow a standard debiased construction. We briefly illustrate that our proposed estimators for the overall performance bounds also converge quickly and confidence intervals based on the asymptotic normal approximation have good coverage properties. For $\underline{\Gamma}=2/3$, $\overline{\Gamma}=3/2$ and focusing on the lower bound on mean square error and the upper bound on accuracy, Online Appendix Table A2 reports the average bias of our estimators and the estimated coverage rate of 95% nominal confidence intervals. Our proposed estimators are approximately unbiased for the true bounds. Their estimated standard errors slightly underestimate the true standard errors when the size of the evaluation dataset is small, but are quite accurate for $n\geq 1000$. Consequently, confidence intervals based on the asymptotic normal approximation have approximately 95% coverage. Our proposed estimators are approximately normally distributed in finite samples and concentrate around the true bounds (Online Appendix Figure A4).
$n$ Bias SD $\widehat{\sigma}$ Coverage
500 -0.001 0.011 0.011 0.950
1000 -0.001 0.008 0.008 0.939
2500 -0.001 0.004 0.008 0.939
(a) Lower bound on mean square error
$n$ Bias SD $\widehat{\sigma}$ Coverage
500 0.001 0.011 0.011 0.950
1000 0.001 0.008 0.008 0.934
2500 0.001 0.005 0.005 0.942
(b) Upper bound on accuracy
Table A2: Bias and coverage properties with observed outcome bounds.
(a) Upper bound on mean square error
(b) Lower bound on accuracy
Figure A4: Distribution of overall performance estimators across Monte Carlo simulations with observed outcome bounds.
Appendix G Additional Tables for Empirical Application
In this section, we report an additional table for the empirical application that is referenced in Section 5 of the main text. Table A3 provides detailed descriptions of the variable names in right panel of Figure 1 in Section 5.1 of the main text. Personal loans applications can have multiple listed applicants, so some variables refer to just the first listed applicant.
Variable name Detailed description Total net income Total net income for all applicants on the personal loan application Occupation type Industry code of 1st applicant’s occupation. Mos in current employment Number of months 1st applicant has held current job. Max delinquency in 12 mos Maximum delinquency over last 12 months (home loan, personal loan or credit card). Exposure to loan amount Exposure to requested personal loan amount. Existing personal loan balance Existing personal loan balance of applicants. Current days in debt Current number of days in debt of all applicants. Credit bureau score External credit score. Accommodation status Type of accommodation applicant currently occupies (e.g., owned, rented, etc). # of credit card apps in 12 mos (all applicants) Number of credit card applications submitted by all applications in last 12 months. # of credit card apps in 12 mos (1st applicant) Number of credit card applications submitted by 1st applicant in last 12 months. # of check acct payment reversals in 6 mos (all applicants) Number of checking account payment reversals by all applicants in last 6 months. # of check acct payment reversals in 6 mos (1st applicants) Number of checking account payment reversals by first applicant in last 6 months.
Table A3: Detailed description of variable names in right panel of Table 1.